### v1

# Bounds on the number of squares in recurrence sequences

Abstract.
We investigate the number of squares in a very broad family of binary recurrence sequences with $u_{0}=1$. We show that there are at most two distinct squares in such sequences (the best possible result), except under such very special conditions where we prove there are at most three such squares.
Key words and phrases:
binary recurrence sequences; Diophantine approximations.
1. Introduction
1.1. Background
The study of the arithmetic properties of recurrence sequences has a long history. Important problems include the open question of whether there are infinitely many primes in such sequences, lower bounds for the largest prime divisor of the $n$-th element [13], the zero-multiplicity of such sequences (how often zero occurs as an element) [10] and also the occurrence of powers in them [4].
While some progress on arithmetic questions has been achieved for sequences of Lucas numbers (numbers of the form $u_{n}=\left(\alpha^{n}-\beta^{n}\right)/(\alpha-\beta)$, where $\alpha$ and $\beta$ are algebraic numbers such that $\alpha+\beta$ and $\alpha\beta$ are non-zero coprime rational numbers and $\alpha/\beta$ is not a root of unity) and their associated sequences ($v_{n}=\alpha^{n}+\beta^{n}$), much less is known for other binary recurrence sequences.
As with sequences of Lucas numbers, other binary recurrence sequences are important for the solution of Diophantine equations such as $aX^{2}-bY^{4}=c$, which is important in its own right, but is also a quartic model of elliptic curves.
This paper is concerned with such problems, in particular how many distinct squares can occur in non-degenerate binary recurrence sequences.
1.2. Notation
To define the sequences we are interested in and to express our results, we start with some notation.
Let $a$, $b$ and $d$ be positive integers such that $d$ is not a square. Suppose $\alpha=a+b^{2}\sqrt{d}$ has $N_{\alpha}=N_{\mathbb{Q}\left(\sqrt{d}\right)/\mathbb{Q}}(\alpha)=a^{2}-b^{4}d$ and let $\varepsilon=\left(t+u\sqrt{d}\right)/2$ be a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ positive integers.
We define the two sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ by
(1.1) $x_{k}+y_{k}\sqrt{d}=\alpha\varepsilon^{2k}.$
Observe that $x_{0}=a$, $y_{0}=b^{2}$,
[…]
and that both sequences satisfy the recurrence relation
(1.3) $u_{k+1}=\frac{t^{2}+du^{2}}{2}u_{k}-u_{k-1},$
for all $k\in\mathbb{Z}$. Note that $\left(t^{2}+du^{2}\right)/2=\operatorname{Tr}_{\mathbb{Q}\left(\sqrt{d}\right)/\mathbb{Q}}\left(\varepsilon^{2}\right)$.
To relate such sequences back to the quadratic equations mentioned in the previous section, observe that from (1.1),
$x_{k}^{2}-dy_{k}^{2}=N_{\alpha}.$
We restrict to the coefficient of $\sqrt{d}$ in $\alpha$ being a square since here we are interested in squares in the sequence of $y_{k}$’s. For such problems, we will choose $\alpha$ such that $b^{2}$ is the smallest square among the $y_{k}$’s.
For any non-zero integer, $n$, let $\operatorname{sf}(n)$ be the unique squarefree integer such that $n/\operatorname{sf}(n)$ is a square. We will put $\operatorname{sf}(1)=1$.
1.3. Conjectures
We start with some conjectures regarding squares in the sequence of $y_{k}$’s. The dependence on the arithmetic of $N_{\alpha}$ is noteworthy.
Conjecture 1.1.
There are at most four distinct integer squares among the $y_{k}$’s.
If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then there are at most three distinct integer squares among the $y_{k}$’s.
Furthermore, if $\left|N_{\alpha}\right|$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
When $N_{\alpha}<0$ and $b=1$, a stronger result appears to hold.
Conjecture 1.2.
Suppose that $b=1$ and $N_{\alpha}<0$. There are at most three distinct integer squares among the $y_{k}$’s.
If $-N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
If we replace $\varepsilon^{2k}$ in (1.1) with $\varepsilon^{k}$, it appears that the following is true.
Conjecture 1.3.
There are at most four distinct integer squares among the $y_{k}$’s.
If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=p$ where $p$ is prime, or if $\left|N_{\alpha}\right|$ is a perfect square, then there are at most three distinct integer squares among the $y_{k}$’s.
We used PARI/GP [9] to search for squares among $y_{-60},\ldots,y_{60}$ from Conjecture 1.3, where $\varepsilon=\left(t+u\sqrt{d}\right)/2>1$ is the smallest algebraic integer that is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t,u\in\mathbb{Z}$ and $u\leq 10^{30}$. Note that this also includes searching among $y_{-30},\ldots,y_{30}$ from Conjectures 1.1 and 1.2. The range of $b$ and $d$ in this search was $1\leq b,d\leq 1000$ with $d$ squarefree. For $a$, we searched over two ranges: (1) $\max\left(1,\lfloor\sqrt{db^{4}}\rfloor-500\right)\leq a\leq\lceil\sqrt{db^{4}}\rceil+500$ (so $\left|N_{\alpha}\right|$ is small) and (2) $1\leq a\leq 1000$. For both ranges of $a$, we considered only $\gcd\left(a,b^{2}\right)$ squarefree. No counterexamples to any of these conjectures were found in this search.
The largest value of $|k|$ found in this search such that $y_{k}$ is a square was $k=-42$. One such example was found: $y_{-42}=12223^{2}$ for $(a,b,d,t,u)=(502,15,5,1,1)$. We also observed that no squares were found among the $y_{k}$’s with $|k|>30$ for $d>5$ and with $|k|>18$ for $d>13$.
Remark.
The distinctness condition in these conjectures, and in our results below, is important, as such sequences can have repeated elements. E.g., $(a,b,d,t,u)=(42,4,7,16,6)$ where $y_{-k}=y_{k-1}$ for all $k\geq 1$. But this can only happen when $\alpha$ divided by its algebraic conjugate is a unit in the ring of integers.
1.4. Results
In this paper, we prove Conjecture 1.2 when $-N_{\alpha}$ is a square, except for a very limited set of sequences, where we prove a slightly weaker result.
Theorem 1.4.
Let $b=1$, $a$ and $d$ be positive integers, where $d$ is not a square, $N_{\alpha}<0$ and $-N_{\alpha}$ is a square.
(a) If $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$, $\gcd\left(a^{2},d\right)=1,4$ and one of $y_{\pm 1}$ is a perfect square, then there are at most three distinct squares among the $y_{k}$’s.
(b) If $u=2$, $t^{2}-du^{2}=-4$, $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)=1$ and one of $y_{\pm 1}$ is a perfect square, then there are at most three distinct squares among the $y_{k}$’s.
(c) Otherwise, there are at most two distinct squares among the $y_{k}$’s.
1.5. Our method of proof
Our proof has its basis in the work of Siegel [12] (also see [6]). Foremost with our use of hypergeometric functions. But also with how we use them here. The definition and use of our quantity $r_{0}$ in Section 4 has similarities to the definition and use of $\ell_{1}$ and $\ell_{2}$ from Lemma 7 onwards in [6]. The idea is to treat ranges of hypothetical squares beyond those that can be treated by our gap principle alone.
We have state and prove our results in Section 3 more generally than required here. This is because they will be useful in further work on binary recurrence sequences. An example is [18] where we prove Conjecture 1.2 when $N_{\alpha}$ is a prime power.
When $N_{\alpha}>0$ is not a square, we can apply Thue’s Fundamentaltheorem (see [14] and [15]) to obtain similar results.
1.6. Structure of this article
Section 2 concerns results on diophantine approximation, with a focus on the use of hypergeometric functions to do so. In Section 3, we collect various facts that we will require about elements of the sequence $\left(y_{k}\right)_{k=-\infty}^{\infty}$ defined by (1.1). These two sections are independent of each other. Their results are brought together in Section 4, where we state and prove Proposition 4.1. Our theorem follows from this proposition. Its proof is given in Section 5. Finally, in Section 6, we give examples showing that our results and conjectures are best possible.
2. Diophantine Approximation via Hypergeometric Functions
Recall that by an effective irrational measure for an irrational number, $\alpha$, we mean an inequality of the form
$\left|\alpha-\frac{p}{q}\right|>\frac{c}{|q|^{\mu}},$
for all $p/q\in\mathbb{Q}$ with $\gcd(p,q)=1$ and $|q|>Q$, where $c$, $Q$ and $\mu$ are all effectively computable.
By Liouville’s famous result [7], where he constructed the first examples of numbers proven to be transcendental, we have such effective irrational measures for algebraic numbers of degree $n$, with $\mu=n$. But for practically all applications we require $\mu<n$.
We can sometimes use the hypergeometric method to obtain effective irrationality measures that improve on Liouville’s result for the algebraic numbers that arise here. This was first done by Baker [2, 3] building on earlier applications of hypergeometric functions to diophantine approximation such as those of Siegel [11]. However, that does not suffice for us to prove our results here. The problem here arises not because of the exponent, $\mu$, in the effective irrationality measure, but because the constant, $c$, is too large. Upon investigating this further, we found that we can complete the proof of our results if we use not the effective irrationality measures from the hypergeometric method, but rather consider more carefully the actual results that we obtain from the use of hypergeometric functions.
The means of doing so is the following lemma.
Lemma 2.1.
Let $\theta\in\mathbb{C}$ and let $\mathbb{K}$ be an imaginary quadratic field. Suppose that there exist $k_{0},\ell_{0}>0$ and $E,Q>1$ such that for all non-negative integers $r$, there are algebraic integers $p_{r}$ and $q_{r}$ in $\mathbb{K}$ with $\left|q_{r}\right|<k_{0}Q^{r}$ and $\left|q_{r}\theta-p_{r}\right|\leq\ell_{0}E^{-r}$ satisfying $p_{r}q_{r+1}\neq p_{r+1}q_{r}$.
For any algebraic integers $p$ and $q$ in $\mathbb{K}$, let $r_{0}$ be the smallest positive integer such that $\left(Q-1/E\right)\ell_{0}|q|/\left(Q-1\right)<cE^{r_{0}}$, where $10<c<$ 1.
(a) We have
$\left|q\theta-p\right|>\frac{1-c/E}{k_{0}Q^{r_{0}+1}}.$
(b) When $p/q\neq p_{r_{0}}/q_{r_{0}}$, we have
$\left|q\theta-p\right|>\frac{1-c}{k_{0}Q^{r_{0}}}.$
Remark.
This is a modification of Lemma 6.1 of [15] (and other similar results).
First, we replace $1/\left(2k_{0}\right)$ in both parts and define $r_{0}$ somewhat differently too.
Second, the terms in the lower bounds in both parts are no longer converted into ones involving $|q|^{-(\kappa+1)}$. For our use here, where an effective irrationality measure is not required, this results in further improvements.
Proof.
Let $p$, $q$ be algebraic integers in $\mathbb{K}$. Let $R=q\theta-p$ and $R_{r}=q_{r}\theta-p_{r}$. We can write
$q_{r}R=q_{r}q\theta-q_{r}p=q\left(q_{r}\theta-p_{r}\right)+p_{r}q-q_{r}p=qR_{r}+p_{r}q-q_{r}p.$
We consider two cases according to whether $p_{r_{0}}q-q_{r_{0}}p=0$ or not.
If $p_{r}q-q_{r}p\neq 0$, then $\left|p_{r}q-q_{r}p\right|\geq 1$, since it is an algebraic integer in an imaginary quadratic field. From our upper bounds for $\left|q_{r}\right|$ and $\left|q_{r}\theta-p_{r}\right|$ in the statement of the lemma, we have
$k_{0}Q^{r}|R|>1-\ell_{0}E^{-r}|q|.$
With $r=r_{0}$, from the definition of $r_{0}$, we have
$k_{0}Q^{r}|R|>1-c\frac{Q-1}{Q-1/E}>1-c.$
Part (b) now follows from the upper bound for $\left|q_{r_{0}}\right|$ in the statement of the lemma.
If $p_{r_{0}}q-q_{r_{0}}p=0$, then we use $r=r_{0}+1$. From the definition of $r_{0}$, we have
$k_{0}Q^{r_{0}+1}|R|>1-c\frac{Q-1}{E(Q-1/E)}=\frac{QE-1-cQ+c}{QE-1}>1-c/E.$
Part (a) now follows. ∎
2.1. Construction of Approximations
Let $t^{\prime}$, $u_{1}$ and $u_{2}$ be rational integers with $t^{\prime}<0$  [^1] To avoid confusion with $t$ in the expression for $\varepsilon$, we use $t^{\prime}$ here to denote what was $t$ in our previous works like [16]. Similarly, we will use $d^{\prime}$ where $d$ was used in previous works.. We let $u=\left(u_{1}+u_{2}\sqrt{t^{\prime}}\right)/2$ be an algebraic integer in $\mathbb{K}=\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$ with $\sigma(u)=\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/2$ as its algebraic (and complex) conjugate. Put $\omega=u/\sigma(u)$ and write $\omega=e^{i\varphi}$, where $-\pi<\varphi\leq\pi$. For any real number $\nu$, we shall put $\omega^{\nu}=e^{i\nu\varphi}$ – unless otherwise stated, we will use this convention throughout this paper.
Suppose that $\alpha$, $\beta$ and $\gamma$ are complex numbers and $\gamma$ is not a non-positive integer. We denote by ${}_{2}F_{1}(\alpha,\beta,\gamma,z)$ the classical (or Gauss) hypergeometric function of the complex variable $z$.
For integers $m$ and $n$ with $0<m<n$, $(m,n)=1$ and $r$ a non-negative integer, put $\nu=m/n$ and
$X_{m,n,r}(z)={}_{2}F_{1}(-r-\nu,-r,1-\nu,z),\quad Y_{m,n,r}=z^{r}X_{m,n,r}\left(z^{-1}\right)$
and
$R_{m,n,r}(z)=(z-1)^{2r+1}\frac{\nu\cdots(r+\nu)}{(r+1)\cdots(2r+1)}{}_{2}F_{1}\left(r+1-\nu,r+1;2r+2;1-z\right).$
We collect here some facts about these functions that we will require.
Lemma 2.2.
(a) Suppose that $|\omega-1|<1$. We have
$\omega^{\nu}Y_{m,n,r}(\omega)-X_{m,n,r}(\omega)=R_{m,n,r}(\omega).$
(b) We have
$X_{m,n,r}(\omega)Y_{m,n,r+1}(\omega)\neq X_{m,n,r+1}(\omega)Y_{m,n,r}(\omega).$
(c) If $|\omega|=1$ and $|\omega-1|<1$, then
$\left|R_{m,n,r}(\omega)\right|\leq\frac{\Gamma(r+1+\nu)}{r!\Gamma(\nu)}|\varphi|\left|1-\sqrt{\omega}\right|^{2r}.$
(d) If $|\omega|=1$ and $|\omega-1|<1$, then
$\left|X_{m,n,r}(\omega)\right|=\left|Y_{m,n,r}(\omega)\right|<1.072\frac{r!\Gamma(1-\nu)}{\Gamma(r+1-\nu)}\left|1+\sqrt{\omega}\right|^{2r}.$
(e) For $|\omega|=1$ and $\operatorname{Re}(\omega)\geq 0$, we have
$\left|{}_{2}F_{1}\left(r+1-\nu,r+1;2r+2;1-\omega\right)\right|\geq 1,$
with the minimum value occurring at $\omega=1$.
Proof.
Part (a) is established in the proof of Lemma 2.3 of [5].
Part (b) is Lemma 4 of [3].
Part (c) is Lemma 2.5 of [5].
Part (d) is Lemma 4 of [17].
Part (e) is Lemma 5 of [17]. ∎
We let $D_{n,r}$ denote the smallest positive integer such that $D_{n,r}X_{m,n,r}(x)\in\mathbb{Z}[x]$ for all $m$ as above. For $d^{\prime}\in\mathbb{Z}$, we define $N_{d^{\prime},n,r}$ to be the largest integer such that $\left(D_{n,r}/N_{d^{\prime},n,r}\right)X_{m,n,r}\left(1-\sqrt{d^{\prime}}\,x\right)\in\mathbb{Z}\left[\sqrt{d^{\prime}}\right][x]$, again for all $m$ as above. We will use $v_{p}(x)$ to denote the largest power of a prime $p$ which divides into the rational number $x$. We put
(2.1) $\mathcal{N}_{d^{\prime},n}=\prod_{p|n}p^{\min(v_{p}(d^{\prime})/2,v_{p}(n)+1/(p-1))}.$
In what follows, we shall restrict our attention to $m=1$ and $n=4$, so $\nu=1/4$.
Lemma 2.3.
(a) We have
(2.2) $\frac{\Gamma(3/4)\,r!}{\Gamma(r+3/4)}\frac{D_{4,r}}{N_{d^{\prime},4,r}}<\mathcal{C}_{4,1}\left(\frac{\mathcal{D}_{4}}{\mathcal{N}_{d^{\prime},4}}\right)^{r}\text{ and }\hskip 2.84526pt\frac{\Gamma(r+5/4)}{\Gamma(1/4)r!}\frac{D_{4,r}}{N_{d^{\prime},4,r}}<\mathcal{C}_{4,2}\left(\frac{\mathcal{D}_{4}}{\mathcal{N}_{d^{\prime},4}}\right)^{r}$
for all non-negative integers $r$, where $\mathcal{C}_{4,1}=0.83$, $\mathcal{C}_{4,2}=0.2$ and $\mathcal{D}_{4}=e^{1.68}$.
(b) For any positive integer, $r$, we have
[…]
Proof.
(a) From Lemma 7.4(c) of [15], we have
$\max\left(1,\frac{\Gamma(3/4)\,r!}{\Gamma(r+3/4)},4\frac{\Gamma(r+5/4)}{\Gamma(1/4)r!}\right)\frac{D_{4,r}}{N_{d^{\prime},4,r}}<100\left(\frac{e^{1.64}}{\mathcal{N}_{d^{\prime},4}}\right)^{r}$
However, the value $100$ in this inequality results in us requiring a lot of computation to complete the proof of our results here (in particular, Proposition 4.1). Therefore, we seek a smaller value at the expense of replacing $1.64$ by a larger value, whose value has less of an impact on our proof. For $r\geq 156$, $100\exp(1.64r)<0.2\exp(1.68r)$, so we compute directly the left-hand sides of (2.2) for $r\leq 155$. We find that the maximum values of the left-hand sides of (2.2) divided by $\exp(1.68r)$ both occur for $r=3$. Part (a) follows.
(b) We prove both of these by induction. A quick calculation shows that we have equality in both cases for $r=1$. Then we take the ratio of the expression for $r+1$ and divide by the expression for $r$. For the first inequality, this ratio is $(1/4)\left(r+5/4\right)/\left(r+3/2\right)$, so the inequality will hold if $\left(r+5/4\right)/\left(r+3/2\right)>\left(r/\left(r+1\right)\right)^{1/4}$. We take the fourth-power of both sides and subtract them, this gives us $\left(224r^{3}+944r^{2}+1329r+625\right)/\left(16\left(2r+3\right)^{4}\left(r+1\right)\right)$, which is clearly positive for $r\geq 1$. Since both $\left(r+5/4\right)/\left(r+3/2\right)$ and $\left(r/\left(r+1\right)\right)^{1/4}$ are positive real numbers for $r\geq 1$, the required inequality hlds.
We proceed in the very same way to prove the second inequality. ∎
Put
$p_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega),$
$q_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}Y_{1,4,r}(\omega)\hskip 8.53581pt\text{ and}$
$R_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}R_{1,4,r}(\omega),$
where $r$ is any non-negative integer and $d^{\prime}$ will be determined below.
From Lemma 2.2(a), we have
$q_{r}\omega^{1/4}-p_{r}=R_{r}.$
By the definitions of $D_{4,r}$ and $N_{d^{\prime},4,r}$, we easily observe that $p_{r}^{\prime}$ and $q_{r}^{\prime}$ are algebraic integers of $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. We can see that as follows. $X_{1,4,r}(z)$ is a polynomial of degree $r$ and
$p_{r}^{\prime}$ $=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega)$
$=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
Since $d^{\prime}=u_{2}^{2}t^{\prime}$, $p_{r}^{\prime}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. The analogous expression for $q_{r}^{\prime}$ shows that it is also an algebraic integer.
In fact, there may be further common factors. As in [16, Theorem 1], put
(2.4) $g_{1}$ $=\gcd\left(u_{1},u_{2}\right),$
$g_{2}$ $=\gcd\left(u_{1}/g_{1},t^{\prime}\right),$
(2.8) $g_{3}$ =$=\left\{\begin{array}[]{ll}1&\text{if $t^{\prime}\equiv 1\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 2&\text{if $t^{\prime}\equiv 3\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 4&\text{otherwise,}\end{array}\right.$
$g$ $=g_{1}\sqrt{g_{2}/g_{3}}.$
Then we can put
(2.9) $p_{r}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega),$
$q_{r}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}Y_{1,4,r}(\omega)\hskip 8.53581pt\text{ and}$
$R_{r}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}R_{1,4,r}(\omega),$
where
(2.10) $d^{\prime}=\left(u-\sigma(u)\right)^{2}/g^{2}=u_{2}^{2}t^{\prime}/g^{2}.$
From Lemma 2.2(a), we have
$q_{r}\omega^{1/4}-p_{r}=R_{r}.$
Since $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/g$ is an algebraic integer, the argument above still applies to show that $p_{r}$ and $q_{r}$ are algebraic integers.
So by applying these quantities, with Lemma 2.2(d) and Lemma 2.3(a), in Lemma 2.1, we can take
(2.11) $Q=\frac{\mathcal{D}_{4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{|g|\mathcal{N}_{d^{\prime},4}}$
and
(2.12) $k_{0}<1.072\mathcal{C}_{4,1}<0.89.$
Using Lemma 2.2(c) instead of Lemma 2.2(d), we also have
(2.13) $E=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}$
and
$\ell_{0}=\mathcal{C}_{4,2}|\varphi|=0.2|\varphi|.$
3. Lemmas about $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$
3.1. Representation Proposition
The proposition in this section plays a crucial role in this work. It is the expression for $f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ in this lemma that permits us to use the hypergeometric method described in the previous section. For this reason, we call it our representation proposition.
It is also because we have $\varepsilon^{2}$, rather than $\varepsilon$, in (3.1) that we use $\varepsilon^{2k}$ in (1.1).
For any non-zero integer, $n$, we let $\operatorname{rad}(n)$ be the product of all distinct prime divisors of $n$. We will put $\operatorname{rad}(\pm 1)=1$.
Proposition 3.1.
Let $a\neq 0$,  [^2] PV: why do I want $a\neq 0$? $a=0$ will give me the squares among the $x_{k}$’s from $\varepsilon^{2k}$. $b>0$ and $d$ be rational integers such that $d$ is not a square. Put $\alpha=a+b^{2}\sqrt{d}$ and denote $N_{\mathbb{Q}(\sqrt{d})/\mathbb{Q}}(\alpha)$ by $N_{\alpha}$. Suppose that $N_{\alpha}$ is not a square, $x\neq 0$ and $y>0$ are rational integers with
(3.1) $x+y^{2}\sqrt{d}=\alpha\varepsilon^{2},$
where $\varepsilon=\left(t+u\sqrt{d}\right)/2\in\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ non-zero rational integers, norm $N_{\varepsilon}$ and $\gcd\left(u,N_{\varepsilon}\right)=1$.
(a) We can write
[…]
(3.2) $fy$ $=b\left(r^{2}-\operatorname{sf}\left(N_{\alpha}\right)s^{2}\right),$
for some integers $f$, $r$ and $s$ satisfying $f\neq 0$,
$f|\left(4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
where $f^{\prime}>0$, $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $f^{\prime}<\sqrt{\operatorname{sf}\left(\left|N_{\alpha}\right|\right)}$, if $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)>1$ and $f^{\prime}=1$ otherwise.
(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha},N_{\varepsilon}\right)\right)\right)$. However, to do so, the first relationship in part (a) may become
$-f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
(c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
Remark.
For work with binary recurrence sequences, $\varepsilon$ will be a power of an algebraic integer $\varepsilon^{\prime}$. So $\operatorname{rad}\left(N_{\varepsilon}\right)=\operatorname{rad}\left(N_{\varepsilon^{\prime}}\right)$. This makes our representation proposition useful for more general sequences than those considered in this paper.
The condition that $\gcd\left(u,N_{\varepsilon}\right)=1$ is no restriction. If $g=\gcd\left(u,N_{\varepsilon}\right)$, then $g|t$. By equation (3.1), $g^{2}|x$ and $g^{2}|y^{2}$.
To prove this proposition, we will use the following three lemmas.
Lemma 3.2.
Suppose that $a,b,d,t,u,x,y,\alpha$ and $\varepsilon$ are as in Proposition 3.1. Put $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}^{\prime}=-u\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$.
Then $\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right)$ is a divisor of $s_{1}^{\prime 2}$.
Proof.
Put $v_{p}(b)=k_{1}$ and $v_{p}\left(r_{1}\right)=k_{2}$ with $k_{1},k_{2}\geq 0$.
Suppose first that $p$ is an odd prime.
With $g_{1}^{2}=\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right)$, we have
(3.3) $v_{p}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2},&\text{if $k_{2}$ is even and $2k_{1}\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}\geq k_{2}$,}\\ 2k_{1}+k_{2}-1,&\text{if $k_{2}$ is odd and $2k_{1}-1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
From the definitions of $r_{1}$ and $k_{2}$, it follows that $v_{p}(au)\geq\min\left(k_{2},v_{p}\left(tb^{2}\pm 2by\right)\right)$.
Expanding the expression for $y^{2}$ in (3.1), we obtain
(3.4) $4y^{2}=2atu+b^{2}t^{2}+b^{2}du^{2}.$
Applying this equation, for odd primes, $p$, we have
$2v_{p}(y)$ $\geq\min\left(v_{p}(atu),v_{p}\left(b^{2}t^{2}\right),v_{p}\left(b^{2}du^{2}\right)\right)$
$\geq\min\left(v_{p}\left(t^{2}b^{2}\pm 2bty\right),k_{2}+v_{p}(t),v_{p}\left(b^{2}t^{2}\right),v_{p}\left(b^{2}du^{2}\right)\right)$
$\geq\min\left(v_{p}\left(bty\right),k_{2}+v_{p}(t),v_{p}\left(b^{2}t^{2}\right),v_{p}\left(b^{2}du^{2}\right)\right)$
$\geq\min\left(k_{1}+v_{p}(t)+v_{p}(y),k_{2}+v_{p}(t),2k_{1}+2v_{p}(t),2k_{1}+v_{p}(d)+2v_{p}(u)\right).$
This implies that $v_{p}(y)\geq\min\left(k_{1},k_{2}/2\right)$.
We consider separately the two possibilities for the gcd defining $g_{1}^{2}$ for each odd prime, $p$.
(1) Suppose that $v_{p}\left(r_{1}^{2}\right)\leq v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)=v_{p}\left(b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)$.
If $k_{2}$ is even, then we have $v_{p}\left(r_{1}^{2}\right)\leq v_{p}\left(b^{2}r_{1}\right)$. I.e., $k_{2}\leq 2k_{1}$. Similarly, if $k_{2}$ is odd, then we have $v_{p}\left(r_{1}^{2}\right)\leq v_{p}\left(b^{2}r_{1}\right)-1$. So $k_{2}\leq 2k_{1}-1$. In both cases, we have $k_{1}\geq k_{2}/2$ and thus $v_{p}(y)\geq k_{2}/2$.
(1-a) If $v_{p}(a)<2v_{p}(b)+v_{p}(d)/2$, then $v_{p}\left(N_{\alpha}\right)=v_{p}\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)=2v_{p}(a)$. So $v_{p}\left(s_{1}^{\prime 2}\right)=2v_{p}(au)$.
From $v_{p}(y)\geq k_{2}/2$ and $k_{1}\geq k_{2}/2$, it follows that $v_{p}\left(tb^{2}\pm 2by\right)\geq k_{2}$. So from the definition of $r_{1}$, we have $v_{p}(au)\geq k_{2}$. Hence $v_{p}\left(s_{1}^{\prime 2}\right)\geq 2k_{2}$. From (3.3), we have $v_{p}\left(g_{1}^{2}\right)=2k_{2}\leq v_{p}\left(s_{1}^{\prime 2}\right)$, as desired.
(1-b) If $v_{p}(a)\geq 2v_{p}(b)+v_{p}(d)/2$, then $v_{p}\left(N_{\alpha}\right)\geq v_{p}\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\geq 4v_{p}(b)$. So $v_{p}\left(s_{1}^{\prime 2}\right)\geq 4v_{p}(b)+2v_{p}(u)$.
From (3.3), we have $v_{p}\left(g_{1}^{2}\right)=2k_{2}$. Also $v_{p}\left(s_{1}^{\prime 2}\right)\geq 4v_{p}(b)+2v_{p}(u)\geq 2k_{2}+2v_{p}(u)$. So $v_{p}\left(g_{1}^{2}\right)\leq v_{p}\left(s_{1}^{\prime 2}\right)$.
Now suppose that $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$.
Put $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$. We shall prove the lemma in this case (i.e., $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$) for $r_{1}=r_{1}^{\prime}$. The proof is identical for $r_{1}=r_{1}^{\prime\prime}$.
If $k_{2}$ is even, then we have $2k_{1}<k_{2}$. If $k_{2}$ is odd, then we have $2k_{1}-1<k_{2}$. So in both cases, $2k_{1}\leq k_{2}$ (i.e., $v_{p}\left(4b^{2}\right)\leq v_{p}\left(r_{1}^{\prime 2}\right)$) and $v_{p}(y)\geq k_{1}$.
We now show that in this case we have $v_{p}\left(4b^{2}\right)<v_{p}\left(r_{1}^{\prime\prime}\right)$.
Since $v_{p}(y)\geq v_{p}(b)$, from the definition of $r_{1}^{\prime}$ and $v_{p}\left(r_{1}^{\prime}\right)>v_{p}\left(4b^{2}\right)$, we also have $v_{p}(au)>v_{p}\left(b^{2}\right)$. So the $p$-adic valuation of each term in the definition of $r_{1}^{\prime}$ is greater than $v_{p}\left(b^{2}\right)$. Thus $v_{p}\left(r_{1}^{\prime\prime}\right)>v_{p}\left(b^{2}\right)$ too.
Combining this with $r_{1}^{\prime}r_{1}^{\prime\prime}=u^{2}N_{\alpha}$, we have $v_{p}\left(b^{2}r_{1}^{\prime}\right)<v_{p}\left(r_{1}^{\prime}r_{1}^{\prime\prime}\right)=v_{p}\left(u^{2}N_{\alpha}\right)$, as desired.
The proof when $p=2$ is similar. But we will use the following information later in the proof of Proposition 3.1. We have
(3.5) $v_{2}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2}+2,&\text{if$k_{2}$ is even and $2k_{1}+2\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}+2\geq k_{2}$,}\\ 2k_{1}+k_{2}+1,&\text{if $k_{2}$ is odd and $2k_{1}+1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
As above, we have $v_{2}(au)\geq\min\left(k_{2},v_{2}\left(tb^{2}\pm 2by\right)\right)$.
Applying this, (3.4) and since $t^{2}+du^{2}=4N_{\varepsilon}+2du^{2}$, we obtain
$2\displaystyle 2v_{2}(y)+$ 2 $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
$\geq\min\left(v_{2}\left(2t^{2}b^{2}\pm 4bty\right),k_{2}+v_{p}(t)+1,v_{2}\left(2b^{2}\right)\right)$
$\geq\min\left(v_{2}\left(4bty\right),k_{2}+v_{2}(t)+1,2k_{1}+1\right)$
$\geq\min\left(k_{1}+v_{2}(t)+v_{2}(y)+2,k_{2}+v_{2}(t)+1,2k_{1}+1\right).$
Since $v_{2}(y)$ is an integer, this implies that
(3.6) $v_{2}(y)\geq\min\left(k_{1},\left(k_{2}-1\right)/2\right)$
and $v_{2}(y)\geq\min\left(k_{1},k_{2}/2\right)$, if $k_{2}$ is even. ∎
Lemma 3.3.
In the notation of Proposition 3.1, put $r_{1}^{\prime}=tb^{2}+au+2by$, $g_{1}^{\prime 2}=\gcd\left(4b^{2}r_{1}^{\prime}/\operatorname{sf}\left(r_{1}^{\prime}\right),r_{1}^{\prime 2}\right)$, $r_{1}^{\prime\prime}=tb^{2}+au-2by$ and $g_{1}^{\prime\prime 2}=\gcd\left(4b^{2}r_{1}^{\prime\prime}/\operatorname{sf}\left(r_{1}^{\prime\prime}\right),r_{1}^{\prime\prime 2}\right)$.
(a) For all primes, $p$,
$\max\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right).$
(b) For all primes, $p$, if $v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)\right)>0$, then
(3.7) $\min\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq 0\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
(c) If $r_{1}^{\prime}$ is even, then $v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<0\leq v_{2}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. The same result holds for $r_{1}^{\prime\prime}$.
Proof.
(a) We will show that for all primes, $p$, we have $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. From the definition of $\operatorname{rad}(\cdot)$, we have $v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\leq 1$.
From (3.11), we can write
(3.8) $\frac{r_{1}^{\prime}}{g_{1}^{\prime 2}}=\frac{\operatorname{sf}\left(r_{1}^{\prime}\right)}{\gcd\left(4b^{2},r_{1}^{\prime}\operatorname{sf}\left(r_{1}^{\prime}\right)\right)}.$
Since $v_{p}\left(\operatorname{sf}\left(r_{1}^{\prime}\right)\right)\leq 1$, for all primes, $p$, it follows from this expression that $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq 1$, for all primes, $p$. So it remains to show that if $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=1$, then we also have $v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)=1$.
From the expression above for $r_{1}^{\prime}/g_{1}^{\prime 2}$, we see that we are considering primes, $p$, such that $p|\operatorname{sf}\left(r_{1}^{\prime}\right)$ and $p\nmid\left(4b^{2}\right)$ (i.e., $p\nmid(2b)$).
From the observation that $r_{1}^{\prime}r_{1}^{\prime\prime}=N_{\alpha}u^{2}$, we have $r_{1}^{\prime}|\left(N_{\alpha}u^{2}\right)$.
We start by showing that if $v_{p}\left(r_{1}^{\prime}/g_{1}^{2}\right)=1$, then $p|\left(N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)$.
If $p|\operatorname{sf}\left(r_{1}^{\prime}\right)$, but $p\nmid\gcd\left(r_{1}^{\prime},r_{1}^{\prime\prime}\right)$, then $v_{p}\left(u^{2}N_{\alpha}\right)$ must be odd. Hence $p|\operatorname{sf}\left(N_{\alpha}\right)$.
So we need only consider what happens when $p|\operatorname{sf}\left(r_{1}^{\prime}\right)$, $p\nmid(2b)$ and $p|\gcd\left(r_{1}^{\prime},r_{1}^{\prime\prime}\right)$.
Suppose that $v_{p}\left(r_{1}^{\prime}\right)=k_{1}$ and $v_{p}\left(r_{1}^{\prime\prime}\right)=k_{2}$ for positive integers $k_{1}$ and $k_{2}$. Then $v_{p}\left(u^{2}N_{\alpha}\right)=k_{1}+k_{2}$. If $k_{1}+k_{2}$ is odd, then $p|\operatorname{sf}\left(N_{\alpha}\right)$.
We also know that $k_{1}$ is odd (otherwise, $p\nmid\operatorname{sf}\left(r_{1}^{\prime}\right)$). So we need only consider when $k_{2}$ is also odd.
Since $p|\left(r_{1}^{\prime}-r_{1}^{\prime\prime}\right)$, $\left(r_{1}^{\prime}-r_{1}^{\prime\prime}\right)=4by$ and $p\nmid(2b)$, it follows that $p|y$. We also have $p$ is a divisor of $r_{1}^{\prime}+r_{1}^{\prime\prime}=2au+2b^{2}t$.
We can write $4y^{2}=2atu+\left(t^{2}+du^{2}\right)b^{2}$ and find that $p$ is a divisor of $t\left(r_{1}^{\prime}+r_{1}^{\prime\prime}\right)-4y^{2}=b^{2}\left(t^{2}-du^{2}\right)$. Since $p\nmid b$, we have $p|N_{\varepsilon}$, completing the proof that $p|\left(N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)$.
Now we show that $p|\left(uN_{\alpha}\right)$ holds if $v_{p}\left(r_{1}^{\prime}/g_{1}^{2}\right)=1$. But this is immediate from $r_{1}^{\prime}r_{1}^{\prime\prime}=u^{2}N_{\alpha}$.
This completes the proof of part (a).
(b) Suppose $v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)\right)>0$. That is, $v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)\right)=1$. We will show that at least one of $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)$ or $v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)$ must be non-positive. Otherwise, by (3.8), $v_{p}\left(\operatorname{sf}\left(r_{1}^{\prime}\right)\right)=v_{p}\left(\operatorname{sf}\left(r_{1}^{\prime\prime}\right)\right)=1$. So $v_{p}\left(r_{1}^{\prime}r_{1}^{\prime\prime}\right)$ is even. But from $r_{1}^{\prime}r_{1}^{\prime\prime}=u^{2}N_{\alpha}$ and $v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)\right)=1$, we see that $v_{p}\left(r_{1}^{\prime}r_{1}^{\prime\prime}\right)$ is odd. Hence at least one of $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)$ or $v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)$ is non-positive, as required.
(c) This follows immediately from the expression on the right-hand side of (3.8). ∎
We shall need the following lemma for the proofs of parts (b) and (c) of Proposition 3.1.
Lemma 3.4.
With the same notation as in Lemma 3.3, $\gcd\left(r_{1}^{\prime},r_{1}^{\prime\prime}\right)$ is even, unless $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
Proof.
From the expressions for $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$, we need only show that $tb^{2}+au$ is even unless $N_{\alpha}\not\equiv 1\pmod{4}$ and $4\nmid d$. We do so by showing that if either (i) $tb^{2}$ even and $au$ odd or (ii) $tb^{2}$ odd and $au$ even, then $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
(i) We first suppose that $t$ is even in case (i). Since $u$ is odd and $\left(t^{2}-du^{2}\right)/4\in\mathbb{Z}$, we have $4|d$. But then $N_{\alpha}=a^{2}-db^{4}\equiv 1\pmod{4}$, since $a$ must be odd.
Now suppose that $t$ is odd. Since $\left(t^{2}-du^{2}\right)/4\in\mathbb{Z}$ and $u$ is odd, it follows that $d\equiv 1\pmod{4}$. Here $b$ must be even, since $tb^{2}$ is even. Again $N_{\alpha}=a^{2}-db^{4}\equiv 1\pmod{4}$, since $a$ must be odd. Applying these to the right-hand side of (3.4), we find that $4y^{2}\equiv 2\pmod{4}$, so this case is not possible.
(ii) If $u$ is even, then since $t$ is odd, $N_{\varepsilon}$ cannot be an integer. So we must have $a$ even and $b$, $t$ and $u$ all odd. Since $t$ and $u$ are both odd, $d\equiv 1\pmod{4}$. So we have $N_{\alpha}\equiv 3\pmod{4}$. Hence $\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}\equiv 2\pmod{4}$. But this quantity is $b^{2}$ times the right-hand side of (3.4), so it is $4b^{2}y^{2}$. Thus (ii) is not possible. ∎
Proof of Proposition 3.1.
(a) We start by writing $\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(a+\sqrt{N_{\alpha}}\right)$ as a square.
By rationalising the denominator of $\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(a+\sqrt{N_{\alpha}}\right)$ and then applying the expressions for $N_{\varepsilon}$ and $N_{\alpha}$, we find that
$\frac{x+N_{\varepsilon}\sqrt{N_{\alpha}}}{a+\sqrt{N_{\alpha}}}=\frac{4ax-a^{2}t^{2}+b^{4}dt^{2}+a^{2}du^{2}-b^{4}d^{2}u^{2}+\left(at^{2}-adu^{2}-4x\right)\sqrt{N_{\alpha}}}{4b^{4}d}.$
Substituting
(3.9) $4x=a\left(t^{2}+du^{2}\right)+2b^{2}dtu,$
from the expression for $x+y^{2}\sqrt{d}$ in (3.1), we obtain
$\frac{x+N_{\varepsilon}\sqrt{N_{\alpha}}}{a+\sqrt{N_{\alpha}}}=\frac{2a^{2}u^{2}+2ab^{2}tu+b^{4}t^{2}-db^{4}u^{2}-2u\left(au+b^{2}t\right)\sqrt{N_{\alpha}}}{4b^{4}}.$
We can write $2a^{2}u^{2}+2ab^{2}tu+b^{4}t^{2}-db^{4}u^{2}=\left(au+b^{2}t\right)^{2}+u^{2}N_{\alpha}$, so
$\frac{x+N_{\varepsilon}\sqrt{N_{\alpha}}}{a+\sqrt{N_{\alpha}}}=\left(\frac{\left(au+b^{2}t\right)-u\sqrt{N_{\alpha}}}{2b^{2}}\right)^{2}.$
Applying (3.4), a routine calculation shows that with $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}=-u$, we have
$\left(r_{1}+s_{1}\sqrt{N_{\alpha}}\right)^{2}=2r_{1}\left(\left(au+b^{2}t\right)-u\sqrt{N_{\alpha}}\right).$
Hence
$\left(4b^{2}r_{1}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{1}+s_{1}\sqrt{N_{\alpha}}\right)^{4}$
(3.10) $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{1}+s_{1}^{\prime}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
where $s_{1}^{\prime}=s_{1}\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$.
From this relationship, we now show that we can find such a relationship for $x+N_{\varepsilon}\sqrt{N_{\alpha}}$ with the conditions on $f$ in part (a) satisfied.
We want to take the largest common factor, $g_{1}$, of $r_{1}$ and $s_{1}^{\prime}$ such that $4b^{2}r_{1}/g_{1}^{2}$ is also an integer. We have
$g_{1}^{2}=\gcd\left(\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right),s_{1}^{\prime 2}\right).$
First, note that we can write
(3.11) $\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right)=r_{1}\gcd\left(4b^{2},r_{1}\operatorname{sf}\left(r_{1}\right)\right)/\operatorname{sf}\left(r_{1}\right).$
From Lemma 3.2, we see that
$g_{1}^{2}=\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right).$
From Lemma 3.3(a), $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of $4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$. We put $r=r_{1}/g_{1}$, for either choice of $r_{1}$, $s=s_{1}^{\prime}/g_{1}$ and $f=4b^{2}r_{1}/g_{1}^{2}$. By the definition of $g_{1}^{2}$, we see that $f,r,s\in\mathbb{Z}$.
Lastly, we choose $r_{1}$. Let $d^{\prime}$ be the product of all primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$, such that the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ and $d^{\prime\prime}$ be the product of the remaining primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$ (so the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime\prime}/{g_{1}^{\prime\prime}}^{2}$, but not for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ for such primes). If $d^{\prime}>d^{\prime\prime}$, then we let $r_{1}=r_{1}^{\prime}$ and let $r_{1}=r_{1}^{\prime\prime}$ otherwise. This choice ensures that $f^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)/\max\left(d^{\prime},d^{\prime\prime}\right)$ satisfies the desired conditions.
(b) If $f$ is odd, then there is nothing to prove, as $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$ follows from $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. So we need only consider when $f$ is even.
We consider the four cases for the parity of $r$ and $s$.
(b-i) Suppose that $r$ and $s$ are both odd.
Here we will replace $r$ and $s$.
We can write $(r+si)/(1+i)=((r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
$-\left(\frac{2b^{2}r}{g_{1}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
Since $f=4b^{2}r/g_{1}$ is even, we know that $2b^{2}r/g_{1}\in\mathbb{Z}$.
Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$. Hence $2b^{2}r/g_{1}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$.
(b-ii) Suppose that $r=r_{1}/g_{1}$ is odd and $s=s_{1}^{\prime}/g_{1}$ is even.
Since $r$ is odd, we have $v_{2}\left(g_{1}\right)=v_{2}\left(r_{1}\right)$. So from the definition of $g_{1}^{2}$ in the proof of part (a), it follows that $v_{2}\left(4b^{2}/\operatorname{sf}\left(r_{1}\right)\right)\geq v_{2}\left(r_{1}\right)$.
For this case and case (b-iii), if $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, then we must have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$, since $r_{1}^{\prime\prime}=r_{1}^{\prime}+4by$.
(b-ii-1) Suppose first that $v_{2}\left(r_{1}\right)$ is even. In this case, $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$. If we have equality, then $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and so $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required. So we consider $v_{2}\left(4b^{2}\right)>v_{2}\left(r_{1}\right)$.
Since both of these quantities are even, we must have $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)+2$. That is, $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$.
Since $s$ is even, we have $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)>2v_{2}\left(g_{1}\right)=2v_{2}\left(r_{1}^{\prime}\right)$. So $v_{2}\left(r_{1}^{\prime\prime}\right)>v_{2}\left(r_{1}^{\prime}\right)$. As we stated near the start of this case ((b-ii)), since $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
We saw in the proof of Lemma 3.2 that $v_{2}(y)\geq\min\left(v_{2}(b),v_{2}\left(r_{1}^{\prime}\right)/2\right)$. Since $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$, it follows that $v_{2}(y)\geq v_{2}\left(r_{1}^{\prime}\right)/2$. Hence $v_{2}(2by)\geq 1+v_{2}\left(r_{1}\right)/2+v_{2}\left(r_{1}\right)/2=v_{2}\left(r_{1}\right)+1$. Hence $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)\geq v_{2}\left(r_{1}\right)+1$, which implies that $v_{2}\left(r_{1}\right)>v_{2}\left(r_{1}\right)$. Therefore, the case when $v_{2}\left(r_{1}\right)$ is even and $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$ never occurs and we find that in case (b-ii-1), $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
(b-ii-2) Now suppose that $v_{2}\left(r_{1}\right)$ is odd. Here $v_{2}\left(2b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
As in case (b-ii-1), since $s$ is even, we have $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$. Hence $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)$.
Recall from (3.6) that $v_{2}(y)\geq\min\left(v_{2}(b),\left(v_{2}\left(r_{1}\right)-1\right)/2\right)$. Hence, $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)\geq v_{2}\left(r_{1}\right)$. Since $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)$, we have $v_{2}\left(r_{1}\right)\geq v_{2}\left(2by\right)+1>v_{2}\left(r_{1}\right)$, which is not possible.
This completes the proof for case (b-ii).
(b-iii) Suppose that $r=r_{1}/g_{1}$ is even and $s=s_{1}^{\prime}/g_{1}$ is odd.
Since $r$ is even, we must have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{2}\left(r_{1}^{2}\right)$.
Since $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(g_{1}^{2}\right)<2v_{2}\left(r_{1}\right)$. Hence $v_{2}\left(r_{1}^{\prime\prime}\right)\neq v_{2}\left(r_{1}\right)$, so by the comment at the start of case (b-ii), we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
(b-iii-1) If $v_{2}\left(r_{1}\right)$ is even, then $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}\right)$. Hence $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
(b-iii-2) If $v_{2}\left(r_{1}\right)$ is odd, then $v_{2}\left(2b^{2}r_{1}\right)<v_{2}\left(r_{1}^{2}\right)$. So $v_{2}(b)<\left(v_{2}\left(r_{1}\right)-1\right)/2$. Recall from (3.6) that $v_{2}(y)\geq\min\left(v_{2}(b),\left(v_{2}\left(r_{1}\right)-1\right)/2\right)$. So $v_{2}(y)\geq v_{2}(b)$.
Hence, $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)\geq 2v_{2}(b)+1$. Since $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)$, we have $v_{2}\left(r_{1}\right)\geq v_{2}\left(2by\right)+1\geq 2v_{2}(b)+2$. Using the same argument, we have $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+2$ too.
But $v_{2}\left(2b^{2}r_{1}\right)=v_{2}\left(g_{1}^{2}\right)=v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)$. Hence $v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(2b^{2}\right)$. But this contradicts $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+2$. Hence $v_{2}\left(r_{1}\right)$ cannot be odd.
(b-iv) Suppose that $r$ and $s$ are both even.
Since $r$ is even, we know that $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{2}\left(r_{1}^{2}\right)$. So $v_{2}(f)=0$ or $1$, depending on whether $v_{2}\left(r_{1}\right)$ is even or odd. So, as stated at the start of the proof of part (b), we need only consider the latter case.
Since $s$ is even, we have $2\leq v_{2}\left(s^{2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)-v_{2}\left(2b^{2}\right)$. So $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+3$.
Similarly, $2\leq v_{2}\left(r_{1}^{\prime 2}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime}\right)-v_{2}\left(2b^{2}\right)$, so $v_{2}\left(r_{1}^{\prime}\right)\geq 2v_{2}(b)+3$.
A consequence of these two inequalities is that $64|\left(u^{2}N_{\alpha}\right)$.
From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4bN_{\varepsilon}$. So $atu+du^{2}$ must be even. I.e., $atu$ and $du^{2}$ have the same parity.
If $du^{2}$ is odd, then $a$ and $t$ are also odd. Since $4N_{\varepsilon}$ is divisible by $4$, it must be the case that $d\equiv 1\pmod{4}$. If $d\equiv 1\pmod{8}$, then $N_{\varepsilon}$ is even. Furthermore, since $abd$ is odd, $N_{\alpha}$ is even, so $v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\geq 1$ and the desired conclusion holds.
If $d\equiv 5\pmod{8}$, then $N_{\alpha}=a^{2}-db^{4}\equiv 4\pmod{8}$. Since $u$ is odd, we get a contradiction with $64|\left(u^{2}N_{\alpha}\right)$.
We now consider $du^{2}$ even. Here we will prove that both $uN_{\alpha}$ and $N_{\varepsilon}$ are even. This suffices to prove part (b) in this case.
To prove that $N_{\varepsilon}$ is even, we will assume that it is odd and obtain a contradiction.
Since $4N_{\varepsilon}=t^{2}-du^{2}\in\mathbb{Z}$ and $du^{2}$, we have $t$ is even and hence $4|\left(du^{2}\right)$.
If $u$ is odd, then $4|d$. Also $N_{\alpha}$ is even, so $a$ is even. As a result we find that $v_{2}(2atu),v_{2}\left(2b^{2}du^{2}\right)\geq 3$. So under our assumption that $N_{\varepsilon}$ is odd, we have $v_{2}\left(4y^{2}\right)=2$. I.e., $y$ is odd.
The same reasoning shows that $y$ is odd if $u$ is even under the assumption that $N_{\varepsilon}$ is odd.
Suppose that $v_{2}\left(r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)$. Then $v_{2}\left(4by\right)=v_{2}\left(\pm\left(r_{1}^{\prime}-r_{1}^{\prime\prime}\right)\right)>v_{2}\left(r_{1}^{\prime}\right)$. So $v_{2}(y)>v_{2}\left(r_{1}^{\prime}\right)-v_{2}(b)-2>v_{2}(b)+1$.
Suppose that $v_{2}\left(r_{1}^{\prime}\right)<v_{2}\left(r_{1}^{\prime\prime}\right)$. Then $v_{2}\left(4by\right)=v_{2}\left(r_{1}^{\prime}\right)$. So $v_{2}(y)\geq v_{2}(b)+1$.
In both cases, we find that $y$ is even. This contradicts the assumption that $N_{\varepsilon}$ is odd. Hence we know that in the case when $du^{2}$ is even, $N_{\varepsilon}$ is even and the assertion in part (b) of the lemma holds.
(c) We consider two cases.
(c-i) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=p$, for an odd prime, $p$. Then we have $f^{\prime}=1$ in part (a) and so $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$. Hence the result holds if $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
From Lemma 3.4, $r$ is even unless $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$. So, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$. With $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$.
(c-ii) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=2p$, where $p$ is an odd prime. Here the worst case is when the minimum in (3.7) occurs for one of $r_{1}^{\prime}/g_{1}^{\prime 2}$ and $r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}$ for $2$ and occurs for the other for $p$. In this case, we let $r$ be the one of $r_{1}^{\prime}$ or $r_{1}^{\prime\prime}$ for which the minimum in (3.7) for $p$ occurs. Thus $v_{p}\left(r/g_{1}^{2}\right)\leq v_{p}\left(\operatorname{rad}\left(N_{\varepsilon}\right)\right)$ for all odd primes, $p$. From Lemma 3.4, we see that $r$ is even. Hence, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$. So $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$. ∎
3.2. Lower bounds for $y_{k}$’s
Next we bound the $y_{k}$’s from below.
Lemma 3.5.
[…]
(a) Put $\overline{\alpha}=a-b^{2}\sqrt{d}$. We have
(3.12) $y_{k}>\left\{\begin{array}[]{ll}\dfrac{\alpha\varepsilon^{2k}}{2\sqrt{d}}&\text{for $k\geq 0$,}\\ \dfrac{-\overline{\alpha}\,\varepsilon^{2|k|}}{2\sqrt{d}}&\text{for $k<0$.}\end{array}\right.$
(b) For all $k$, $2y_{k}$ is a positive integer. The sequences $\left(y_{k}\right)_{k\geq 0}$ and $\left(y_{k}\right)_{k\leq K+1}$ are increasing sequences of positive numbers.
(c) We have
(3.13) $y_{k}>\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{K-k}&\text{for $k\leq K$.}\end{array}\right.$
In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $52du^{2}/$ 5 by $85du^{2}/$ 8 and if $N_{\varepsilon}=1$, then we can replace $52du^{2}/$ 5 by $du^{2}$.
Remark.
If $N_{\alpha}>0$, then $y_{k}<0$ can occur for $k<0$.
Proof.
(a) From (1.1), we can write
$y_{k}=\frac{\alpha\varepsilon^{2k}-\overline{\alpha}\,\overline{\varepsilon}^{2k}}{2\sqrt{d}},$
where $\overline{\varepsilon}=\left(t-u\sqrt{d}\right)/2$. Since $N_{\alpha}=\alpha\overline{\alpha}<0$ and $\alpha>0$, we have $\overline{\alpha}<0$. So the inequality (3.12) follows.
(b) From part (a), it follows that all the $y_{k}$’s are positive. Since $\alpha\varepsilon^{2k}$ is an algebraic integer, we have $2y_{k}\in\mathbb{Z}$.
Since $d$, $t^{2}$ and $u^{2}$ are all positive integers, a quick search over small values of $d$, $t$ and $u$ with $t^{2}-du^{2}=\pm 4$ shows that $t^{2}+du^{2}\geq 6$ (the minimum occurs for $(d,t,u)=(5,\pm 1,\pm 1)$). From this, (1.3) and $y_{k}>0$, we have $y_{k+1}\geq 3y_{k}-y_{k-1}$ for $k\geq 1$.
From the expression for $y_{1}$ after (1.1) and $t^{2}+du^{2}\geq 6$, we have $y_{1}>(3/2)b^{2}>b^{2}=y_{0}$. So using induction and $y_{k+1}\geq 3y_{k}-y_{k-1}$, we find that $y_{k+1}>y_{k}$ for all $k\geq 0$.
By the definition of $K$, $y_{K}>y_{K+1}$. Also $y_{K}>b^{2}>0$, so we can proceed as in the case of $k\geq 0$, using $y_{k-1}\geq 3y_{k}-y_{k+1}$ and induction.
(c) From (1.2), we find that
(3.14) $4b^{2}y_{1}=b^{4}\left(t^{2}+du^{2}\right)+2ab^{2}tu=\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}.$
From the second equality in (3.14), equation (3.13) follows for $k=1$.
Writing $\varepsilon^{2k}=t_{k}+u_{k}\sqrt{d}$ for $k\geq 1$, we see that $\left(u_{k}\right)_{k\geq 1}$ is an increasing sequence. Hence
$4b^{2}y_{|K|}$ $=b^{4}\left(t_{|K|}^{2}+du_{|K|}^{2}\right)-2ab^{2}t_{|K|}u_{|K|}=\left(b^{2}t_{|K|}-au_{|K|}\right)^{2}-N_{\alpha}u_{|K|}^{2}$
$\geq\left|N_{\alpha}\right|u^{2}.$
So equation (3.13) holds for $k=K$.
Recall the recurrence in (1.3) for $k\geq 0$ and using the monotonicity established in part (b):
$y_{k+1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k-1}\geq\left(du^{2}+2N_{\varepsilon}-1\right)y_{k}.$
If $N_{\varepsilon}=1$, then the stronger form of the desired inequality holds, so we need only consider $N_{\varepsilon}=-1$. We always have $du^{2}\geq 5$ (with equality only if $(d,t,u)=(5,1,1)$. Otherwise, $du^{2}\geq 8$). Hence the desired inequality holds.
This is not too far from the best possible result from this argument. For $(a,d,t,u)=(1,5,1,1)$ and $b$ arbitrarily large, $y_{2}/y_{1}$ approaches $7/3$ from above, whereas we use $2$ here.
Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain the desired result for $k\leq K$ too. ∎
3.3. Gap Principle
In Lemma 3.7 below, we establish a gap principle separating distinct squares in the sequence (1.1). We first need the following technical lemma to help us prove our gap principle.
Lemma 3.6.
Let $\omega=e^{i\theta}$ with $-\pi<\theta\leq\pi$.
(a) Put $\omega^{1/4}=e^{i\theta/4}$. If
$0<\left|\omega^{1/4}-z\right|<c_{1},$
for some $z\in\mathbb{C}$ with $|z|=1$ and $0<c_{1}<\sqrt{2}$, then
$\left|\omega-z^{4}\right|>c_{2}\left|\omega^{1/4}-z\right|,$
where $c_{2}=\left(2-c_{1}^{2}\right)\sqrt{4-c_{1}^{2}}$.
(b) If
$0<\left|\omega-z^{4}\right|<c_{0},$
for some $z\in\mathbb{C}$ with $|z|=1$ and $0<c_{0}\leq 2$, then
$0<\left|\omega^{1/4}-z\right|<c_{3},$
for some choice of $\omega^{1/4}$, where $c_{3}$ is the smallest positive real root of $x^{8}-8x^{6}+20x^{4}-16x^{2}+c_{0}^{2}$.
Proof.
(a) We can write
$\left|\omega-z^{4}\right|=\left|\omega^{1/4}-z\right|\times\prod_{k=1}^{3}\left|\omega^{1/4}-e^{2\pi ik/4}z\right|.$
Multiplying by $\omega^{-1/4}$ and expanding the resulting expression, the product above equals
$\prod_{k=1}^{3}\left|e^{2\pi ik/4-i\theta/4}z-1\right|=\left|e^{3i\varphi}+e^{2i\varphi}+e^{i\varphi}+1\right|,$
for $-\pi<\varphi\leq\pi$ such that $e^{i\varphi}=e^{-i\theta/4}z$. Squaring this quantity and simplifying, we obtain
(3.15) $8\cos^{2}(\varphi)\left(\cos(\varphi)+1\right).$
If $\left|\omega^{1/4}-z\right|=\left|1-\omega^{-1/4}z\right|=c_{1}$, we have $2-2\cos(\varphi)=c_{1}^{2}$ and, substituting this expression for $\cos(\varphi)$ into (3.15), we find that $\left|\omega-z^{4}\right|=c_{2}\left|\omega^{1/4}-z\right|$. Since $c_{2}$ is a decreasing function of $c_{1}$, part (a) of the lemma holds (i.e., if $\left|\omega^{1/4}-z\right|<c_{1}$).
(b) Using the same argument as in part (a), but supposing that $\left|\omega-z^{4}\right|=\left|1-\omega^{-1}z^{4}\right|=c_{0}$, then we have $2-2\cos(4\varphi)=c_{0}^{2}$. Thus
$\left|\omega^{1/4}-z\right|^{2}$ $=\left(\frac{\left|\omega-z^{4}\right|}{\left|e^{3i\varphi}+e^{2i\varphi}+e^{i\varphi}+1\right|}\right)^{2}=\frac{c_{0}^{2}}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}$
$=\frac{2-\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$. Part (b) now follows and with inequalities since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$. ∎
Lemma 3.7.
[…]
(a) Suppose that $-N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are both squares with $y_{j}>y_{i}\geq 4b^{6}$, then
$y_{j}>\frac{57.6d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
(b) Suppose $-N_{\alpha}$ is not a square. If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4b^{6}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
$y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3},$
where $f_{i}$ and $f_{j}$ are the values of $f$ in Proposition 3.1 associated with $y_{i}$ and $y_{j}$, respectively.
If $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then we can replace $15.36$ with $182$.
Proof.
We start by considering the two parts together and what is common to their proofs.
By Proposition 3.1, there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
±$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
(3.16) $f_{j}\sqrt{y_{j}}$ $=b\left(r_{j}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{j}^{2}\right),$
where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$.
For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $y_{i},y_{j}>1$, we will assume that the $\pm$ on the left-hand side is always $+$. That is,
(3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
$f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
as the argument for the other cases is exactly the same. Subtracting the complex conjugate of one of these equations from the equation itself, we obtain
$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
and the analogous equation with the index $j$ replaced by $i$ also holds.
Putting $\omega=\left(a-\sqrt{N_{\alpha}}\right)/\left(a+\sqrt{N_{\alpha}}\right)$, by (3.3) and the relationship for $f_{j}\sqrt{y_{j}}$ in Proposition 3.1(a), we have
(3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ =|±$=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
$=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}<\frac{2b^{2}}{y_{j}}\leq\frac{1}{2},$
the last inequality holding because $y_{j}\geq 4b^{6}$.
Let $\zeta_{4}^{(j)}$ be the $4$-th root of unity such that
(3.20) $\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq k\leq 3}\left|\omega^{1/4}-e^{2k\pi i/4}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
From (3.19), we can apply Lemma 3.6(b) with $c_{0}=1/2$ to obtain
$\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.1263.$
Applying Lemma 3.6(a) with $c_{1}=0.1263$, we obtain
$\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.96\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
and combining this with the equalities in (3.19) yields
(3.21) $\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.5051\frac{b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}}\,\frac{1}{y_{j}}.$
In the same way, we define $\zeta_{4}^{(i)}$ for any square $y_{i}>1$ in our sequence and (3.21) also holds with $j$ replaced by $i$ for such $i$. Hence, for any two distinct squares, $y_{i},y_{j}\geq 4b^{2}$, among the $y_{k}$’s, we have
$\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$\leq$ $\left|\omega^{1/4}-\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|+\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$<$ $0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\left(\frac{1}{y_{i}}+\frac{1}{y_{j}}\right).$
Next we obtain a lower bound for this same quantity. We first show that it cannot be zero.
If
$\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}=\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}},$
then from our expressions for $f_{i}\sqrt{y_{i}}$ and $f_{j}\sqrt{y_{j}}$ from Proposition 3.1 we have
$\frac{\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}{f_{i}^{2}y_{i}}=\pm\frac{\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}{f_{j}^{2}y_{j}}.$
From (3.17), it follows that
$\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{j}=\pm\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{i}.$
Comparing the imaginary parts of both sides of this equation, we find that $y_{i}=y_{j}$, but this contradicts our assumption that $y_{j}>y_{i}$. Hence the left-hand side of (3.3) cannot be 00.
Let $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$. We can write
(3.23) $\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}$
$=$ $\frac{2\zeta_{4}^{(i)}x-\left(\zeta_{4}^{(i)}+\zeta_{4}^{(j)}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}.$
The numerator on the right-hand side is
(3.24) $\begin{array}[]{rl}-2\zeta_{4}^{(i)}y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}&\text{if $\zeta_{4}^{(j)}=\zeta_{4}^{(i)}$,}\\ 2\zeta_{4}^{(i)}x&\text{if $\zeta_{4}^{(j)}=-\zeta_{4}^{(i)}$,}\\ \zeta_{4}^{(i)}\left(1-\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=\sqrt{-1}\zeta_{4}^{(i)}$ and}\\ \zeta_{4}^{(i)}\left(1+\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=-\sqrt{-1}\zeta_{4}^{(i)}$.}\end{array}$
Here we use $\sqrt{-1}$ to denote $\exp\left(2\pi i/4\right)$.
Using (3.2), we find that
(3.25) $\left|r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|\left|r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|=\frac{\sqrt{f_{i}f_{j}}\left(y_{i}y_{j}\right)^{1/4}}{b}.$
At this point, our proofs of the two parts of the lemma separate.
(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that $1+\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
$\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$\geq$ $\frac{|1+\sqrt{-1}|}{\left|\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\right|}=\frac{\sqrt{2}\,b}{\sqrt{f_{i}f_{j}}\left(y_{i}y_{j}\right)^{1/4}},$
using (3.25) above.
Furthermore, from Proposition 3.1(b), we know that $f_{i},f_{j}|b^{2}$, since $-N_{\alpha}$ is a square. So
$\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\geq\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}.$
Combining this with (3.3), we have
(3.26) $\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\left(\frac{1}{y_{i}}+\frac{1}{y_{j}}\right).$
From $y_{j}>y_{i}$, it follows that $1/y_{i}+1/y_{j}<2/y_{i}$ and so
$y_{j}>\frac{4}{1.0102^{4}b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
Since $N_{\alpha}=a^{2}-b^{4}d<0$, we have
(3.27) $y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3},$
We can use this gap principle to improve its constant term. Combining (3.27) with $y_{i}\geq 4b^{6}>4b^{2}\left|N_{\alpha}\right|/d$ and $\left|N_{\alpha}\right|<b^{4}d$, we obtain
$y_{j}>\frac{3.84}{b^{12}}y_{i}^{3}\geq 61.44y_{i},$
Applying this to (3.26) yields
$\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\frac{1+1/61.44}{y_{i}}.$
This implies that
$y_{j}>\frac{57.61}{b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
completing our proof of part (a).
(b) We again use (3.23) and (3.24). Notice that $x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}$ on the left-hand side of (3.24) for $\zeta_{4}^{(j)}=\pm\sqrt{-1}\zeta_{4}^{(i)}$ can be as small as $1/\left(2\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}y\right)$ when $-N_{\alpha}$ is not a square, so the proof of part (a) breaks down here.
Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other. Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $i<j$.
Thus,
$\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\geq\frac{2b}{\sqrt{f_{i}f_{j}}\left(y_{i}y_{j}\right)^{1/4}}.$
The argument is the same as in the proof of part (a) except we have an extra factor of $\sqrt{2}$ on the right-hand side of this inequality. Thus
(3.28) $\frac{2b}{\sqrt{f_{i}f_{j}}\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\left(\frac{1}{y_{i}}+\frac{1}{y_{j}}\right).$
As in the proof of part (a), this gives
$y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
If $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, then this lower bound for $y_{j}$ yields $y_{j}>1.09y_{i}$. Applying this to (3.28), we obtain
$y_{j}>18.25\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
Repeating this process eight more times yields
$y_{j}>182\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
Since $f_{i}f_{j}\leq\left(4b^{2}\right)^{2}\left|N_{\alpha}\right|\leq 16b^{8}d$, if $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, as required. ∎
3.4. Miscellaneous Lemmas
Here we collect some results that we will need for bounding quantities that arise in the proof of the main result in the following section (Section 4).
Lemma 3.8.
(a) Let $y_{k}\geq 4b^{2}$ be a square and put
$\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)=e^{i\varphi}$
with $-\pi<\varphi\leq\pi$. Then
$\left|\varphi\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4b^{2}$, $\omega$ be as in part (a) and put $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1. Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.7 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$. Then
$\min_{0\leq j\leq 3}\left|\omega^{1/4}-\zeta_{4}^{j}\frac{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
occurs for either $j=0$ or $j=2$.
Proof.
(a) We can write $\omega=\left(\left|x_{k}\right|\pm\sqrt{N_{\alpha}}\right)^{2}/\left(x_{k}^{2}+\left|N_{\alpha}\right|\right)$, so with $\omega=e^{i\varphi}$, we have $\left|\tan\left(\varphi\right)\right|=\left|2x_{k}\sqrt{\left|N_{\alpha}\right|}/\left(x_{k}^{2}+N_{\alpha}\right)\right|$. From (1.1) and $N_{\alpha}>-b^{4}d$, we have
$x_{k}^{2}+N_{\alpha}$ $=dy_{k}^{2}+2N_{\alpha}>d\left(y_{k}^{2}-2b^{4}\right)\geq dy_{k}^{2}\left(1-\frac{2b^{4}}{y_{k}^{2}}\right)$
$\geq 0.875dy_{k}^{2}>0.875x_{k}^{2},$
since $y_{k}\geq 4b^{2}$ and hence $1-2b^{4}/y_{k}^{2}\geq 1-2b^{4}/\left(16b^{12}\right)\geq 0.875$.
From $|\varphi|\leq|\tan(\varphi)|$, we have
$|\varphi|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
Since $y_{k}\geq 4b^{2}$, we have $dy_{k}^{2}\geq 16db^{4}>16\left|N_{\alpha}\right|$ and so
$x_{k}^{2}=dy_{k}^{2}+N_{\alpha}>15\left|N_{\alpha}\right|,$
which yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
(b) Recall that from (3.3) in the proof of Lemma 3.7, we can write
$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{N_{\alpha}}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{N_{\alpha}}\right)^{4}=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
for $i=k,\ell$. There are other cases according to the signs in the result in Proposition 3.1, but the argument for them is identical to the argument that follows below.
Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)}{f_{k}^{2}}$
$f_{k}^{2}y_{k}^{2}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)$.
Applying equation (3.3) with $j=\ell$ and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
(3.29) $|f(x,y)|$ $=\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
$=2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}\sqrt{\left|N_{\alpha}\right|}.$
Let $\zeta_{4}$ be the $4$-th root of unity satisfying
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (3.1) (which implies that $\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ $=\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
$=\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}<\frac{2b^{2}}{y_{\ell}}\leq\frac{1}{2}.$
since $y_{\ell}\geq 4b^{2}$.
By Lemma 3.6(b) with $c_{0}=1/2$,
(3.31) $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
From (3.3), along with $y_{k},y_{\ell}\geq 4b^{2}$, we have
$\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|$ $=\left|\zeta_{4}^{(k)}\frac{r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(\ell)}\frac{r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$<0.5051b^{2}\frac{\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}}\left(\frac{1}{y_{k}}+\frac{1}{y_{\ell}}\right)<0.253.$
From part (a), we have $|\varphi|<0.6$, so $\left|\omega^{1/4}-1\right|<0.15$ and
$\left|\omega^{1/4}-\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\leq\left|\omega^{1/4}-1\right|+\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|<0.403.$
Recalling (3.31), it follows that
$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|=\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
$\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
for $\zeta_{4}^{\prime}\neq\zeta_{4}$. Since this exceeds $0.403$, it follows that $\zeta_{4}=\zeta_{4}^{(k)}/\zeta_{4}^{(\ell)}=\pm 1$, the last equality holding by our assumption in the statement of this lemma. ∎
In Lemma 3.9, we establish a gcd result for elements of a generalisation of our sequences. This will help us prove Lemmas 3.10 and 3.11. Lemma 3.11 is used in the proof of Proposition 4.1, as well as in the proof of Theorem 1.4. Lemma 3.11 will be particularly important for showing that the possible exceptions to Conjecture 1.2 in the statement of Theorem 1.4 can only possibly occur for $u=1$ or $u=2$.
Lemma 3.9.
Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon\neq\pm 1$ is a unit in $\mathbb{Q}\left(\sqrt{d}\right)$.
If $x_{k}$ and $y_{k}$ are both integers, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
Proof.
We write $\varepsilon=\left(t+u\sqrt{d}\right)/2$, where $t$ and $u$ are integers.
We may assume without loss of generality that $\gcd(a,b)=1$. Otherwise, consider below the sequences $\left(x_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ and $\left(y_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ instead of $\left(x_{k}\right)_{k=0}^{k=\infty}$ and $\left(y_{k}\right)_{k=0}^{k=\infty}$.
We show that
(3.32) $ux_{k}-ty_{k}=-2N_{\varepsilon}y_{k-1}\quad\text{ and }\quad tx_{k}-duy_{k}=2N_{\varepsilon}x_{k-1}$
for $k\geq 1$.
For $k=1$, we have $x_{1}=(at+bdu)/2$ and $y_{1}=(au+bt)/2$, so we find $ux_{1}-ty_{1}=-2bN_{\varepsilon}=-2y_{0}N_{\varepsilon}$. Similarly, $tx_{1}-duy_{1}=2N_{\varepsilon}x_{0}$ holds.
Similarly, we can prove the relationship holds for $k=2$.
For $k\geq 3$, we use induction and the recurrence relations for $x_{k}$ and $y_{k}$:
$ux_{k}-ty_{k}$ $=u\operatorname{Tr}\left(\varepsilon\right)x_{k-1}-uN_{\varepsilon}x_{k-2}-t\operatorname{Tr}\left(\varepsilon\right)y_{k-1}+tN_{\varepsilon}y_{k-2}$
$=\operatorname{Tr}\left(\varepsilon\right)(-2)N_{\varepsilon}y_{k-2}-N_{\varepsilon}(-2)N_{\varepsilon}y_{k-3}$
$=(-2)N_{\varepsilon}y_{k-1}.$
The proof of $tx_{k}-duy_{k}=2N_{\varepsilon}x_{k-1}$ is identical, so we omit it here.
We will also need
(3.33) $\frac{tu}{N_{\varepsilon}}x_{k}-\frac{t^{2}+du^{2}}{2N_{\varepsilon}}y_{k}$ $=-2N_{\varepsilon}y_{k-2}\quad\text{ and }$
$\frac{t^{2}+du^{2}}{2N_{\varepsilon}}x_{k}-\frac{dtu}{N_{\varepsilon}}y_{k}$ $=2N_{\varepsilon}x_{k-2},$
for $k\geq 2$ and
(3.34) $\left(\frac{u\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=-N_{\varepsilon}y_{k-3}\quad\text{ and }$
$\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{du\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=N_{\varepsilon}x_{k-3},$
for $k\geq 3$, which both follow from (3.32).
Since $3t^{2}+du^{2}=-N_{\varepsilon}+4t^{2}\equiv 4+4\equiv 0\pmod{8}$ and $t^{2}+3du^{2}=N_{\varepsilon}+4u^{2}\equiv 4+4\equiv 0\pmod{8}$, all the coefficients of $x_{k}$ and $y_{k}$ in the relationships in (3.34) are integers.
(i) Suppose that $t$ and $u$ are both even. Then $x_{k}$ and $y_{k}$ are both integers. Furthermore in this case, the coefficients of $x_{k}$ and $y_{k}$ in (3.32) can all be divided by $2$ to eliminate the factors of $2$ on the right-hand sides. So in this case, we find that
$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-1},y_{k-1}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1.$
(ii) Next suppose that $t$ and $u$ are both odd. In this case, we must have $d\equiv 5\pmod{8}$, since $\varepsilon$ is a unit and hence $t^{2}-du^{2}\equiv 4\pmod{8}$.
(ii-a) Suppose that $a$ and $b$ have opposite parities. In this case, $y_{1}=(au+bt)/2$ cannot be an integer. Similarly, $2x_{1}$ is odd.
Since $\left(\left(t+u\sqrt{d}\right)/2\right)^{2}=\left(t^{2}+u^{2}d\right)/4+(tu/2)\sqrt{d}=\left(t_{2}+u_{2}\sqrt{d}\right)/2$ with $t_{2}$ and $u_{2}$ both odd, we see that $x_{2}$ and $y_{2}$ are not integers either.
However, $\left(\left(t+u\sqrt{d}\right)/2\right)^{3}=\left(t^{3}+3tu^{2}d\right)/8+\left(\left(3t^{2}u+u^{3}d\right)/8\right)\sqrt{d}=\left(t_{3}+u_{3}\sqrt{d}\right)/2$ with both $t_{3}$ and $u_{3}$ even. Hence $x_{3}$ and $y_{3}$ are both integers.
By an inductive argument, one finds that $x_{k}$ and $y_{k}$ are both integers if and only if $3|k$, when $t$ and $u$ are both odd and $a$ and $b$ have opposite parities. So we consider only $\gcd\left(x_{3k},y_{3k}\right)$.
By (3.34), we have $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-3},y_{k-3}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1$.
(ii-b) If $a$ and $b$ have the same parity, then $y_{1}=(au+bt)/2\in\mathbb{Z}$. Similarly, $x_{1}\in\mathbb{Z}$. Using the recurrence relations, we find that all the $x_{k}$’s and $y_{k}$’s are integers.
Using (3.34), we find that if $k\equiv k_{1}\pmod{3}$ with $0\leq k_{1}\leq 2$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k_{1}},y_{k_{1}}\right)$.
So if $3|k$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right)$.
If $k\equiv 1\pmod{3}$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{1},y_{1}\right)$. From (3.32),
$\gcd\left(x_{1},y_{1}\right)|2\gcd\left(x_{0},y_{0}\right).$
If $k\equiv 2\pmod{3}$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{2},y_{2}\right)$. From (3.32),
$\gcd\left(x_{2},y_{2}\right)|\left(2\gcd\left(x_{1},y_{1}\right)\right)|\left(4\gcd\left(x_{0},y_{0}\right)\right).$
We now consider the parity of $x_{k}$ and $y_{k}$. Since $\gcd(a,b)=1$, at least one of $x_{0}$ or $y_{0}$ is odd.
Suppose first that $y_{0}$ is odd. Since $t$ is odd, we find that $y_{2}=ty_{1}+y_{0}$ is even if $y_{1}$ is odd and is odd otherwise.
In the first case (when $y_{1}$ is odd), we have $\gcd\left(x_{1},y_{1}\right)|\left(\gcd\left(x_{0},y_{0}\right)\right)$ and hence
$\gcd\left(x_{2},y_{2}\right)|\left(2\gcd\left(x_{1},y_{1}\right)\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
In the second case (when $y_{1}$ is even), we have $\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{1},y_{1}\right)$ and hence
$\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
Now we suppose that $x_{0}$ is odd. Since $t$ is odd, we find that $x_{2}=tx_{1}+x_{0}$ is even if $x_{1}$ is odd and is odd otherwise. As above, we find that
$\gcd\left(x_{2},y_{2}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
(iii) Now suppose that $t$ is even and $u$ is odd. In this case, $4|d$. In fact, if $4|t$, then $d\equiv 4\pmod{8}$ since $\varepsilon$ is a unit, and if $t\equiv 2\pmod{4}$, then $d\equiv 0\pmod{8}$.
Since $y_{1}=(au+bt)/2$, we see that $y_{1}\in\mathbb{Z}$ if and only if $a$ is even. Hence all the $y_{k}$’s are integers if $a$ is even. Since $t$ is even, using the recurrence relation, we see that for $a$ odd, $y_{k}\in\mathbb{Z}$ when $k$ is even and $y_{k}\not\in\mathbb{Z}$ otherwise.
Since $x_{1}=(at+bdu)/2$, we see that $x_{1}\in\mathbb{Z}$. Hence by the recurrence relation, $x_{k}\in\mathbb{Z}$ always holds.
As in case (ii) above, we use (3.34). If $k\equiv 0\pmod{3}$, then
$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right).$
If $k\equiv 1\pmod{3}$, then
$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{1},y_{1}\right)&\text{if $a$ is even},\\ \gcd\left(x_{4},y_{4}\right)&\text{if $a$ is odd}.\end{array}\right.$
We need to stop at $\gcd\left(x_{4},y_{4}\right)$ when $a$ is odd because $y_{1}$ is not an integer in this case. By (3.32), we see that the lemma holds if $a$ is even. By (3.33) and observing that $2|t$ and $4|d$, we see that $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-2},y_{k-2}\right)$. So $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right)$ when $a$ is odd.
If $k\equiv 2\pmod{3}$, then from (3.32) and (3.33), along with the observation when $k\equiv 1\pmod{3}$ and $a$ is odd, we have
$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{0},y_{0}\right).$
(iv) We cannot have $t$ odd and $u$ even because $t^{2}-du^{2}\equiv 0\pmod{4}$. ∎
We will use Lemma 3.10 in the proof of Lemma 3.11.
Lemma 3.10.
Let the sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be as defined in (1.1). Suppose that $k\neq 0$, $\gcd(a,b)=1$ and that $x_{k}$ and $y_{k}$ are both integers. Using the notation of Subsection 2.1 with $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$, $u_{1}=2x_{k}$ and $u_{2}=\pm 2\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$, then
$g^{2}/\gcd\left(a^{2},db^{4}\right)=2^{m},\text{ where $m\geq 0$.}$
Remark.
With more work, one can show that $m=0,1,2,4$.
Proof.
We break the proof into several cases.
(1) Suppose that $p>2$ is prime.
Put $v_{p}\left(g^{2}\right)=\ell\geq 0$. We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$. We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$. Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (recall from (2.4) that $g_{2}|t^{\prime}$ and $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$ here). From (2.4), we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$. Hence $\ell=\ell_{1}+\ell_{2}$.
Since $g_{1}|u_{2}$, by the definition of $g_{1}$, and from the expression for $u_{2}$ here, we have $p^{\ell_{1}}|\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)$. Similarly, from $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$, it follows that $p^{\ell_{2}}|\operatorname{sf}\left(N_{\alpha}\right)$. So $p^{\ell}|N_{\alpha}$.
In the same way, we find that $p^{\ell_{1}+2\ell_{2}}|x_{k}^{2}$.
By Lemma 3.9, since $p$ is odd, $p\nmid\gcd\left(x_{k},y_{k}\right)$. Hence, from $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $p^{\ell}|d$. Combining this with $p^{\ell}|N_{\alpha}$ and $N_{\alpha}=a^{2}-db^{4}$, we find that $p^{\ell}|a^{2}$. So $p^{\ell}|\gcd\left(a^{2},d\right)$.
We now show that $p^{\ell+1}\nmid\gcd\left(a^{2},d\right)$, so that the denominator of $g^{2}/\gcd\left(a^{2},d\right)$ has no odd prime factors.
Suppose that $p^{\ell+1}|\gcd\left(a^{2},d\right)$. Then $p^{\ell+1}|N_{\alpha}$. Also $p^{\ell+1}|x_{k}^{2}$ follows from $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$. By the definition of $\ell_{1}$ and $g_{1}$, this means that $p^{\ell_{1}}\parallel N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)$.
If $p\nmid\operatorname{sf}\left(N_{\alpha}\right)$, then $p^{\ell_{1}}\parallel N_{\alpha}$, but $\ell_{1}<\ell+1=\ell_{1}+\ell_{2}+1$, contradicting $p^{\ell+1}|N_{\alpha}$.
If $p\mid\operatorname{sf}\left(N_{\alpha}\right)$, then $p^{\ell_{1}+1}\parallel N_{\alpha}$. We must have $\ell_{2}=0$, since otherwise $\ell_{1}+1<\ell_{1}$ and we get a contradictions to $p^{\ell+1}|N_{\alpha}$. In this case, $p\nmid\left(x_{k}^{2}/g_{1}^{2}\right)$. So $p^{\ell_{1}}\parallel x_{k}^{2}$. But this contradicts $p^{\ell+1}|x_{k}^{2}$. Hence $p^{\ell+1}\nmid\gcd\left(a^{2},d\right)$. Therefore, for all primes $p>2$, $v_{p}\left(g^{2}/\gcd\left(a^{2},d\right)\right)=0$.
(2) We now consider $p=2$.
Put $v_{2}\left(\gcd\left(a^{2},db^{4}\right)\right)=\ell$. Since $\gcd(a,b)=1$, we also have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=\ell$. So $2^{\ell}|N_{\alpha}$ too. From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we find that $2^{\ell}|x_{k}^{2}$. Hence $2^{\ell}|\gcd\left(x_{k}^{2},N_{\alpha}\right)$.
Furthermore, if $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then $2^{\ell+1}|N_{\alpha}$.
(2.1) Suppose that $2\nmid\operatorname{sf}\left(N_{\alpha}\right)$.
We have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}\right)\right)\geq\ell+2$. Also $g_{2}=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}/g_{3}\right)\geq\ell$, recalling that $g_{3}=1,2$ or $4$.
(2.2) Suppose that $2|\left(N_{\alpha}\right)$.
Again, we break this case into subcases.
(2.2.1) If $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then we saw above that $v_{2}\left(N_{\alpha}\right)\geq\ell+1$. So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+2$. Here $g_{2}\geq 1$, so as in case (2.1), we have $v_{2}\left(g^{2}\right)\geq\ell$, as required.
(2.2.2) If $v_{2}\left(a^{2}\right)\neq v_{2}\left(db^{4}\right)$, then $v_{2}\left(N_{\alpha}\right)=\ell$. So $v_{2}\left(g_{1}^{2}\right)=\ell+1$. Since $2|\operatorname{sf}\left(N_{\alpha}\right)$, it follows that $\ell$ must be odd.
From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$ and $v_{2}\left(N_{\alpha}\right)=\ell$, we have either $v_{2}\left(x_{k}^{2}\right)=\ell$ and $v_{2}\left(dy_{k}^{2}\right)\geq\ell+1$; or else $v_{2}\left(x_{k}^{2}\right)\geq\ell+1$ and $v_{2}\left(dy_{k}^{2}\right)=\ell$. The first case is not possible because $\ell$ is odd. In the latter case, we must have $v_{2}(d)=\ell$ and $2\nmid y_{k}$. Also, $v_{2}\left(u_{1}/g_{1}\right)=(1/2)v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)\geq(1/2)\left(\ell+3-(\ell+1)\right)=1$, so $v_{2}\left(g_{2}\right)=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}g_{2}/g_{3}\right)=v_{2}\left(g_{1}^{2}g_{2}/4\right)=\ell$. ∎
Lemma 3.11.
[…]
Proof.
Since $\gcd\left(a^{2},db^{4}\right)\geq 2^{v_{2}\left(\gcd\left(a^{2},db^{4}\right)\right)}$, the inequality in equation (3.35) holds.
To establish the equality in equation (3.35), we use Lemma 3.10. From that lemma, we can write $g^{2}=2^{m}\gcd\left(a^{2},db^{4}\right)$, where $m\geq 0$. So, squaring both sides of the equality in (3.35), applying the definitions of $b^{\prime}$ and $d^{\prime}$ and simplifying, it is equivalent to
$\min\left(6+m,v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)\right)=\min\left(6,v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)\right).$
To show that this equality holds, we prove that if $v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)>6$, then $m=0$ (i.e., $g^{2}=\gcd\left(a^{2},db^{4}\right)$). Putting $\ell=v_{2}\left(\gcd\left(a^{2},db^{4}\right)\right)$, we must have $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=v_{2}\left(d\right)=\ell$. The second-last equality holds because $\gcd(a,b)=1$. From Lemma 3.9, we have $\gcd\left(x_{k},y_{k}\right)=1,2$. We will show that $\gcd\left(x_{k},y_{k}\right)=2$ is not possible here.
From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $\left(x_{k}/2^{\ell/2+1}\right)^{2}-\left(d/2^{\ell}\right)\left(y_{k}/2\right)^{2}N_{\alpha}/2^{\ell+2}$. Since $v_{2}\left(N_{\alpha}\right)\geq\ell+5$, it follows that $x_{k}/2^{\ell/2+1}$ and $y_{k}/2$ are both odd. Arguing mod $8$, we find that $d\equiv 1\pmod{8}$ and that $t$ and $u$ must both be even.
From (1.2), we can write $4y_{k}=\left(b^{2}\left(2du^{2}\pm 4\right)\pm 2atu\right)$ for some positive integers $t$ and $u$ with $t^{2}-du^{2}=\pm 4$. Since $t$ and $u$ are both even, it follows that $v_{2}\left(4y_{k}\right)=v_{2}\left(2du^{2}\pm 4\right)=2$.
If $a$ is even, then $b$ is odd. So if $u$ is also odd, then $4|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=1$. Hence $2\nmid y_{k}$.
If $u$ is even, then $16|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$. Hence $2\nmid y_{k}$.
We saw above that $\gcd\left(a^{2}\right)=\gcd\left(db^{4}\right)$, so if $a$ is odd, then so is $b$. Hence $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$, since $u$ is even. Here $8|(2atu)$, so $v_{2}\left(4y_{k}\right)=2$.
In none of these cases did we find that $2|y_{k}$. So
$v_{2}\left(g_{1}^{2}\right)=\min\left(v_{2}\left(4x_{k}^{2},4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4x_{k}^{2}\right)=\ell+2.$
This also implies that $v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)=0$, so $v_{2}\left(g_{2}\right)=0$.
Furthermore, since $v_{2}\left(4x_{k}^{2}\right)<v_{2}\left(4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)$, we also have $g_{3}=4$. Hence $g^{2}=\gcd\left(a^{2},db^{4}\right)$, as desired and the lemma follows. ∎
The quantities in Lemma 3.12 are related to the quantities $E$ and $Q$ that we will use in the proof of Proposition 4.1.
Lemma 3.12.
Suppose that $b=1$, $k\neq 0$, that $x_{k}$ and $y_{k}$ are both integers and that $g$ and $\mathcal{N}_{d^{\prime},4}$ are as above.
If $d\geq 105$, then
(3.36) $\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}>1.13$
and
(3.37) $\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>217.$
Proof.
We first obtain an analytic lower bound for $d$ such that (3.36) and (3.37) hold for larger $d$.
We have $|g|\mathcal{N}_{d^{\prime},4}\geq 2^{1+\min\left(v_{2}(N_{\alpha})/2,2\right)}\geq 2$ from (3.35). Since $b=1$, from Lemma 3.5(c), we find that $y_{k}>\left|N_{\alpha}\right|/4$ holds, so
$\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}>0.091\sqrt{d}.$
For $d\geq 155$, we find that $1.130.091\sqrt{d}>$ 1.13 holds.
Similarly,
$\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>\frac{21.12\sqrt{d}\,\left(\left|N_{\alpha}\right|/4\right)}{8\sqrt{\left|N_{\alpha}\right|}}\geq 0.66\sqrt{d},$
where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.35) to establish the first inequality. For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
We complete the proof computationally, checking all the remaining pairs, $(a,d)$. Since $d$ is bounded and $N_{\alpha}<0$ (so $a^{2}<d$), there are only finitely many such pairs.
For each pair, we check (3.36) and (3.37) by computing their left-hand sides for all $k\neq 0$ such that $20.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<$ 2 and $30021.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<$ 300. No counterexamples to (3.36) and (3.37) were found. The script took 112 seconds to run on a laptop with an Intel i7-9750H processor and 16gb of RAM.
The smallest value of the left-hand side of (3.36) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, $k=-1$, $x_{k}=-59$ and $y_{k}=5$. The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, $k=-1$, $x_{k}=-130$ and $y_{k}=11$.
Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$. ∎
4. Proposition 4.1 and its proof
Our results in Subsection 1.4 follow from the next proposition.
Proposition 4.1.
Let $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be defined by (1.1).
If $b=1$ and $-N_{\alpha}$ is a square, then there is at most one integer square among all distinct elements of $\left(y_{k}\right)_{k=-\infty}^{\infty}$ which satisfies
$y_{k}>\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right).$
Conjecture 1.2 would follow for all sequences with $b=1$ and $-N_{\alpha}$ a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$. This is due to Lemma 3.5.
4.1. Prerequisites
In this subsection, we collect some inequalities what will be required in the subsections that follow.
We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$.
We shall initially assume that
(4.1) $d\geq 128.$
This arises from Lemma 3.12 and will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof, in Subsection 4.6.
As in Lemma 3.8, we put $\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|,$
where $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k}$, $r_{\ell}$, $s_{k}$ and $s_{\ell}$ as in Proposition 3.1, which are associated with $\left(x_{k},y_{k}\right)$ and $\left(x_{\ell},y_{\ell}\right)$, respectively. From (3.31) in the proof of Lemma 3.8(b), we have
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.06.$
Thus we can apply Lemma 3.6(a) with $c_{1}=0.06$ to find that
(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
The equality on the left-hand side is from the equalities in (3.19).
By our choice of $k$ and $\ell$, and by Lemma 3.8(b) when $-N_{\alpha}$ is not a square, $\zeta_{4}=\pm 1\in\mathbb{Q}\left(\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$. This is important for us here as $\zeta_{4}\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)/\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ must be in an imaginary quadratic field in order to apply Lemma 2.1 to obtain a lower bound for the rightmost quantity in (4.2).
We need to derive a lower bound for the far-right quantity in (4.2). To do so, we shall use the lower bounds in Lemma 2.1 with a sequence of good approximations $p_{r}/q_{r}$ obtained from the hypergeometric functions. So we collect here the required quantities.
Since $y_{k}\geq 4$ and $N_{\alpha}>-db^{4}=-d$ with $b=1$ here, we obtain
(4.3) $x_{k}^{2}=dy_{k}^{2}+N_{\alpha}>d\left(y_{k}^{2}-1\right)\geq 0.9375dy_{k}^{2}.$
So
(4.4) $\sqrt{x_{k}^{2}-N_{\alpha}}=\sqrt{dy_{k}^{2}}<1.04\left|x_{k}\right|.$
Using the notation of Subsection 2.1, let $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$, $u_{1}=2x_{k}$, $u_{2}=2\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$ and $d^{\prime}$ is as defined in (2.10).
Substituting these quantities along with $\mathcal{D}_{4}=e^{1.68}$ from Lemma 2.3(a) into the definition of $E$ in (2.13) and applying (4.3), we have
(4.5) $E$ $=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}>\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{4e^{1.68}\left|N_{\alpha}\right|}$
$>\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|x_{k}\right|+\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{10.74\left|N_{\alpha}\right|}>\frac{|g|\mathcal{N}_{d^{\prime},4}\left(1+\sqrt{0.9375}\right)\sqrt{d}\,y_{k}}{10.74\left|N_{\alpha}\right|}$
$>\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}.$
From Lemma 3.5(c), $|g|\mathcal{N}_{d^{\prime},4}\geq 2$ and (4.1), we have $E>1$, as required for its use with Lemma 2.1.
Similarly, using (4.3) and (3.37) in Lemma 3.12 we have
(4.6) $Q>\frac{2e^{1.68}\left(1+\sqrt{0.9375}\right)\sqrt{d}y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>217,$
so the condition $Q>1$ in Lemma 2.1 is satisfied.
From $N_{\alpha}<0$ and the equality in (4.4), we have $x_{k}<\sqrt{x_{k}^{2}-N_{\alpha}}=\sqrt{d}\,y_{k}$, so from Lemma 2.3(a) we have
(4.7) $Q=\frac{e^{1.68}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{|g|\mathcal{N}_{d^{\prime},4}}<4\frac{e^{1.68}\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}<\frac{21.47\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}.$
Recall from (2.12) that we take $k_{0}=0.89$.
Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.8(a), we can take
(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.46\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
Also from Lemma 3.8(a), we have $|\varphi|<0.6$, so the condition $|\omega-1|<1$ in Lemma 2.2 is satisfied.
Let $q=x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ and $p=x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$. Recall from (3.25) with $b=1$ that
(4.9) $|q|=\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}.$
We are now ready to deduce the required contradiction from the assumption that there are two sufficiently large squares when $-N_{\alpha}$ is a square. We will break the proof into five parts.
With $r_{0}$ as in Lemma 2.1, we separate the case of $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$, from the case of $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$. In the first case, Lemma 2.1 provides a suitable lower bound for the approximation. But in the second case, Lemma 2.1 is not strong enough, so we work directly with the approximations themselves.
4.2. $r_{0}=1$ and $\zeta_{4}p/q\neq p_{1}/q_{1}$ for all $4$-th roots of unity, $\zeta_{4}$
We start by determining an upper bound for $y_{\ell}$ for all $r_{0}\geq 1$ when $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$, since we will also need such a result in Subsection 4.4.
From equation (4.2), along with Lemma 2.1(b) and (3.25) with $b=1$, we have
(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.99(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
Applying (2.12) and (4.7) to (4.10), we obtain
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{3.99(1-c)}{0.89\left(21.47\sqrt{d}\,y_{k}/\left(|g|\mathcal{N}_{d^{\prime},4}\right)\right)^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
After taking the fourth power of both sides and rearranging, we find that
(4.11) $\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}>y_{\ell}^{3}.$
Specialising to the case when $r_{0}=1$ and using $|g|\mathcal{N}_{d^{\prime},4}\geq 2$ from (3.35), we have
(4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
We will now combine (4.12) with the gap principle in Lemma 3.7(a). to show that this case cannot occur.
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.7(a), we find that
$57.6^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
which we can rewrite as
$y_{k}<\frac{0.24}{1-c}\frac{\left|N_{\alpha}\right|^{2}}{d^{3/2}}<\frac{\left|0.24N_{\alpha}\right|}{(1-c)d^{1/2}}.$
The last inequality holds because $\left|N_{\alpha}\right|=d-a^{2}<d$.
By (4.1), we have $0.250.24/\left((1-c)d^{1/2}\right)<$ 0.25 if $c<0.9$. We will see in Subsection 4.4 that $c=0.75$ is a good choice. So $y_{k}<\left|N_{\alpha}\right|$/4, which contradicts Lemma 3.5(c). Hence this case cannot hold.
4.3. $r_{0}=1$ and $\zeta_{4}p/q=p_{1}/q_{1}$ for some $4$-th root of unity, $\zeta_{4}$
As in Subsection 4.2, we start by proving an upper bound for $y_{\ell}$ that holds for all $r_{0}\geq 1$ with $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$.
From the definitions of $p_{r_{0}}$, $q_{r_{0}}$ and $R_{r_{0}}$ in (2.9), along with parts (a) and (e) of Lemma 2.2, we have
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega^{1/4}-p_{r_{0}}\right|$
$=\left|\frac{N_{d,4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d,4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}\right|$
$=\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
(4.13) $\geq\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
From Lemma 2.2(d), we have
(4.14) $\left|Y_{1,4,r_{0}}(\omega)\right|<1.072\frac{r_{0}!\Gamma(3/4)}{\Gamma(r_{0}+3/4)}\left|1+\sqrt{\omega}\right|^{2r_{0}}.$
We can write $\omega-1$ as $\left(2N_{\alpha}+2x_{k}N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}^{2}-N_{\alpha}\right)$, so
$|\omega-1|=\sqrt{\frac{4\left|N_{\alpha}\right|}{x_{k}^{2}-N_{\alpha}}}=2\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
Since $\left|\sqrt{\omega}+1\right|<2$, it follows that
$\left|\sqrt{\omega}-1\right|=\frac{\left|\omega-1\right|}{\left|\sqrt{\omega}+1\right|}>\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
From these two inequalities, we also obtain
$\left|\frac{(\omega-1)^{2r_{0}+1}}{\left(1+\sqrt{\omega}\right)^{2r_{0}}}\right|=\left|(\omega-1)\left(\sqrt{\omega}-1\right)^{2r_{0}}\right|>2\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}(\omega)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.291}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
Applying (4.2) with this inequality, we obtain
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{1.16}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1},$
so
(4.15) $1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>y_{\ell}.$
We now specialise to the case of $r_{0}=1$ and apply the gap principle in Lemma 3.7.
Our gap principle in Lemma 3.7(a) with $b=1$, along with (4.15), implies that
$6.92\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.6\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.6\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible.
4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
Here we establish a stronger gap principle for $y_{k}$ and $y_{\ell}$ than the one in Lemma 3.7. We then use this with the upper bound for $y_{\ell}$ in (4.11) to obtain a contradiction.
We start by deriving a lower bound for $y_{\ell}$ that holds in both this step and in the next step.
From the definition of $r_{0}$ in Lemma 2.1, along with (4.6) and $E>1$, we have
(4.16) $|q|\geq\frac{c(Q-1)}{\ell_{0}(Q-1/E)}E^{r_{0}-1}>0.995cE^{r_{0}-1}/\ell_{0}.$
Recall that $|q|=\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}$ by (4.9). Thus
$\left(y_{k}y_{\ell}\right)^{1/4}>\frac{0.995cE^{r_{0}-1}}{\ell_{0}\sqrt{f_{k}f_{\ell}}}.$
Applying (4.5) and (4.8), and then (4.3), to this inequality, we obtain
$\left(y_{k}y_{\ell}\right)^{1/4}$ $>\frac{0.995c\left|x_{k}\right|}{0.46\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}$
$>\frac{2.1c\sqrt{d}\,y_{k}}{\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}.$
Taking the fourth power of both sides and rearranging, we find that this inequality implies
(4.17) $y_{\ell}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}.$
With this lower bound for $y_{\ell}$, we now focus for the rest of this subsection on when $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$.
We now take the third power of both sides of this inequality and combine it with the upper bound for $y_{\ell}^{3}$ in (4.11), obtaining
(4.18) $\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}$
$>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{12}c^{12}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{12r_{0}}d^{6r_{0}}y_{k}^{12r_{0}-3}.$
Using elementary calculus, $c^{12}(1-c)^{4}$ is monotonically increasing for $0<c\leq 0.75$. So we put $c=0.75$ and find that for such $c$, $c^{12}(1-c)^{4}>0.000124$. Applying this to (4.18) and simplifying, we have
(4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
and then
$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
Using the equality in (3.35) of Lemma 3.11, we can compute that $\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}/d^{4}\leq 2^{20}/17^{4}$ (the max value occurs when $d=17a^{2}$ so that $\left|N_{\alpha}\right|=16a^{2}$) to obtain
(4.20) $0.01>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
But if
(4.21) $y_{k}>59.2\left|N_{\alpha}\right|^{3/2}/\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right),$
then the right-hand side is greater than $1$, so this is not possible.
4.5. $r_{0}>1$ and $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$
We now combine our upper bound for $y_{\ell}$ in (4.15) with our lower bound for $y_{\ell}$ in (4.17). Thus
$1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}$
and so
$1.73r_{0}^{1/2}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}^{2}f_{\ell}^{2}}}\right)^{4}c^{4}\left(\frac{0.1832^{4}|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{4\left|N_{\alpha}\right|^{3}}\right)^{r_{0}}d^{r_{0}}y_{k}^{2r_{0}-2}.$
We can show that $0.1832^{4r_{0}}/r_{0}^{1/2}>0.175^{4r_{0}}$, with the minimum being attained at $r_{0}=3$. Applying this, along with $c=0.75$, yields
$1>\frac{0.739d}{f_{k}^{2}f_{\ell}^{2}\left|N_{\alpha}\right|}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
Combined with $d>N_{\alpha}$, this implies that
(4.22) $1>\frac{0.739}{f_{k}^{2}f_{\ell}^{2}}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
We now proceed similarly to the way we did in Subsection 4.4.
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b). Also, as $r_{0}-1\geq 1$, it follows that $0.739\cdot 0.0002344^{r_{0}-1}>\left(0.0001732\right)^{r_{0}-1}$, the above inequality implies
$1>\left(\frac{0.0001732|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}>\left(y_{k}^{2}\frac{d\left(|g|\mathcal{N}_{d,4}\right)^{4}}{5774\left|N_{\alpha}\right|^{3}}\right)^{r_{0}-1}.$
But if
(4.23) $y_{k}\geq\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d,4}\right)^{2}},$
then the right-hand side is greater than $1$, so this is not possible.
4.6. Small $d$
To complete the proof of Proposition 4.1, we need to remove the assumption on $d$ in (4.1).
We check directly all pairs $(a,d)$ of positive integers with $2\leq d\leq 127$ not a square and $-N_{\alpha}=d-a^{2}$ a square. Note that there are only finitely many such pairs, since $N_{\alpha}=a^{2}-d<0$.
If $y_{k}=y^{2}$ is a square, then from (1.1), we know that $x_{k}^{2}-dy^{4}=N_{\alpha}$. By Theorem 1.1 of [1], our theorem holds for $N_{\alpha}=-1$, so we may also assume that $N_{\alpha}\leq-4$. There are 74 such pairs $(a,d)$.
For each such pair, we solved $x^{2}-dy^{4}=N_{\alpha}$ using Magma (version V2.28-2) [8] and its IntegralQuarticPoints() function. For the 74 equations, this calculation took 17.66 seconds using MAGMA’s online calculator. 21 of the equations had at least two solutions in positive integers. Only two of these had three solutions in positive integers:
$x^{2}-17y^{4}=-16$ has the solutions $(x,y)=(1,1),(16,2)$, $(103,5)$,
$x^{2}-68y^{4}=-64$ has the solutions $(x,y)=(2,1),(32,2)$, $(206,5)$.
None of the equations had more solutions.
For $x^{2}-17y^{4}=-16$, note that $\left(103+5^{2}\sqrt{17}\right)/\left(1+\sqrt{17}\right)=\left(161+39\sqrt{17}\right)/8$, so the solutions $(103,5)$ and $(1,1)$ arise from different sequences. Hence Proposition 4.1 holds for $(a,d)=(1,17)$. Similarly, Proposition 4.1 holds for $(a,d)=(2,68)$ too. This completes the proof of Proposition 4.1.
5. Proof of Theorem 1.4
If $N_{\alpha}$ is even, then $|g|\mathcal{N}_{d^{\prime},4}\geq 4$ by Lemma 3.11. So the right-hand side of the inequality in Proposition 4.1 is at most $4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But for $u\geq 5$, we have $y_{k}>25\left|N_{\alpha}\right|/4$, by Lemma 3.5(c). So Theorem 1.4 holds for $u\geq 5$ and we need only consider $1\leq u\leq 4$.
Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}>81\left|N_{\alpha}\right|/4>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$ and we need only consider $1\leq u\leq 8$.
Lemma 5.1.
If $y_{\ell}>y_{k}>1$ are two squares, then $k=\pm 1$.
Proof.
Suppose that $y_{k}$ is a square for some $|k|\geq 2$.
From Lemma 3.5(c), $\left|N_{\alpha}\right|=d-a^{2}<d$ and that by Theorem 1.1 of [1] we may assume $\left|N_{\alpha}\right|\geq 4$, we have $y_{k}>\left|N_{\alpha}\right|^{2}u^{4}/8$. So, for $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 8\left|N_{\alpha}\right|>8\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
However, we saw above that for $N_{\alpha}$ even, if $y_{k}>4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$. So we need only consider $u=1$ when $N_{\alpha}$ is even.
Similarly, for $N_{\alpha}$ odd, if $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 18\left|N_{\alpha}\right|>18\left|N_{\alpha}\right|^{2}>90\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But, we saw above that for $N_{\alpha}$ even, if $y_{k}>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$. So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
For $u=2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}$ for $|k|\geq 2$. So if $\left|N_{\alpha}\right|\geq 11$ is odd, then we have $y_{k}\geq 22\left|N_{\alpha}\right|>20\left|N_{\alpha}\right|$. So, for odd $N_{\alpha}$, we need only consider $N_{\alpha}=-9$. In this case, $N(\varepsilon)=\left(t^{2}-4d\right)/4=\pm 1$, so $t^{2}-4a^{2}=-4N_{\alpha}\pm 4=32,40$. This means that $\left(t,a,N_{\alpha}\right)=(\pm 6,1,-9)$. Here we have $y_{\pm 2}\geq 493>20\left|N_{\alpha}\right|$.
For $u=1$, we proceed similarly. Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/8$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $N_{\alpha}$ is even and $9\leq\left|N_{\alpha}\right|\leq 121$ when $-N_{\alpha}$ is an odd square.
For each value of $N_{\alpha}$, the norm of $\varepsilon$ leads us to an equation of the form $t^{2}-a^{2}=-N_{\alpha}\pm 4$. We solve each of these equations over the integers using PARI/GP and calculate $y_{k}$ and $y_{-k}$ for $k\geq 2$ until the lower bound from Proposition 4.1 is exceeded (we never had to go beyond $k=7$). No squares were found. ∎
Lemma 5.2.
Suppose that $N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
(a) If $t^{2}-du^{2}=4$, then $t\equiv 2\pmod{4}$ and $u\equiv 0\pmod{4}$.
(b) If $t^{2}-du^{2}=-4$, then $u\equiv 2\pmod{4}$.
Proof.
For $t^{2}-du^{2}=\pm 4$, if $t$ is odd, then $d\equiv u^{2}\equiv 1\pmod{4}$. Since $N_{\alpha}$ is odd, it follows that $a$ is even. Therefore $4y_{\pm 1}=2du^{2}\pm 2atu\pm 4\equiv 2\pmod{4}$, which is not possible since $y_{\pm 1}\in\mathbb{Z}$. Therefore, if $N_{\alpha}$ is odd, $t$ must be even.
(a) Next we show that if $t^{2}-du^{2}=4$, then $4\nmid t$. If $4\mid t$, then $du^{2}\equiv 12\pmod{16}$. Since $-N_{\alpha}$ is an odd square, we have $-N_{\alpha}\equiv 1,9\pmod{16}$. If $a$ is odd, then $a^{2}\equiv 1\pmod{8}$. So $d\equiv 2\pmod{8}$ and hence $u^{2}\equiv 6\pmod{8}$, which is not possible. If $a$ is even, then $a^{2}\equiv 0,4\pmod{16}$. Since $-N_{\alpha}=d-a^{2}\equiv 1,9\pmod{16}$, it follows that $d\equiv 1\pmod{8}$. So $u^{2}\equiv 12\pmod{16}$, which is also impossible. So $4\nmid t$.
Therefore, we must have $t\equiv 2\pmod{4}$, if $t^{2}-du^{2}=4$ and $N_{\alpha}$ is an odd square.
From $t\equiv 2\pmod{4}$ and $t^{2}-du^{2}=4$, we have $du^{2}\equiv 0\pmod{16}$.
Since $N_{\alpha}$ is odd, we have $-N_{\alpha}\equiv 1,9\pmod{16}$. Also $a^{2}\equiv 0,1,4,9\pmod{16}$. Hence $d\equiv 1,2,5,9,10,13\pmod{16}$ and so $8|u^{2}$. Thus $4|u$.
(b) We proceed similarly. Suppose that $4\mid t$, then $du^{2}\equiv 4\pmod{16}$. As in the proof of part (a), if $a$ is odd, then $d\equiv 2\pmod{8}$. Hence $u^{2}\equiv 2\pmod{8}$, which is not possible.
If $a$ is even, as in the proof of part (a), then $d\equiv 1\pmod{8}$. So $u^{2}\equiv 4\pmod{16}$. I.e., $u\equiv 2\pmod{4}$.
If $t\equiv 2\pmod{4}$, then $du^{2}\equiv 8\pmod{16}$. Again, if $a$ is odd, then $d\equiv 2\pmod{8}$ and so $u^{2}\equiv 4\pmod{8}$.
If $a$ is even, then $d\equiv 1\pmod{8}$ and $u\equiv 8\pmod{16}$, which is not possible. ∎
We now prove a lemma concerning $N_{\alpha}$ odd.
Lemma 5.3.
If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $\gcd\left(a^{2},d\right)=1$ and $u=2$, $t^{2}-du^{2}=-4$.
Proof.
(1) We start by considering $t^{2}-du^{2}=4$.
From Lemma 5.2(a), if $t^{2}-du^{2}=4$ and $-N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
(1-i) First, we eliminate $u=4$. We can write $N_{\alpha}=-(2n+1)^{2}$, so $t^{2}-\left(a^{2}+(2n+1)^{2}\right)u^{2}=4$ becomes $t^{2}-64n^{2}-64n=16a^{2}+20$. But $t^{2}\equiv 20\pmod{64}$, has no solution, so $a$ must be odd. Hence $t^{2}-64n^{2}-64n=16a^{2}+20$ implies $t^{2}\equiv 36\pmod{64}$. So $t\equiv\pm 6\pmod{16}$.
Expanding the expressions for $4y_{\pm 1}$, with $a=2a_{1}+1$, $N_{\alpha}=-(2n+1)^{2}$ and $t=16t_{1}\pm 6$, we find that $4y_{\pm 1}\equiv 20\pmod{32}$. But this congruence has no solution with $y_{\pm 1}$ a square. Hence $u=4$ is not possible.
(1-ii) Now we eliminate $u=8$. Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$. From this, $-N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
We will show that $d\equiv 5\pmod{9}$ is not possible. In this case, from $t^{2}-64d=4$, we find that $t^{2}\equiv 0\pmod{9}$. So $t\equiv 0\pmod{3}$. Since $d\equiv 5\pmod{9}$), we have $y_{\pm 1}=1+32d\pm 4at\equiv 2\pm 4at\pmod{3}$. Since $3|t$, it follows that $y_{\pm 1}\equiv 2\pmod{3}$, which can never be square. Hence $9|\gcd\left(a^{2},d\right)$.
(Since we will use this same argument in several places in the proof of this lemma and the next, we wrote programs (in both Maple and PARI/GP) to automate these steps and prove that $\gcd\left(a^{2},d\right)$ has certain factors (typically, powers of $2$ and $3$). In fact, using this same code, we could have eliminated case (1-i) above too.)
Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 36$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>16\left|N_{\alpha}\right|$ here. So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
Thus for $N_{\alpha}$ odd, we do not need to consider $t^{2}-du^{2}=4$.
(2) Now we consider $t^{2}-du^{2}=-4$.
(2-i) From Lemma 5.2(b), we have $u\equiv 2\pmod{4}$. For $u=6$, $t^{2}-du^{2}=-4$ is not possible modulo $9$. So we must have $u=2$ if $t^{2}-du^{2}=-4$.
(2-ii) Finally, we show that $\gcd\left(a^{2},d\right)=1$ in this case.
We suppose otherwise. since $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)$ must be odd and since $d=\left(t^{2}+4\right)/4$, any odd prime factor, $p$, of the gcd must also be a factor of $t^{2}+4$. This means that $t^{2}\equiv-4\pmod{p}$. Hence $p\equiv 1\pmod{4}$. Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$. Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 100$. Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|$ here. So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$. ∎
We now prove an analogous lemma for $N_{\alpha}$ even.
Lemma 5.4.
If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
Proof.
As in the proof of Lemma 5.3, we proceed in steps.
(1) We start by considering $t^{2}-du^{2}=4$.
(1-i) If $u=4$ and $t^{2}-du^{2}=4$, then we can argue as for $u=8$ in the proof of Lemma 5.3 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$. Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 144$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq(76/144)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>4\left|N_{\alpha}\right|$ here. So we can exclude $u=4$ and $t^{2}-du^{2}=4$ from consideration.
(1-ii) Suppose that $u=3$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. We will again argue as for $u=8$ in the proof of Lemma 5.3 and use the Maple program mentioned there modulo $32$ to show that $64|\gcd\left(a^{2},d\right)$.
Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 256$. Hence the theorem holds if $y_{\pm 1}>(76/256)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}>(9/4)\left|N_{\alpha}\right|$ here.
(1-iii) Suppose that $u=2$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. Using the argument for $u=8$ in the proof of Lemma 5.3 and the Maple program mentioned there modulo $16$, we obtain $16|\gcd\left(a^{2},d\right)$.
Doing the same modulo $9$, we obtain $9|\gcd\left(a^{2},d\right)$. Combining this with $16|\gcd\left(a^{2},d\right)$ and applying Lemma 3.11, we have $|g|\mathcal{N}_{d,4}\geq 24$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/24^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|u^{2}/4=\left|N_{\alpha}\right|$ here. So we can exclude $u=2$ and $t^{2}-du^{2}=4$ from consideration.
(1-iv) Suppose that $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. Here too we use the argument and Maple program from the proof of Lemma 5.3, first modulo $9$ and then modulo $16$. From the latter, we obtain $64|\gcd\left(a^{2},d\right)$. So by Lemma 3.11, $|g|\mathcal{N}_{d,4}\geq 48$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/48^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|/4$ here. So we can exclude $u=1$ and $t^{2}-du^{2}=4$ from consideration.
(2) We now suppose that $t^{2}-du^{2}=-4$.
Independent of the parity of $N_{\alpha}$, for $u=3$ and $4$, $t^{2}-du^{2}=-4$ is not possible modulo $9$ and $16$, respectively. So we can only have $u=1,2$ with $t^{2}-du^{2}=-4$
(2-i) Consider $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$.
We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $-N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d,4}\geq 16$. So $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}\leq\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. However $76/16^{2}$ is bigger than $1/4$, so we must work a bit harder to eliminate $8|b_{1}$.
If $a^{2}\geq 0.292d$, then
$\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}<\left(76\sqrt{1-0.292}/16^{2}\right)\left|N_{\alpha}\right|<0.2498\left|N_{\alpha}\right|.$
By Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|/4$. So $a^{2}\geq 0.292d$ is excluded.
Suppose that $d\geq 80$. Then $t^{2}=d-4$, so $t^{2}\geq(76/80)d$. If $a^{2}<0.292d$, then $t-a>\sqrt{76d/80}-\sqrt{0.292d}$ and so $(t-a)^{2}/4>0.047d>0.047\left|N_{\alpha}\right|$. We can write $4y_{\pm 1}=\left(t\pm au\right)^{2}-N_{\alpha}u^{2}$. So here with $u=1$, we have $y_{\pm 1}\geq\left(t-a\right)^{2}/4+\left|N_{\alpha}\right|/4>0.297\left|N_{\alpha}\right|$. But $76\left|N_{\alpha}\right|/16^{2}=0.296875\left|N_{\alpha}\right|$. So we can exclude $8|b_{1}$, provided $d\geq 80$.
For $d<80$ with $u=1$ and $-N_{a}$ an even square divisble by $64$, there is just one possibility: $d=68$, $t=8$, $u=1$, $a=2$, so $N_{\alpha}=-64$ and $y_{-1}=25$. But here, $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}<\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}<19$, so this case is excluded too and hence we can exclude $8|b_{1}$ altogether.
Hence $b_{1}\equiv 2\pmod{4}$ and so $-N_{\alpha}\equiv 4\pmod{16}$.
Now we show that $\gcd\left(a^{2},d\right)=1$ or $4$.
We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $-N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 00 in the second case.
We consider first $\gcd\left(a^{2},d\right)>4$ odd. So by Lemma 3.11 and the argument in the previous paragraph, $|g|\mathcal{N}_{d,4}=4\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$. If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $40.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/$ 4, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
Since $d=t^{2}+4$, arguing modulo $3$, we see that $3\nmid d$ and hence $\gcd\left(a^{2},d\right)=9$ is not possible.
Now consider $\gcd\left(a^{2},d\right)>4$ even. Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$). So by Lemma 3.11 and the argument above, $|g|\mathcal{N}_{d,4}=2\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
We saw above that $3\nmid d$, so we can ignore $\gcd\left(a^{2},d\right)=6^{2}$. If $\gcd\left(a^{2},d\right)\geq 100$, then $\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}\leq 0.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/4$, the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
(2-ii) Lastly, we consider $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$. Arguing modulo $64$, we have $d\equiv a^{2}\equiv 1\pmod{16}$ and $8|t$. So $16|N_{\alpha}$ and, in the notation of Lemma 3.11, $16|b^{\prime}$. Thus $|g|\mathcal{N}_{d^{\prime},4}=8\sqrt{\gcd\left(a^{2},d\right)}$, where $\gcd\left(a^{2},d\right)$ is odd.
It follows that $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}\leq\frac{76\left|N_{\alpha}\right|^{3/2}}{64\left(\sqrt{d}\gcd\left(a^{2},d\right)\right)}$. However $76/64$ is bigger than $1$, so we proceed as in case (2-i) above.
If $a^{2}\geq 0.292d$, then
$\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(64\sqrt{d}\gcd\left(a^{2},d\right)\right)}<\frac{76\sqrt{1-0.292}}{64}\left|N_{\alpha}\right|<0.9992\left|N_{\alpha}\right|.$
By Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$. So $a^{2}\geq 0.292d$ is excluded.
Suppose that $d\geq 2$. Then $t^{2}=4d-4$, so $t^{2}\geq 2d$. If $a^{2}<0.292d$, then $t-a>\sqrt{2d}-\sqrt{0.292d}$ and so $(t-a)^{2}/4>0.19d>0.19\left|N_{\alpha}\right|$. We can write $4y_{\pm 1}=\left(t\pm au\right)^{2}-N_{\alpha}u^{2}$. So here with $u=2$, we have $y_{\pm 1}\geq\left(t-a\right)^{2}/4+\left|N_{\alpha}\right|>1.19\left|N_{\alpha}\right|$. But $76\left|N_{\alpha}\right|/64=1.1875\left|N_{\alpha}\right|$. So we can exclude the case where $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$. ∎
Theorem 1.4 now follows from the bounds on $u$ from Proposition 4.1, along with Lemmas 5.1, lem:nrmA-odd and lem:nrmA-even.
6. Examples
In this section, we give examples showing that our conjectures and results are best possible.
6.1. Examples for Conjecture 1.1
For $d=2$ with $t=u=2$.
$a$ $b$ indices, $k$ $\sqrt{y_{k}}$
1 $3^{2}$ [0, -1, 3, -5] [3, 5, 31, 167]
1019 $27^{2}$ [0, 1, -3, -7] [27, 65, 29, 983]
167 $13^{2}$ [0, 1, -3, 4] [13,29,71, 407]
157 $29^{2}$ [0, -1, 3, -4] [29, 47, 307, 649]
1 $41^{2}$ [0, -1, -9, 11] [41, 71, 80753, 470861]
1633 $65^{2}$ [0, -1, -4, 7] [65, 97, 1331, 24791]
48479 $211^{2}$ [0, -3, 4, -7] [211, 1007, 6743, 34205]
45649 $677^{2}$ [0, -1, -4, 4] [677, 1133, 15679, 16825]
1940147 $1217^{2}$ [0, -3, 4, -11] [1217, 3289, 40573, 3794239]
600589 $2213^{2}$ [0, -1, -4, 4] [2213, 3673, 50801, 55415]
20509501 $8689^{2}$ [0, -1, 3, -4] [8689, 13619, 94393, 187603]
255488029 $13457^{2}$ [0, -1, 3, -4] [13457, 5683, 189241, 15821]
409660129 $17023^{2}$ [0, -1, -4, -8] [17023, 7073, 7949, 269495]
3032771269 $46313^{2}$ [0, -1, -4, -8] [46313, 19213, 15269, 516625]
We also found such examples with $(d,t,u,a,b)=(3,4,2,672,8281)$, $(6,10,4,78,49)$, $(6,10,4,34986,22201)$, $(6,10,4,3663828,5094049)$, $(30,22,4,826320,1234321)$ and $(37,12,2,138,25)$.
At least for $d=2$ and $d=6$, it appears there may be infinitely many such examples.
6.2. Examples for Conjecture 1.2
Let $n\geq 5$ be an odd integer. Put $a=\left(n^{2}-9\right)/4$, $d=\left(n^{4}-2n^{2}+17\right)/16$, so that $N_{\alpha}=4-n^{2}$. With $t=\left(n^{2}-1\right)/2$ and $u=2$, $\varepsilon=\left(t+u\sqrt{d}\right)/2$ is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$. We have $y_{1}=\left(n^{2}-3\right)^{2}/4$ and $y_{-1}=n^{2}$. This example shows that Conjecture 1.2 is best possible, if true.
Let $n>5$ satisfy $n\equiv 1\pmod{4}$ and not divisible by $5$. Put $a=\left(n-5\right)/4$, $d=\left(n^{2}+6n+25\right)/16$. Here $N_{\alpha}=-n$. With $t=2a+4$ and $u=2$, $\varepsilon=\left(t+u\sqrt{d}\right)/2$ is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ and we have $y_{1}=(n+1)^{2}/4$. If $n$ is a square, then this example shows that Conjecture 1.2 is best possible when $\left|N_{\alpha}\right|$ is a perfect square.
6.3. Examples for Conjecture 1.3
The examples in Subsection 6.1 also apply in this more general case to show there can be at least four distinct squares.
6.4. Examples for Lemma 3.5
We can show that the lower bound in Lemma 3.5(c) is actually best possible. Let $n$ be a positive integer such that $a=2n^{2}-3$ is not divisible by $5$ and put $d=4n^{4}-8n^{2}+8$. We have $N_{\alpha}=1-4n^{2}$. Then $(t,u)=(a+1,1)$, $y_{-1}=n^{2}$ and so $y_{-1}/\left|N_{\alpha}\right|\rightarrow 1/4$ from above as $n\rightarrow+\infty$. There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $-N_{\alpha}$ is a square too.

### v2

@@
 We investigate the number of squares in a very broad family of binary recurrence sequences with $u_{0}=1$.
-We show that there are at most two distinct squares in such sequences (the best possible result), except under such very special conditions where we prove there are at most three such squares.
+We show that there are at most two distinct squares in such sequences (the best possible result), except under very special conditions where we prove there are at most three such squares.
 Key words and phrases:
@@
 While some progress on arithmetic questions has been achieved for sequences of Lucas numbers (numbers of the form $u_{n}=\left(\alpha^{n}-\beta^{n}\right)/(\alpha-\beta)$, where $\alpha$ and $\beta$ are algebraic numbers such that $\alpha+\beta$ and $\alpha\beta$ are non-zero coprime rational numbers and $\alpha/\beta$ is not a root of unity) and their associated sequences ($v_{n}=\alpha^{n}+\beta^{n}$), much less is known for other binary recurrence sequences.
-As with sequences of Lucas numbers, other binary recurrence sequences are important for the solution of Diophantine equations such as $aX^{2}-bY^{4}=c$, which is important in its own right, but is also a quartic model of elliptic curves.
+As with sequences of Lucas numbers, other binary recurrence sequences are connected with the solutions of Diophantine equations such as $aX^{2}-bY^{4}=c$.
+Such equations are important in their own right, and also as quartic models of elliptic curves.
 This paper is concerned with such problems, in particular how many distinct squares can occur in non-degenerate binary recurrence sequences.
@@
 for all $k\in\mathbb{Z}$. Note that $\left(t^{2}+du^{2}\right)/2=\operatorname{Tr}_{\mathbb{Q}\left(\sqrt{d}\right)/\mathbb{Q}}\left(\varepsilon^{2}\right)$.
-To relate such sequences back to the quadratic equations mentioned in the previous section, observe that from (1.1),
+To relate such sequences back to the quartic equations mentioned in the previous section, observe that from (1.1),
 $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}.$
@@
 If $-N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
-If we replace $\varepsilon^{2k}$ in (1.1) with $\varepsilon^{k}$, it appears that the following is true.
+In fact, a more general result than Conjecture 1.1 also appears to be true.
 Conjecture 1.3.
+Let $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be the sequence formed by replacing $\varepsilon^{2k}$ in (1.1) with $\varepsilon^{k}$.
 There are at most four distinct integer squares among the $y_{k}$’s.
-If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=p$ where $p$ is prime, or if $\left|N_{\alpha}\right|$ is a perfect square, then there are at most three distinct integer squares among the $y_{k}$’s.
-We used PARI/GP [9] to search for squares among $y_{-60},\ldots,y_{60}$ from Conjecture 1.3, where $\varepsilon=\left(t+u\sqrt{d}\right)/2>1$ is the smallest algebraic integer that is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t,u\in\mathbb{Z}$ and $u\leq 10^{30}$.
-Note that this also includes searching among $y_{-30},\ldots,y_{30}$ from Conjectures 1.1 and 1.2.
-The range of $b$ and $d$ in this search was $1\leq b,d\leq 1000$ with $d$ squarefree.
-For $a$, we searched over two ranges: (1) $\max\left(1,\lfloor\sqrt{db^{4}}\rfloor-500\right)\leq a\leq\lceil\sqrt{db^{4}}\rceil+500$ (so $\left|N_{\alpha}\right|$ is small) and (2) $1\leq a\leq 1000$.
+If $\left|N_{\alpha}\right|$ is a prime power or a perfect square, then there are at most three distinct integer squares among the $y_{k}$’s.
+Examples showing that if these conjectures are true, then they are best possible are provided in Subsections 6.1, 6.2 and 6.3.
+We used PARI/GP [10] to search for squares among $y_{-80},\ldots,y_{80}$ from Conjecture 1.3, where $\varepsilon=\left(t+u\sqrt{d}\right)/2>1$ is the fundamental unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t,u\in\mathbb{Z}$ and $u\leq 10^{30}$.
+Note that this also includes searching among $y_{-40},\ldots,y_{40}$ from Conjectures 1.1 and 1.2.
+The range of $b$ and $d$ in this search was $1\leq b\leq 2000$ and $1\leq d\leq 1000$ with $d$ squarefree.
+For $a$, we searched over two ranges: (1) $\max\left(1,\lfloor\sqrt{db^{4}}\rfloor-1000\right)\leq a\leq\lceil\sqrt{db^{4}}\rceil+1000$ (so $\left|N_{\alpha}\right|$ is small) and (2) $1\leq a\leq 2000$.
 For both ranges of $a$, we considered only $\gcd\left(a,b^{2}\right)$ squarefree.
+The search across these ranges took just under $319$ hours on a Windows 11 laptop with an Intel i7-13700H 2.40 GHz processor and 32GB of RAM.
 No counterexamples to any of these conjectures were found in this search.
-The largest value of $|k|$ found in this search such that $y_{k}$ is a square was $k=-42$.
-One such example was found: $y_{-42}=12223^{2}$ for $(a,b,d,t,u)=(502,15,5,1,1)$.
-We also observed that no squares were found among the $y_{k}$’s with $|k|>30$ for $d>5$ and with $|k|>18$ for $d>13$.
 Remark.
@@
 (c) Otherwise, there are at most two distinct squares among the $y_{k}$’s.
+In Subsection 6.4, we present an infinite family of examples of sequences satisfying the conditions in part (a) with $y_{1}$ roughly one-tenth the size of the bounds in Proposition 4.1.
+The same is possible for $y_{-1}$ in part (a) and $y_{\pm 1}$ in part (b).
 1.5. Our method of proof
@@
 The idea is to treat ranges of hypothetical squares beyond those that can be treated by our gap principle alone.
-We have state and prove our results in Section 3 more generally than required here.
+We state and prove our results in Section 3 more generally than required here.
 This is because they will be useful in further work on binary recurrence sequences.
 An example is [18] where we prove Conjecture 1.2 when $N_{\alpha}$ is a prime power.
-When $N_{\alpha}>0$ is not a square, we can apply Thue’s Fundamentaltheorem (see [14] and [15]) to obtain similar results.
 1.6. Structure of this article
-Section 2 concerns results on diophantine approximation, with a focus on the use of hypergeometric functions to do so.
+Section 2 contains results on diophantine approximation, with a focus on the use of hypergeometric functions to do so.
 In Section 3, we collect various facts that we will require about elements of the sequence $\left(y_{k}\right)_{k=-\infty}^{\infty}$ defined by (1.1).
@@
 Finally, in Section 6, we give examples showing that our results and conjectures are best possible.
+1.7. Acknowledgements
+The author deeply appreciates all the time and effort spent by Mihai Cipu, very carefully reading an earlier version of this work.
+His extensive comments, questions and our follow-up discussions led to significant improvements in the presentation of this work, as well as numerous corrections.
+The author is also grateful to the referee for their notes and suggestions, which improved this paper too.
 2. Diophantine Approximation via Hypergeometric Functions
@@
 However, that does not suffice for us to prove our results here.
-The problem here arises not because of the exponent, $\mu$, in the effective irrationality measure, but because the constant, $c$, is too large.
+The problem here arises not because of the exponent, $\mu$, in the effective irrationality measure, but because the constant, $c$, is too small.
 Upon investigating this further, we found that we can complete the proof of our results if we use not the effective irrationality measures from the hypergeometric method, but rather consider more carefully the actual results that we obtain from the use of hypergeometric functions.
@@
 For $d^{\prime}\in\mathbb{Z}$, we define $N_{d^{\prime},n,r}$ to be the largest integer such that $\left(D_{n,r}/N_{d^{\prime},n,r}\right)X_{m,n,r}\left(1-\sqrt{d^{\prime}}\,x\right)\in\mathbb{Z}\left[\sqrt{d^{\prime}}\right][x]$, again for all $m$ as above.
-We will use $v_{p}(x)$ to denote the largest power of a prime $p$ which divides into the rational number $x$.
+We will use $v_{p}(x)$ to denote the largest power of a prime $p$ which divides the rational number $x$.
 We put
@@
 (b) For any positive integer, $r$, we have
-[…]
+(2.3) $\frac{5}{24\cdot 4^{r}r^{1/4}}\leq\frac{(1/4)\cdots(r+1/4)}{(r+1)\cdots(2r+1)}\quad\text{and}\quad\frac{r!\Gamma(3/4)}{\Gamma(r+3/4)}\leq 4r^{1/4}/3.$
 Proof.
@@
 Therefore, we seek a smaller value at the expense of replacing $1.64$ by a larger value, whose value has less of an impact on our proof.
-For $r\geq 156$, $100\exp(1.64r)<0.2\exp(1.68r)$, so we compute directly the left-hand sides of (2.2) for $r\leq 155$.
+For $r\geq 156$, we have $100\exp(1.64r)<0.2\exp(1.68r)$, so we compute directly the left-hand sides of (2.2) for $r\leq 155$.
 We find that the maximum values of the left-hand sides of (2.2) divided by $\exp(1.68r)$ both occur for $r=3$.
@@
 A quick calculation shows that we have equality in both cases for $r=1$.
-Then we take the ratio of the expression for $r+1$ and divide by the expression for $r$.
+Then we take the expression for $r+1$ and divide it by the expression for $r$.
 For the first inequality, this ratio is $(1/4)\left(r+5/4\right)/\left(r+3/2\right)$, so the inequality will hold if $\left(r+5/4\right)/\left(r+3/2\right)>\left(r/\left(r+1\right)\right)^{1/4}$.
-We take the fourth-power of both sides and subtract them, this gives us $\left(224r^{3}+944r^{2}+1329r+625\right)/\left(16\left(2r+3\right)^{4}\left(r+1\right)\right)$, which is clearly positive for $r\geq 1$.
+We take the fourth-power of both sides and subtract them, this gives us
+$\frac{224r^{3}+944r^{2}+1329r+625}{16\left(2r+3\right)^{4}\left(r+1\right)},$
+which is clearly positive for $r\geq 1$.
 Since both $\left(r+5/4\right)/\left(r+3/2\right)$ and $\left(r/\left(r+1\right)\right)^{1/4}$ are positive real numbers for $r\geq 1$, the required inequality hlds.
 We proceed in the very same way to prove the second inequality. ∎
-Put
-$p_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega),$
-$q_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}Y_{1,4,r}(\omega)\hskip 8.53581pt\text{ and}$
-$R_{r}^{\prime}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}R_{1,4,r}(\omega),$
-where $r$ is any non-negative integer and $d^{\prime}$ will be determined below.
-From Lemma 2.2(a), we have
-$q_{r}\omega^{1/4}-p_{r}=R_{r}.$
-By the definitions of $D_{4,r}$ and $N_{d^{\prime},4,r}$, we easily observe that $p_{r}^{\prime}$ and $q_{r}^{\prime}$ are algebraic integers of $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$.
-We can see that as follows.
-$X_{1,4,r}(z)$ is a polynomial of degree $r$ and
-$p_{r}^{\prime}$ $=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega)$
-$=\frac{D_{1,4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
-Since $d^{\prime}=u_{2}^{2}t^{\prime}$, $p_{r}^{\prime}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$.
-The analogous expression for $q_{r}^{\prime}$ shows that it is also an algebraic integer.
-In fact, there may be further common factors. As in [16, Theorem 1], put
+As in [16, Theorem 1], put
 (2.4) $g_{1}$ $=\gcd\left(u_{1},u_{2}\right),$
@@
 $q_{r}\omega^{1/4}-p_{r}=R_{r}.$
-Since $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/g$ is an algebraic integer, the argument above still applies to show that $p_{r}$ and $q_{r}$ are algebraic integers.
+By the definitions of $D_{4,r}$ and $N_{d^{\prime},4,r}$, we see that $X_{1,4,r}(1-z)$ is a polynomial of degree $r$ with coefficients in $\mathbb{Z}\left[\sqrt{d^{\prime}}\right]$ and
+$p_{r}$ $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}(\omega)$
+$=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
+Since $d^{\prime}=u_{2}^{2}t^{\prime}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$.
+The analogous expression for $q_{r}$ shows that it is also an algebraic integer.
 So by applying these quantities, with Lemma 2.2(d) and Lemma 2.3(a), in Lemma 2.1, we can take
@@
 The proposition in this section plays a crucial role in this work.
-It is the expression for $f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ in this lemma that permits us to use the hypergeometric method described in the previous section.
+It is the expression for $f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ in this proposition that permits us to use the hypergeometric method described in the previous section.
 For this reason, we call it our representation proposition.
@@
 (3.1) $x+y^{2}\sqrt{d}=\alpha\varepsilon^{2},$
-where $\varepsilon=\left(t+u\sqrt{d}\right)/2\in\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ non-zero rational integers, norm $N_{\varepsilon}$ and $\gcd\left(u,N_{\varepsilon}\right)=1$.
+where $\varepsilon=\left(t+u\sqrt{d}\right)/2\in\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$ with $t$ and $u$ non-zero rational integers, norm $N_{\varepsilon}$.
 (a) We can write
-[…]
+$f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
 (3.2) $fy$ $=b\left(r^{2}-\operatorname{sf}\left(N_{\alpha}\right)s^{2}\right),$
 for some integers $f$, $r$ and $s$ satisfying $f\neq 0$,
-$f|\left(4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
-where $f^{\prime}>0$, $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $f^{\prime}<\sqrt{\operatorname{sf}\left(\left|N_{\alpha}\right|\right)}$, if $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)>1$ and $f^{\prime}=1$ otherwise.
-(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha},N_{\varepsilon}\right)\right)\right)$.
-However, to do so, the first relationship in part (a) may become
-$-f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
-(c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
+$f|\left(4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
+where $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $0<f^{\prime}<\max\left(2,\sqrt{\left|\operatorname{sf}\left(N_{\alpha}\right)\right|}\right)$.
+(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
+±$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
+(c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
 Remark.
@@
 This makes our representation proposition useful for more general sequences than those considered in this paper.
-The condition that $\gcd\left(u,N_{\varepsilon}\right)=1$ is no restriction. If $g=\gcd\left(u,N_{\varepsilon}\right)$, then $g|t$. By equation (3.1), $g^{2}|x$ and $g^{2}|y^{2}$.
 To prove this proposition, we will use the following three lemmas.
@@
 Expanding the expression for $y^{2}$ in (3.1), we obtain
-(3.4) $4y^{2}=2atu+b^{2}t^{2}+b^{2}du^{2}.$
+(3.4) $4y^{2}=2atu+b^{2}t^{2}+b^{2}du^{2}=2atu+4b^{2}N_{\varepsilon}+2b^{2}du^{2}.$
 Applying this equation, for odd primes, $p$, we have
@@
 So $v_{p}\left(g_{1}^{2}\right)\leq v_{p}\left(s_{1}^{\prime 2}\right)$.
-Now suppose that $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$.
+(2) Now suppose that $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$.
 Put $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$.
-We shall prove the lemma in this case (i.e., $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$) for $r_{1}=r_{1}^{\prime}$.
+We shall prove the lemma in this case (i.e., when $v_{p}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{p}\left(r_{1}^{2}\right)$) for $r_{1}=r_{1}^{\prime}$.
 The proof is identical for $r_{1}=r_{1}^{\prime\prime}$.
@@
 If $k_{2}$ is odd, then we have $2k_{1}-1<k_{2}$.
-So in both cases, $2k_{1}\leq k_{2}$ (i.e., $v_{p}\left(4b^{2}\right)\leq v_{p}\left(r_{1}^{\prime 2}\right)$) and $v_{p}(y)\geq k_{1}$.
+So in both cases, $2k_{1}\leq k_{2}$ (i.e., $v_{p}\left(4b^{2}\right)\leq v_{p}\left(r_{1}^{\prime}\right)$) and $v_{p}(y)\geq k_{1}$.
 We now show that in this case we have $v_{p}\left(4b^{2}\right)<v_{p}\left(r_{1}^{\prime\prime}\right)$.
@@
 As above, we have $v_{2}(au)\geq\min\left(k_{2},v_{2}\left(tb^{2}\pm 2by\right)\right)$.
-Applying this, (3.4) and since $t^{2}+du^{2}=4N_{\varepsilon}+2du^{2}$, we obtain
+Applying this and the right-hand expression in (3.4) for $4y^{2}$, we obtain
 $2\displaystyle 2v_{2}(y)+$ 2 $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
@@
 (b) For all primes, $p$, if $v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)\right)>0$, then
-(3.7) $\min\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq 0\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
-(c) If $r_{1}^{\prime}$ is even, then $v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<0\leq v_{2}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$.
-The same result holds for $r_{1}^{\prime\prime}$.
+(3.7) $\min\left(v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)\leq 0.$
+(c) If $r_{1}^{\prime}$ is even, then $v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<0$. The same result holds for $r_{1}^{\prime\prime}$.
 Proof.
-(a) We will show that for all primes, $p$, we have $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$.
-From the definition of $\operatorname{rad}(\cdot)$, we have $v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\leq 1$.
+(a) We will prove that for all primes, $p$, we have $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$.
+The proof is the same for $r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}$.
 From (3.11), we can write
@@
 Hence at least one of $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)$ or $v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)$ is non-positive, as required.
-(c) This follows immediately from the expression on the right-hand side of (3.8). ∎
+(c) This follows from (3.8).
+We have $v_{2}\left(\operatorname{sf}\left(r_{1}^{\prime}\right)\right)=0$ or $1$.
+In both cases, since $r_{1}^{\prime}$ is even, we find that $v_{2}$ of the denominator of the right-hand side of (3.8) is strictly greater than $v_{2}$ of its numerator.
+∎
 We shall need the following lemma for the proofs of parts (b) and (c) of Proposition 3.1.
@@
 Proof.
-From the expressions for $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$, we need only show that $tb^{2}+au$ is even unless $N_{\alpha}\not\equiv 1\pmod{4}$ and $4\nmid d$.
+From the expressions for $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$, we need only show that $tb^{2}+au$ is even unless $N_{\alpha}\equiv 1\pmod{4}$ and $4\mid d$.
 We do so by showing that if either (i) $tb^{2}$ even and $au$ odd or (ii) $tb^{2}$ odd and $au$ even, then $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
@@
 Again $N_{\alpha}=a^{2}-db^{4}\equiv 1\pmod{4}$, since $a$ must be odd.
-Applying these to the right-hand side of (3.4), we find that $4y^{2}\equiv 2\pmod{4}$, so this case is not possible.
+Applying these to the middle expression in (3.4), we find that $4y^{2}\equiv 2\pmod{4}$, so this case is not possible.
 (ii) If $u$ is even, then since $t$ is odd, $N_{\varepsilon}$ cannot be an integer.
@@
 Hence $\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}\equiv 2\pmod{4}$.
-But this quantity is $b^{2}$ times the right-hand side of (3.4), so it is $4b^{2}y^{2}$.
+But this quantity is $b^{2}$ times the middle expression in (3.4), so it is $4b^{2}y^{2}$.
 Thus (ii) is not possible.
@@
 $\frac{x+N_{\varepsilon}\sqrt{N_{\alpha}}}{a+\sqrt{N_{\alpha}}}=\left(\frac{\left(au+b^{2}t\right)-u\sqrt{N_{\alpha}}}{2b^{2}}\right)^{2}.$
-Applying (3.4), a routine calculation shows that with $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}=-u$, we have
+With $r_{1}=tb^{2}+au\pm 2by$ and $s_{1}=-u$, a routine calculation, along with (3.4), shows that
 $\left(r_{1}+s_{1}\sqrt{N_{\alpha}}\right)^{2}=2r_{1}\left(\left(au+b^{2}t\right)-u\sqrt{N_{\alpha}}\right).$
@@
 From this relationship, we now show that we can find such a relationship for $x+N_{\varepsilon}\sqrt{N_{\alpha}}$ with the conditions on $f$ in part (a) satisfied.
-We want to take the largest common factor, $g_{1}$, of $r_{1}$ and $s_{1}^{\prime}$ such that $4b^{2}r_{1}/g_{1}^{2}$ is also an integer. We have
+We want to take the largest common factor, $g_{1}$, of $r_{1}$ and $s_{1}^{\prime}$ such that $4b^{2}r_{1}/g_{1}^{2}$ is also an integer. That is,
 $g_{1}^{2}=\gcd\left(\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right),s_{1}^{\prime 2}\right).$
@@
 $g_{1}^{2}=\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right).$
-From Lemma 3.3(a), $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of $4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$.
-We put $r=r_{1}/g_{1}$, for either choice of $r_{1}$, $s=s_{1}^{\prime}/g_{1}$ and $f=4b^{2}r_{1}/g_{1}^{2}$.
-By the definition of $g_{1}^{2}$, we see that $f,r,s\in\mathbb{Z}$.
-Lastly, we choose $r_{1}$.
-Let $d^{\prime}$ be the product of all primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$, such that the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ and $d^{\prime\prime}$ be the product of the remaining primes, $p|\operatorname{sf}\left(N_{\alpha}\right)$ (so the minimum in Lemma 3.3(b) is attained for $r_{1}^{\prime\prime}/{g_{1}^{\prime\prime}}^{2}$, but not for $r_{1}^{\prime}/{g_{1}^{\prime}}^{2}$ for such primes).
-If $d^{\prime}>d^{\prime\prime}$, then we let $r_{1}=r_{1}^{\prime}$ and let $r_{1}=r_{1}^{\prime\prime}$ otherwise.
-This choice ensures that $f^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)/\max\left(d^{\prime},d^{\prime\prime}\right)$ satisfies the desired conditions.
-(b) If $f$ is odd, then there is nothing to prove, as $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$ follows from $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$.
-So we need only consider when $f$ is even.
+From Lemma 3.3(a), $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of
+$4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=4b^{2}\operatorname{rad}\left(\operatorname{sf}\left(N_{\alpha}\right)\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right).$
+We put $r=r_{1}/g_{1}$, for either choice of $r_{1}$, $s=s_{1}^{\prime}/g_{1}$ and $f=4b^{2}r_{1}/g_{1}^{2}$. By the definition of $g_{1}^{2}$, we see that $f,r,s\in\mathbb{Z}$.
+Lastly, we choose which of the two possible values of $r_{1}$ to use.
+As in the statement of Lemma 3.3, we let $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$ be the two possibilities.
+We will use Lemma 3.3(b).
+We divide the prime divisors of $\operatorname{sf}\left(N_{\alpha}\right)$ into three disjoint sets:
+$I^{\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
+$I^{\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
+$I^{\prime\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
+With $d^{\prime}=\prod_{p\in I^{\prime}}p$, $d^{\prime\prime}=\prod_{p\in I^{\prime\prime}}p$ and $d^{\prime\prime\prime}=\prod_{p\in I^{\prime\prime\prime}}p$, we have $d^{\prime}d^{\prime\prime}d^{\prime\prime\prime}=\operatorname{sf}\left(N_{\alpha}\right)$.
+We put $r_{1}=r_{1}^{\prime}$ if $d^{\prime}>d^{\prime\prime}$ and $r_{1}=r_{1}^{\prime\prime}$ otherwise. By Lemma 3.3(b),
+$v_{p}\left(r_{1}/g_{1}^{2}\right)\leq 0$ $\leq v_{p}\left(\operatorname{sf}\left(N_{\alpha}\right)/p\right)$
+$\leq v_{p}\left(\operatorname{rad}\left(\left(\operatorname{sf}\left(N_{\alpha}\right)/p\right)\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right),$
+for all $p\in I^{\prime}\cup I^{\prime\prime\prime}$ in the first case and for all $p\in I^{\prime\prime}\cup I^{\prime\prime\prime}$ in the second case.
+Hence with $f^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)/\max\left(d^{\prime}d^{\prime\prime\prime},d^{\prime\prime}d^{\prime\prime\prime}\right)$, we see that $f^{\prime}<\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ and that $4b^{2}r_{1}/g_{1}^{2}$ is a divisor of
+$4b^{2}\operatorname{rad}\left(f^{\prime}\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right).$
+(b) If $4b^{2}r_{1}/g_{1}^{2}$ is odd, then there is nothing to prove, as $4b^{2}r_{1}/g_{1}^{2}$ dividing $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ follows from $4b^{2}r_{1}/g_{1}^{2}$ being a divisor of $4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$.
+So we put $f=4b^{2}r_{1}/g_{1}^{2}$ and need only consider when $4b^{2}r_{1}/g_{1}^{2}$ is even.
 We consider the four cases for the parity of $r$ and $s$.
@@
 Here we will replace $r$ and $s$.
-We can write $(r+si)/(1+i)=((r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
-$-\left(\frac{2b^{2}r}{g_{1}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
-Since $f=4b^{2}r/g_{1}$ is even, we know that $2b^{2}r/g_{1}\in\mathbb{Z}$.
-Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$.
-Hence $2b^{2}r/g_{1}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)$.
+We can write $(r+si)/(1+i)=(r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
+$-\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
+Since $4b^{2}r/g_{1}^{2}$ is even, we know that $2b^{2}r/g_{1}^{2}\in\mathbb{Z}$.
+Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$.
+Hence $2b^{2}r/g_{1}^{2}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ and so we take $f=2b^{2}r/g_{1}^{2}$ in this case.
 (b-ii) Suppose that $r=r_{1}/g_{1}$ is odd and $s=s_{1}^{\prime}/g_{1}$ is even.
+In this case, as well as cases (b-iii) and (b-iv), we will assume that $r_{1}=r_{1}^{\prime}$. The proof is identical using $r_{1}=r_{1}^{\prime\prime}$ instead.
 Since $r$ is odd, we have $v_{2}\left(g_{1}\right)=v_{2}\left(r_{1}\right)$.
@@
 In this case, $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
-If we have equality, then $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and so $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
+If we have equality, then $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and so $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, as required.
 So we consider $v_{2}\left(4b^{2}\right)>v_{2}\left(r_{1}\right)$.
-Since both of these quantities are even, we must have $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)+2$. That is, $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$.
-Since $s$ is even, we have $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)>2v_{2}\left(g_{1}\right)=2v_{2}\left(r_{1}^{\prime}\right)$.
-So $v_{2}\left(r_{1}^{\prime\prime}\right)>v_{2}\left(r_{1}^{\prime}\right)$.
-As we stated near the start of this case ((b-ii)), since $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
+Since both of these quantities are even, we must have $v_{2}\left(4b^{2}\right)\geq v_{2}\left(r_{1}\right)+2$.
+That is, $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$.
+Since $s$ is even, we have
+$v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)>2v_{2}\left(g_{1}\right).$
+Since $r$ is odd, we also have $2v_{2}\left(g_{1}\right)=2v_{2}\left(r_{1}^{\prime}\right)$.
+Hence $v_{2}\left(r_{1}^{\prime\prime}\right)>v_{2}\left(r_{1}^{\prime}\right)$.
+As we stated near the start of this case (case (b-ii)), since $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$, we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
 We saw in the proof of Lemma 3.2 that $v_{2}(y)\geq\min\left(v_{2}(b),v_{2}\left(r_{1}^{\prime}\right)/2\right)$.
 Since $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$, it follows that $v_{2}(y)\geq v_{2}\left(r_{1}^{\prime}\right)/2$.
-Hence $v_{2}(2by)\geq 1+v_{2}\left(r_{1}\right)/2+v_{2}\left(r_{1}\right)/2=v_{2}\left(r_{1}\right)+1$.
-Hence $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)\geq v_{2}\left(r_{1}\right)+1$, which implies that $v_{2}\left(r_{1}\right)>v_{2}\left(r_{1}\right)$.
-Therefore, the case when $v_{2}\left(r_{1}\right)$ is even and $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$ never occurs and we find that in case (b-ii-1), $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
-(b-ii-2) Now suppose that $v_{2}\left(r_{1}\right)$ is odd. Here $v_{2}\left(2b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
+Hence
+$v_{2}(2by)\geq 1+v_{2}\left(r_{1}\right)/2+v_{2}\left(r_{1}\right)/2=v_{2}\left(r_{1}\right)+1.$
+Thus $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)\geq v_{2}\left(r_{1}\right)+1$, which implies that $v_{2}\left(r_{1}\right)>v_{2}\left(r_{1}\right)$.
+Therefore, the case when $v_{2}\left(r_{1}\right)$ is even and $v_{2}\left(b^{2}\right)\geq v_{2}\left(r_{1}\right)$ never occurs and we find that in case (b-ii-1), $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, as required.
+(b-ii-2) Now suppose that $v_{2}\left(r_{1}\right)$ is odd.
+Here $v_{2}\left(2b^{2}\right)\geq v_{2}\left(r_{1}\right)=v_{2}\left(g_{1}\right)$.
 As in case (b-ii-1), since $s$ is even, we have $v_{2}\left(r_{1}^{\prime}\right)\neq v_{2}\left(r_{1}^{\prime\prime}\right)$. Hence $v_{2}\left(au+b^{2}t\right)=v_{2}\left(2by\right)$.
@@
 Since $r$ is even, we must have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right)\right)<v_{2}\left(r_{1}^{2}\right)$.
-Since $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(g_{1}^{2}\right)<2v_{2}\left(r_{1}\right)$.
-Hence $v_{2}\left(r_{1}^{\prime\prime}\right)\neq v_{2}\left(r_{1}\right)$, so by the comment at the start of case (b-ii), we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
+Since $v_{2}\left(s_{1}^{\prime 2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)=v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)$, by the assumption that $s_{1}^{\prime}/g_{1}$ is odd, we have $v_{2}\left(r_{1}^{\prime}\right)+v_{2}\left(r_{1}^{\prime\prime}\right)=v_{2}\left(g_{1}^{2}\right)$.
+Since $r$ is even, we also have $v_{2}\left(g_{1}^{2}\right)<2v_{2}\left(r_{1}\right)$, it follows that $v_{2}\left(r_{1}^{\prime\prime}\right)\neq v_{2}\left(r_{1}\right)$.
+Hence, by the comment at the start of case (b-ii), we have $v_{2}\left(au+b^{2}t\right)=v_{2}(2by)$.
 (b-iii-1) If $v_{2}\left(r_{1}\right)$ is even, then $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4b^{2}r_{1}\right)$.
-Hence $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)$, as required.
+Hence $f=4b^{2}r_{1}/g_{1}^{2}$ is odd and $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$ as required.
 (b-iii-2) If $v_{2}\left(r_{1}\right)$ is odd, then $v_{2}\left(2b^{2}r_{1}\right)<v_{2}\left(r_{1}^{2}\right)$.
@@
 So, as stated at the start of the proof of part (b), we need only consider the latter case.
-Since $s$ is even, we have $2\leq v_{2}\left(s^{2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)-v_{2}\left(2b^{2}\right)$.
+Since $s$ is even, we have
+$2\leq v_{2}\left(s^{2}\right)=v_{2}\left(u^{2}N_{\alpha}\right)-v_{2}\left(2b^{2}r_{1}^{\prime}\right)=v_{2}\left(r_{1}^{\prime\prime}\right)-v_{2}\left(2b^{2}\right).$
 So $v_{2}\left(r_{1}^{\prime\prime}\right)\geq 2v_{2}(b)+3$.
@@
 A consequence of these two inequalities is that $64|\left(u^{2}N_{\alpha}\right)$.
-From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4bN_{\varepsilon}$. So $atu+du^{2}$ must be even. I.e., $atu$ and $du^{2}$ have the same parity.
-If $du^{2}$ is odd, then $a$ and $t$ are also odd.
-Since $4N_{\varepsilon}$ is divisible by $4$, it must be the case that $d\equiv 1\pmod{4}$.
-If $d\equiv 1\pmod{8}$, then $N_{\varepsilon}$ is even.
-Furthermore, since $abd$ is odd, $N_{\alpha}$ is even, so $v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\operatorname{sf}\left(N_{\alpha}\right)\right)\right)\right)\geq 1$ and the desired conclusion holds.
+From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., $atu$ and $b^{2}du^{2}$ have the same parity.
+(b-iv-1) If $b^{2}du^{2}$ is odd, then $a$ and $t$ are also odd. Since $4N_{\varepsilon}=t^{2}-du^{2}$ is divisible by $4$ and $t$ and $u$ are both odd, it must be the case that $d\equiv 1\pmod{4}$.
+If $d\equiv 1\pmod{8}$, then $4N_{\varepsilon}=t^{2}-du^{2}\equiv 0\pmod{8}$ and so $N_{\varepsilon}$ is even.
+Furthermore, since $abd$ is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
 If $d\equiv 5\pmod{8}$, then $N_{\alpha}=a^{2}-db^{4}\equiv 4\pmod{8}$. Since $u$ is odd, we get a contradiction with $64|\left(u^{2}N_{\alpha}\right)$.
-We now consider $du^{2}$ even. Here we will prove that both $uN_{\alpha}$ and $N_{\varepsilon}$ are even. This suffices to prove part (b) in this case.
+(b-iv-2) We now consider $b^{2}du^{2}$ even. If $b$ is even, then
+$v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)\geq 2>v_{2}(f)=1.$
+So the desired conclusion holds.
+Therefore, we may assume that $b$ is odd and $du^{2}$ is even in what follows.
+We showed above that $uN_{\alpha}$ is even.
+We will show here that $N_{\varepsilon}$ is also even.
+These facts will suffice to show that $v_{2}\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)\geq 1=v_{2}(f)$ and hence to prove part (b) in this case.
 To prove that $N_{\varepsilon}$ is even, we will assume that it is odd and obtain a contradiction.
-Since $4N_{\varepsilon}=t^{2}-du^{2}\in\mathbb{Z}$ and $du^{2}$, we have $t$ is even and hence $4|\left(du^{2}\right)$.
+Since $4N_{\varepsilon}=t^{2}-du^{2}\in\mathbb{Z}$ and $du^{2}$ is even, we have $t$ is even and hence $4|\left(du^{2}\right)$.
 If $u$ is odd, then $4|d$.
 Also $N_{\alpha}$ is even, so $a$ is even.
-As a result we find that $v_{2}(2atu),v_{2}\left(2b^{2}du^{2}\right)\geq 3$.
-So under our assumption that $N_{\varepsilon}$ is odd, we have $v_{2}\left(4y^{2}\right)=2$.
+As a result, we find that $v_{2}(2atu),v_{2}\left(2b^{2}du^{2}\right)\geq 3$.
+So under our assumption that $N_{\varepsilon}$ is odd, from (3.4) we have $v_{2}\left(4y^{2}\right)=2$.
 I.e., $y$ is odd.
@@
 In both cases, we find that $y$ is even.
-This contradicts the assumption that $N_{\varepsilon}$ is odd.
-Hence we know that in the case when $du^{2}$ is even, $N_{\varepsilon}$ is even and the assertion in part (b) of the lemma holds.
+This contradicts what we obtained ($y$ is odd) under the assumption that $N_{\varepsilon}$ is odd.
+Hence we know that in the case when $b^{2}du^{2}$ is even, $N_{\varepsilon}$ is even and the assertion in part (b) of the lemma holds.
 (c) We consider two cases.
-(c-i) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=p$, for an odd prime, $p$.
-Then we have $f^{\prime}=1$ in part (a) and so $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$.
+(c-i) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=p$, for a prime, $p$.
+Then we have $f^{\prime}=1$ in part (a) and so $f|\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$.
 Hence the result holds if $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$.
@@
 So, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$.
-With $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$.
+With $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$.
 (c-ii) Suppose that $\left|\operatorname{sf}\left(N_{\alpha}\right)\right|=2p$, where $p$ is an odd prime.
-Here the worst case is when the minimum in (3.7) occurs for one of $r_{1}^{\prime}/g_{1}^{\prime 2}$ and $r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}$ for $2$ and occurs for the other for $p$.
-In this case, we let $r$ be the one of $r_{1}^{\prime}$ or $r_{1}^{\prime\prime}$ for which the minimum in (3.7) for $p$ occurs.
-Thus $v_{p}\left(r/g_{1}^{2}\right)\leq v_{p}\left(\operatorname{rad}\left(N_{\varepsilon}\right)\right)$ for all odd primes, $p$.
-From Lemma 3.4, we see that $r$ is even.
-Hence, from Lemma 3.3(c), we have $v_{2}\left(r/g_{1}^{2}\right)\leq-1$.
-So $f=4b^{2}r/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(N_{\varepsilon}\right)\right)$.
-∎
+From Lemma 3.4, we see that $r_{1}^{\prime}$ and $r_{1}^{\prime\prime}$ are both even.
+Hence, from Lemma 3.3(c), we have
+$\max\left(v_{2}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right),v_{2}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right)<0\leq v_{2}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
+Similar to the end of the proof of part (a), we let $r_{1}=r_{1}^{\prime}$ if $v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\leq v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{2}\right)$ and $r_{1}=r_{1}^{\prime\prime}$ otherwise.
+From Lemma 3.3(b), we find that
+$v_{p}\left(r_{1}/g_{1}^{2}\right)\leq 0\leq v_{p}\left(\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right).$
+So with $f=4b^{2}r_{1}/g_{1}^{2}$, we have $f|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$. ∎
 3.2. Lower bounds for $y_{k}$’s
@@
 (3.12) $y_{k}>\left\{\begin{array}[]{ll}\dfrac{\alpha\varepsilon^{2k}}{2\sqrt{d}}&\text{for $k\geq 0$,}\\ \dfrac{-\overline{\alpha}\,\varepsilon^{2|k|}}{2\sqrt{d}}&\text{for $k<0$.}\end{array}\right.$
-(b) For all $k$, $2y_{k}$ is a positive integer. The sequences $\left(y_{k}\right)_{k\geq 0}$ and $\left(y_{k}\right)_{k\leq K+1}$ are increasing sequences of positive numbers.
+(b) For all $k$, $2y_{k}$ is a positive integer. The sequences $\left(y_{k}\right)_{k\geq 0}$ and $\left(y_{K+1},y_{K},y_{K-1},y_{K-2},\ldots\right)$ are increasing sequences of positive numbers.
 (c) We have
-(3.13) $y_{k}>\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{K-k}&\text{for $k\leq K$.}\end{array}\right.$
+(3.13) $y_{k}\geq\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{\max(0,K-k)}&\text{for $k<0$.}\end{array}\right.$
 In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $52du^{2}/$ 5 by $85du^{2}/$ 8 and if $N_{\varepsilon}=1$, then we can replace $52du^{2}/$ 5 by $du^{2}$.
 Remark.
-If $N_{\alpha}>0$, then $y_{k}<0$ can occur for $k<0$.
+The condition $N_{\alpha}<0$ is needed, since if $N_{\alpha}>0$, then $y_{k}<0$ can occur for $k<0$.
+Also, since $du^{2}\geq 5$ holds under the conditions here, the lower bounds in (3.13) are increasing as $|k|$ increases.
 Proof.
@@
 From the second equality in (3.14), equation (3.13) follows for $k=1$.
-Writing $\varepsilon^{2k}=t_{k}+u_{k}\sqrt{d}$ for $k\geq 1$, we see that $\left(u_{k}\right)_{k\geq 1}$ is an increasing sequence. Hence
-$4b^{2}y_{|K|}$ $=b^{4}\left(t_{|K|}^{2}+du_{|K|}^{2}\right)-2ab^{2}t_{|K|}u_{|K|}=\left(b^{2}t_{|K|}-au_{|K|}\right)^{2}-N_{\alpha}u_{|K|}^{2}$
+Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence.
+Hence, for $k<0$,
+$4b^{2}y_{k}$ $=b^{4}\left(t_{|2k|}^{2}+du_{|2k|}^{2}\right)-2ab^{2}t_{|2k|}u_{|2k|}=\left(b^{2}t_{|2k|}-au_{|2k|}\right)^{2}-N_{\alpha}u_{|2k|}^{2}$
 $\geq\left|N_{\alpha}\right|u^{2}.$
-So equation (3.13) holds for $k=K$.
-Recall the recurrence in (1.3) for $k\geq 0$ and using the monotonicity established in part (b):
+So equation (3.13) holds for $k$ satisfying $K\leq k<0$.
+Recalling the recurrence in (1.3) for $k\geq 0$ and using the monotonicity established in part (b), we obtain
 $y_{k+1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k-1}\geq\left(du^{2}+2N_{\varepsilon}-1\right)y_{k}.$
-If $N_{\varepsilon}=1$, then the stronger form of the desired inequality holds, so we need only consider $N_{\varepsilon}=-1$.
+If $N_{\varepsilon}=1$, then the stronger form of the desired inequality stated after equation (3.13) in the lemma holds, so we need only consider $N_{\varepsilon}=-1$.
 We always have $du^{2}\geq 5$ (with equality only if $(d,t,u)=(5,1,1)$.
-Otherwise, $du^{2}\geq 8$).
+Otherwise, $du^{2}\geq 8$, which we use to establish the other inequality after equation (3.13) in the lemma).
 Hence the desired inequality holds.
-This is not too far from the best possible result from this argument. For $(a,d,t,u)=(1,5,1,1)$ and $b$ arbitrarily large, $y_{2}/y_{1}$ approaches $7/3$ from above, whereas we use $2$ here.
-Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain the desired result for $k\leq K$ too. ∎
+Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain equation (3.13) for $k\leq K$ in the same way. ∎
+We will also need to know when $K$ in Lemma 3.5 is not equal to $-1$ for $b=1$.
+Lemma 3.6.
+[…]
+We have $y_{-1}>1$, except for
+(i) $a\geq 1$, $d=a^{2}+4$, $t=a$ and $u=1$, where $\alpha=2\varepsilon$ and $N_{\alpha}=-4$,
+(ii) $a\geq 1$, $d=a^{2}+1$, $t=2a$ and $u=2$, where $\alpha=\varepsilon$ and $N_{\alpha}=-1$.
+In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, $K=-2$.
+Proof.
+From (3.13), we have $y_{1}\geq\left|N_{\alpha}\right|u^{2}/4$. So if $y_{-1}=1$, we must have either (1) $N_{\alpha}=-4$ and $u=1$, or (2) $N_{\alpha}=-1$ and $u=2$.
+In case (1), we have $d=t^{2}\pm 4$ and $a^{2}-d=-4$. Substituting the expression for $d$ into the second equation, we get $a^{2}-t^{2}=-4\pm 4=0,-8$.
+$a^{2}-t^{2}=-8$ can only occur for $a=1$ and $t=3$ (since we are assuming that $a$ and $t$ are positive integers).
+But in this case, $b^{2}t-au=2$, so $84y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=$ 8.
+Otherwise, we have $t=a$. So $d=a^{2}+4$, $\alpha=2\varepsilon$ and $N_{\varepsilon}=-1$. Here $b^{2}t-au=0$, so $y_{-1}=1$.
+In case (2), we have $d=\left(t^{2}\pm 4\right)/4$ and $a^{2}-d=-1$.
+We proceed similarly, obtaining $84a^{2}-t^{2}=-$ 8 (which is never possible) when $N_{\varepsilon}=1$ and $04a^{2}-t^{2}=$ 0 when $N_{\varepsilon}=-1$.
+Here we have $t=2a$ and $d=a^{2}+1$, so $\alpha=\varepsilon$.
+Once again, $b^{2}t-au=0$, so $y_{-1}=1$.
+A direct calculation, done using Maple, provides the values of $y_{-2}$. ∎
 3.3. Gap Principle
@@
 $\left|\omega-z^{4}\right|=\left|\omega^{1/4}-z\right|\times\prod_{k=1}^{3}\left|\omega^{1/4}-e^{2\pi ik/4}z\right|.$
-Multiplying by $\omega^{-1/4}$ and expanding the resulting expression, the product above equals
+Multiplying by $\omega^{-3/4}$ and expanding the resulting expression, the product above equals
 $\prod_{k=1}^{3}\left|e^{2\pi ik/4-i\theta/4}z-1\right|=\left|e^{3i\varphi}+e^{2i\varphi}+e^{i\varphi}+1\right|,$
@@
 $\left|\omega^{1/4}-z\right|^{2}$ $=\left(\frac{\left|\omega-z^{4}\right|}{\left|e^{3i\varphi}+e^{2i\varphi}+e^{i\varphi}+1\right|}\right)^{2}=\frac{c_{0}^{2}}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}$
-$=\frac{2-\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
+$=\frac{2-2\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
 Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$.
-Part (b) now follows and with inequalities since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$.
+Part (b) now follows and follows with inequalities too, since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$.
 ∎
@@
 […]
-(a) Suppose that $-N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are both squares with $y_{j}>y_{i}\geq 4b^{6}$, then
-$y_{j}>\frac{57.6d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
+(a) Suppose that $-N_{\alpha}$ is a square.
+If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right\}$, then
+$y_{j}>57.32\left(\frac{d}{b^{2}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
 (b) Suppose $-N_{\alpha}$ is not a square.
-If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4b^{6}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
+If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $k_{1},k_{2},k_{3}\neq 0$ and $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
 $y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3},$
 where $f_{i}$ and $f_{j}$ are the values of $f$ in Proposition 3.1 associated with $y_{i}$ and $y_{j}$, respectively.
-If $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then we can replace $15.36$ with $182$.
+If $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right\}$, then we can replace $15.36$ with $182$.
 Proof.
 We start by considering the two parts together and what is common to their proofs.
-By Proposition 3.1, there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
+Since we have assumed that $i\neq 0$ and $j\neq 0$, we can apply Proposition 3.1 to show that there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
 ±$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
@@
 where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$.
-For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $y_{i},y_{j}>1$, we will assume that the $\pm$ on the left-hand side is always $+$. That is,
+Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon}=\pm 1$.
+For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always $+$. That is,
 (3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
@@
 as the argument for the other cases is exactly the same. Subtracting the complex conjugate of one of these equations from the equation itself, we obtain
-$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+(3.18) $\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
 $=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
@@
 (3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ =|±$=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
-$=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}<\frac{2b^{2}}{y_{j}}\leq\frac{1}{2},$
-the last inequality holding because $y_{j}\geq 4b^{6}$.
+$=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}\leq\frac{1}{2},$
+the last inequality holds because $a^{2}+\left|N_{\alpha}\right|=db^{4}$ and $y_{j}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$.
 Let $\zeta_{4}^{(j)}$ be the $4$-th root of unity such that
 (3.20) $\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq k\leq 3}\left|\omega^{1/4}-e^{2k\pi i/4}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
-From (3.19), we can apply Lemma 3.6(b) with $c_{0}=1/2$ to obtain
+From (3.19), we can apply Lemma 3.7(b) with $c_{0}=1/2+0.0001$ to obtain
 $\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.1263.$
@@
 (3.21) $\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.5051\frac{b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}}\,\frac{1}{y_{j}}.$
-In the same way, we define $\zeta_{4}^{(i)}$ for any square $y_{i}>1$ in our sequence and (3.21) also holds with $j$ replaced by $i$ for such $i$.
-Hence, for any two distinct squares, $y_{i},y_{j}\geq 4b^{2}$, among the $y_{k}$’s, we have
-$\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
+In the same way, we define $\zeta_{4}^{(i)}$ for any square $y_{i}$ with $i\neq 0$ in our sequence and (3.21) also holds with $j$ replaced by $i$.
+Hence, for any two distinct squares, $y_{i},y_{j}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, among the $y_{k}$’s, we have
+(3.22) $\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
 $\leq$ $\left|\omega^{1/4}-\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|+\left|\omega^{1/4}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
 The numerator on the right-hand side is
-(3.24) $\begin{array}[]{rl}-2\zeta_{4}^{(i)}y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}&\text{if $\zeta_{4}^{(j)}=\zeta_{4}^{(i)}$,}\\ 2\zeta_{4}^{(i)}x&\text{if $\zeta_{4}^{(j)}=-\zeta_{4}^{(i)}$,}\\ \zeta_{4}^{(i)}\left(1-\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=\sqrt{-1}\zeta_{4}^{(i)}$ and}\\ \zeta_{4}^{(i)}\left(1+\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=-\sqrt{-1}\zeta_{4}^{(i)}$.}\end{array}$
+(3.24) $\begin{array}[]{rl}-2\zeta_{4}^{(i)}y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}&\text{if $\zeta_{4}^{(j)}=\zeta_{4}^{(i)}$,}\\ 2\zeta_{4}^{(i)}x&\text{if $\zeta_{4}^{(j)}=-\zeta_{4}^{(i)}$,}\\ \zeta_{4}^{(i)}\left(1-\sqrt{-1}\right)\left(x-y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=\sqrt{-1}\zeta_{4}^{(i)}$ and}\\ \zeta_{4}^{(i)}\left(1+\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=-\sqrt{-1}\zeta_{4}^{(i)}$.}\end{array}$
 Here we use $\sqrt{-1}$ to denote $\exp\left(2\pi i/4\right)$.
@@
 At this point, our proofs of the two parts of the lemma separate.
-(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that $1+\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
+(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
 $\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
 Since $N_{\alpha}=a^{2}-b^{4}d<0$, we have
-(3.27) $y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3},$
-We can use this gap principle to improve its constant term. Combining (3.27) with $y_{i}\geq 4b^{6}>4b^{2}\left|N_{\alpha}\right|/d$ and $\left|N_{\alpha}\right|<b^{4}d$, we obtain
-$y_{j}>\frac{3.84}{b^{12}}y_{i}^{3}\geq 61.44y_{i},$
+(3.27) $y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
+We can use this gap principle to improve its constant term. Combining (3.27) with $y_{i}\geq b^{2}\left|N_{\alpha}\right|/d$, we obtain
+$y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}\left(\frac{b^{2}\left|N_{\alpha}\right|}{d}\right)^{2}y_{i}=3.84y_{i}.$
 Applying this to (3.26) yields
-$\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\frac{1+1/61.44}{y_{i}}.$
+$\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\frac{1+1/3.84}{y_{i}}.$
 This implies that
-$y_{j}>\frac{57.61}{b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
-completing our proof of part (a).
+$y_{j}>\frac{24.34}{b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}=\frac{24.34d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
+Repeating the process from (3.27) onwards with $3.84$ replaced by $24.34$, we obtain
+$y_{j}>\frac{52.31d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
+Repeating it again with $52.31$ instead of $3.84$, we improve the constant to $56.97$. Repeating it one final time with $56.97$, we obtain the inequality in part (a).
 (b) We again use (3.23) and (3.24).
@@
 Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other.
-Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $i<j$.
+Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $y_{i}<y_{j}$.
 Thus,
@@
 As in the proof of part (a), this gives
-$y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
+$y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3},$
+which establishes the first lower bound for $y_{j}$ in part (b).
 If $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, then this lower bound for $y_{j}$ yields $y_{j}>1.09y_{i}$. Applying this to (3.28), we obtain
@@
 $y_{j}>182\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
-Since $f_{i}f_{j}\leq\left(4b^{2}\right)^{2}\left|N_{\alpha}\right|\leq 16b^{8}d$, if $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, as required.
+Since $f_{i}f_{j}\leq\left(4b^{2}\right)^{2}\left|N_{\alpha}\right|$, if $y_{i}\geq 4.27b^{2}\left|N_{\alpha}\right|^{2}/d$, then $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, as required.
 ∎
@@
 Lemma 3.9.
-(a) Let $y_{k}\geq 4b^{2}$ be a square and put
+[…]
+(a) Let $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ be a square and put
 $\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)=e^{i\varphi}$
@@
 $\left|\varphi\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
-(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4b^{2}$, $\omega$ be as in part (a) and put $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1.
+(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega$ be as in part (a) and put
+$x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$
+with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1.
 Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.8 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$.
@@
 From (1.1) and $N_{\alpha}>-b^{4}d$, we have
-$x_{k}^{2}+N_{\alpha}$ $=dy_{k}^{2}+2N_{\alpha}>d\left(y_{k}^{2}-2b^{4}\right)\geq dy_{k}^{2}\left(1-\frac{2b^{4}}{y_{k}^{2}}\right)$
-$\geq 0.875dy_{k}^{2}>0.875x_{k}^{2},$
-since $y_{k}\geq 4b^{2}$ and hence $1-2b^{4}/y_{k}^{2}\geq 1-2b^{4}/\left(16b^{12}\right)\geq 0.875$.
+$x_{k}^{2}+N_{\alpha}=dy_{k}^{2}+2N_{\alpha}=dy_{k}^{2}\left(1+\frac{2N_{\alpha}}{dy_{k}^{2}}\right)\geq 0.875dy_{k}^{2}>0.875x_{k}^{2},$
+since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ and $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}<0$.
 From $|\varphi|\leq|\tan(\varphi)|$, we have
 $|\varphi|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
-Since $y_{k}\geq 4b^{2}$, we have $dy_{k}^{2}\geq 16db^{4}>16\left|N_{\alpha}\right|$ and so
-$x_{k}^{2}=dy_{k}^{2}+N_{\alpha}>15\left|N_{\alpha}\right|,$
-which yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
+Since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have $dy_{k}^{2}\geq 16\left|N_{\alpha}\right|$ and so
+$x_{k}^{2}=dy_{k}^{2}+N_{\alpha}\geq 15\left|N_{\alpha}\right|.$
+Combining this with the inequality above it yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
 (b) Recall that from (3.18) in the proof of Lemma 3.8, we can write
-$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{N_{\alpha}}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{N_{\alpha}}\right)^{4}=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
+$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+=±$=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
 for $i=k,\ell$. There are other cases according to the signs in the result in Proposition 3.1, but the argument for them is identical to the argument that follows below.
 Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
-$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
-$\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)}{f_{k}^{2}}$
-$f_{k}^{2}y_{k}^{2}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)$.
-Applying equation (3.3) with $j=\ell$ and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
-(3.29) $|f(x,y)|$ $=\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
-$=2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}\sqrt{\left|N_{\alpha}\right|}.$
+$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+$-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+$=$ $\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}}{f_{k}^{2}}\left[\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
+$\hskip 93.89409pt\left.-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right]$
+$=$ $\frac{f_{k}^{2}y_{k}^{2}}{b^{4}}\left[\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
+$\hskip 28.45274pt\left.-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right].$
+Applying equation (3.18) with $j=\ell$ to the last expression and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
+(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
+$\left.-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
+$=$ $\frac{2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}}\sqrt{\left|N_{\alpha}\right|}.$
 Let $\zeta_{4}$ be the $4$-th root of unity satisfying
@@
 From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (1.1) (which implies that $\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
-(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ $=\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
-$=\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}<\frac{2b^{2}}{y_{\ell}}\leq\frac{1}{2}.$
-since $y_{\ell}\geq 4b^{2}$.
-By Lemma 3.6(b) with $c_{0}=1/2$,
+(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
+$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
+$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}\leq\frac{1}{2}.$
+since $y_{\ell}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$.
+By Lemma 3.7(b) with $c_{0}=1/2+0.0001$,
 (3.31) $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
-From (3.3), along with $y_{k},y_{\ell}\geq 4b^{2}$, we have
+From (3.22), along with $y_{k},y_{\ell}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have
 $\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|$ $=\left|\zeta_{4}^{(k)}\frac{r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(\ell)}\frac{r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
 Recalling (3.31), it follows that
-$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|=\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
+$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
+$=$ $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
 $\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
-for $\zeta_{4}^{\prime}\neq\zeta_{4}$.
+for any $4$-th root of unity, $\zeta_{4}^{\prime}$, with $\zeta_{4}^{\prime}\neq\zeta_{4}$.
 Since this exceeds $0.403$, it follows that $\zeta_{4}=\zeta_{4}^{(k)}/\zeta_{4}^{(\ell)}=\pm 1$, the last equality holding by our assumption in the statement of this lemma.
@@
 In Lemma 3.10, we establish a gcd result for elements of a generalisation of our sequences.
-This will help us prove Lemmas 3.10 and 3.11.
+This will help us prove Lemmas 3.11 and 3.12.
 Lemma 3.12 is used in the proof of Proposition 4.1, as well as in the proof of Theorem 1.4.
@@
 Lemma 3.10.
-Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon\neq\pm 1$ is a unit in $\mathbb{Q}\left(\sqrt{d}\right)$.
+Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon=\left(t+u\sqrt{d}\right)/2\neq\pm 1$, with $t$ and $u$ integers, is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$.
 If $x_{k}$ and $y_{k}$ are both integers, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
 Proof.
-We write $\varepsilon=\left(t+u\sqrt{d}\right)/2$, where $t$ and $u$ are integers.
 We may assume without loss of generality that $\gcd(a,b)=1$.
 Otherwise, consider below the sequences $\left(x_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ and $\left(y_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ instead of $\left(x_{k}\right)_{k=0}^{k=\infty}$ and $\left(y_{k}\right)_{k=0}^{k=\infty}$.
-We show that
+We start with some relationships that we will require for the proof itself.
+Throughout the proof, we will use the fact that both the $x_{k}$’s and $y_{k}$’s satisfy the recurrence relation
+(3.32) $u_{k}=\operatorname{Tr}\left(\varepsilon\right)u_{k-1}-N_{\varepsilon}u_{k-2}=tu_{k-1}\pm u_{k-2}.$
+Next, we show that
 (3.33) $ux_{k}-ty_{k}=-2N_{\varepsilon}y_{k-1}\quad\text{ and }\quad tx_{k}-duy_{k}=2N_{\varepsilon}x_{k-1}$
@@
 Similarly, we can prove the relationship holds for $k=2$.
-For $k\geq 3$, we use induction and the recurrence relations for $x_{k}$ and $y_{k}$:
+For $k\geq 3$, we use induction and the recurrence relations in (3.32):
 $ux_{k}-ty_{k}$ $=u\operatorname{Tr}\left(\varepsilon\right)x_{k-1}-uN_{\varepsilon}x_{k-2}-t\operatorname{Tr}\left(\varepsilon\right)y_{k-1}+tN_{\varepsilon}y_{k-2}$
@@
 for $k\geq 2$ and
-(3.34) $\left(\frac{u\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=-N_{\varepsilon}y_{k-3}\quad\text{ and }$
-$\left(\frac{t\left(t^{2}+3du^{2}\right)}{8N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{du\left(3t^{2}+du^{2}\right)}{8N_{\varepsilon}^{2}}\right)y_{k}$ $=N_{\varepsilon}x_{k-3},$
+(3.35) $\left(\frac{u\left(3t^{2}+du^{2}\right)}{4N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{t\left(t^{2}+3du^{2}\right)}{4N_{\varepsilon}^{2}}\right)y_{k}$ $=-2N_{\varepsilon}y_{k-3}\quad\text{ and }$
+$\left(\frac{t\left(t^{2}+3du^{2}\right)}{4N_{\varepsilon}^{2}}\right)x_{k}-\left(\frac{du\left(3t^{2}+du^{2}\right)}{4N_{\varepsilon}^{2}}\right)y_{k}$ $=2N_{\varepsilon}x_{k-3},$
 for $k\geq 3$, which both follow from (3.33).
-Since $3t^{2}+du^{2}=-N_{\varepsilon}+4t^{2}\equiv 4+4\equiv 0\pmod{8}$ and $t^{2}+3du^{2}=N_{\varepsilon}+4u^{2}\equiv 4+4\equiv 0\pmod{8}$, all the coefficients of $x_{k}$ and $y_{k}$ in the relationships in (3.34) are integers.
+Writing $3t^{2}+du^{2}=-4N_{\varepsilon}+4t^{2}$ and $t^{2}+3du^{2}=4N_{\varepsilon}+4du^{2}$, we see that the coefficients of $x_{k}$ and $y_{k}$ in the relationships in (3.35) are integers.
+We divide the remainder of the proof into cases according to the parity of $t$ and $u$.
 (i) Suppose that $t$ and $u$ are both even.
@@
 (ii-a) Suppose that $a$ and $b$ have opposite parities. In this case, $y_{1}=(au+bt)/2$ cannot be an integer. Similarly, $2x_{1}$ is odd.
-Since $\left(\left(t+u\sqrt{d}\right)/2\right)^{2}=\left(t^{2}+u^{2}d\right)/4+(tu/2)\sqrt{d}=\left(t_{2}+u_{2}\sqrt{d}\right)/2$ with $t_{2}$ and $u_{2}$ both odd, we see that $x_{2}$ and $y_{2}$ are not integers either.
-However, $\left(\left(t+u\sqrt{d}\right)/2\right)^{3}=\left(t^{3}+3tu^{2}d\right)/8+\left(\left(3t^{2}u+u^{3}d\right)/8\right)\sqrt{d}=\left(t_{3}+u_{3}\sqrt{d}\right)/2$ with both $t_{3}$ and $u_{3}$ even.
+Since $\left(\left(t+u\sqrt{d}\right)/2\right)^{2}=\left(t^{2}+u^{2}d\right)/4+(tu/2)\sqrt{d}=\left(t_{2}+u_{2}\sqrt{d}\right)/2$, where $t^{2}+u^{2}d=\pm 4+2du^{2}\equiv 2\pmod{4}$, it follows that $t_{2}$ and $u_{2}$ are both odd.
+Hence $x_{2}$ and $y_{2}$ are not integers.
+However,
+$\left(\frac{t+u\sqrt{d}}{2}\right)^{3}=\frac{t^{3}+3tu^{2}d}{8}+\frac{3t^{2}u+u^{3}d}{8}\sqrt{d}=\frac{t_{3}+u_{3}\sqrt{d}}{2},$
+where $t^{3}+3tu^{2}d=t\left(t^{2}+3u^{2}d\right)=t\left(\pm 4+4u^{2}d\right)\equiv 0\pmod{8}$ and $3t^{2}u+u^{3}d=-u\left(-4t^{2}\pm 4\right)\equiv 0\pmod{8}$.
+So both $t_{3}$ and $u_{3}$ are even.
 Hence $x_{3}$ and $y_{3}$ are both integers.
-By an inductive argument, one finds that $x_{k}$ and $y_{k}$ are both integers if and only if $3|k$, when $t$ and $u$ are both odd and $a$ and $b$ have opposite parities.
+By an inductive argument using the recurrence relations for the $x_{k}$’s and $y_{k}$’s in (3.32), one finds that $x_{k}$ and $y_{k}$ are both integers if and only if $3|k$, when $t$ and $u$ are both odd and $a$ and $b$ have opposite parities.
 So we consider only $\gcd\left(x_{3k},y_{3k}\right)$.
-By (3.34), we have $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-3},y_{k-3}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1$.
-(ii-b) If $a$ and $b$ have the same parity, then $y_{1}=(au+bt)/2\in\mathbb{Z}$.
-Similarly, $x_{1}\in\mathbb{Z}$.
-Using the recurrence relations, we find that all the $x_{k}$’s and $y_{k}$’s are integers.
-Using (3.34), we find that if $k\equiv k_{1}\pmod{3}$ with $0\leq k_{1}\leq 2$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k_{1}},y_{k_{1}}\right)$.
+Since $t$ and $du$ are both odd in this case, we have $3t^{2}+du^{2}=-4N_{\varepsilon}+4t^{2}\equiv 0\pmod{8}$ and $t^{2}+3du^{2}=4N_{\varepsilon}+4du^{2}\equiv 0\pmod{8}$, so the coefficients of $x_{k}$ and $y_{k}$ in (3.35) can all be divided by $2$ to eliminate the factors of $2$ on the right-hand sides there.
+Thus, $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-3},y_{k-3}\right)|\cdots|\gcd\left(x_{0},y_{0}\right)=1$.
+(ii-b) If $a$ and $b$ have the same parity, then $y_{1}=(au+bt)/2\in\mathbb{Z}$. Similarly, $x_{1}\in\mathbb{Z}$. Using the recurrence relations in (3.32), the $x_{k}$’s and $y_{k}$’s are integers.
+As in case (ii-a), we can remove the factor of $2$ on the right-hand side of (3.35).
+So we find that if $k\equiv k_{1}\pmod{3}$ with $0\leq k_{1}\leq 2$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k_{1}},y_{k_{1}}\right)$.
 So if $3|k$, then $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right)$.
@@
 We now consider the parity of $x_{k}$ and $y_{k}$. Since $\gcd(a,b)=1$, at least one of $x_{0}$ or $y_{0}$ is odd.
-Suppose first that $y_{0}$ is odd. Since $t$ is odd, we find that $y_{2}=ty_{1}+y_{0}$ is even if $y_{1}$ is odd and is odd otherwise.
-In the first case (when $y_{1}$ is odd), we have $\gcd\left(x_{1},y_{1}\right)|\left(\gcd\left(x_{0},y_{0}\right)\right)$ and hence
+Suppose first that $y_{0}$ is odd. Since $t$ is odd, we find that $y_{2}=ty_{1}\pm y_{0}$ is even if $y_{1}$ is odd and $y_{2}$ is odd otherwise.
+In the first case (when $y_{1}$ is odd), we have $\gcd\left(x_{1},y_{1}\right)|\gcd\left(x_{0},y_{0}\right)$ and hence
 $\gcd\left(x_{2},y_{2}\right)|\left(2\gcd\left(x_{1},y_{1}\right)\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
@@
 $\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
-Now we suppose that $x_{0}$ is odd. Since $t$ is odd, we find that $x_{2}=tx_{1}+x_{0}$ is even if $x_{1}$ is odd and is odd otherwise. As above, we find that
+Now we suppose that $x_{0}$ is odd. Since $t$ is odd, we find that $x_{2}=tx_{1}+x_{0}$ is even if $x_{1}$ is odd and $x_{2}$ is odd otherwise. As above, we find that
 $\gcd\left(x_{2},y_{2}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right).$
-(iii) Now suppose that $t$ is even and $u$ is odd. In this case, $4|d$. In fact, if $4|t$, then $d\equiv 4\pmod{8}$ since $\varepsilon$ is a unit, and if $t\equiv 2\pmod{4}$, then $d\equiv 0\pmod{8}$.
+(iii) Now suppose that $t$ is even and $u$ is odd. In this case, $4|d$.
 Since $y_{1}=(au+bt)/2$, we see that $y_{1}\in\mathbb{Z}$ if and only if $a$ is even.
 Hence all the $y_{k}$’s are integers if $a$ is even.
-Since $t$ is even, using the recurrence relation, we see that for $a$ odd, $y_{k}\in\mathbb{Z}$ when $k$ is even and $y_{k}\not\in\mathbb{Z}$ otherwise.
-Since $x_{1}=(at+bdu)/2$, we see that $x_{1}\in\mathbb{Z}$. Hence by the recurrence relation, $x_{k}\in\mathbb{Z}$ always holds.
-As in case (ii) above, we use (3.34). If $k\equiv 0\pmod{3}$, then
-$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right).$
-If $k\equiv 1\pmod{3}$, then
-$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{1},y_{1}\right)&\text{if $a$ is even},\\ \gcd\left(x_{4},y_{4}\right)&\text{if $a$ is odd}.\end{array}\right.$
-We need to stop at $\gcd\left(x_{4},y_{4}\right)$ when $a$ is odd because $y_{1}$ is not an integer in this case.
-By (3.32), we see that the lemma holds if $a$ is even.
-By (3.33) and observing that $2|t$ and $4|d$, we see that $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{k-2},y_{k-2}\right)$.
-So $\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{0},y_{0}\right)$ when $a$ is odd.
-If $k\equiv 2\pmod{3}$, then from (3.32) and (3.33), along with the observation when $k\equiv 1\pmod{3}$ and $a$ is odd, we have
-$\gcd\left(x_{k},y_{k}\right)|\gcd\left(x_{2},y_{2}\right)|\gcd\left(x_{0},y_{0}\right).$
+Since $t$ is even, using the recurrence relation for the $y_{k}$’s in (3.32), we see that for $a$ odd, $y_{k}\in\mathbb{Z}$ when $k$ is even and $y_{k}\not\in\mathbb{Z}$ otherwise.
+Since $x_{1}=(at+bdu)/2$ and $t$ and $d$ are both even, we see that $x_{1}\in\mathbb{Z}$. Hence, by the recurrence relation for the $x_{k}$’s in (3.32), $x_{k}\in\mathbb{Z}$ always holds.
+Here we use (3.34), observing that the coefficients of $x_{k}$ and $y_{k}$ on the left-hand sides are all even.
+So we have relationships with integer coefficients and without the factor of $2$ on the right-hand sides.
+Hence
+$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{0},y_{0}\right)=1&\text{if $k$ is even},\\ \gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right)=2&\text{if $k$ is odd}.\end{array}\right.$
 (iv) We cannot have $t$ odd and $u$ even because $t^{2}-du^{2}\equiv 0\pmod{4}$. ∎
@@
 Lemma 3.11.
-Let the sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be as defined in (1.1).
-Suppose that $k\neq 0$, $\gcd(a,b)=1$ and that $x_{k}$ and $y_{k}$ are both integers.
+Let the sequences $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be as defined in (1.1), with the notation and assumptions there.
+Suppose that $k\neq 0$, $\gcd(a,b)=1$, $N_{\alpha}<0$ and that $x_{k}$ and $y_{k}$ are both integers.
 Using the notation of Subsection 2.1 with $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$, $u_{1}=2x_{k}$ and $u_{2}=\pm 2\sqrt{N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)}$, then
@@
 We also put $v_{p}\left(g_{1}^{2}\right)=\ell_{1}$ and $v_{p}\left(g_{2}\right)=\ell_{2}$.
-Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (recall from (2.4) that $g_{2}|t^{\prime}$ and $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$ here).
-From (2.4), we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$.
+Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (from the definition of $g_{2}$, we have $g_{2}|t^{\prime}$ and here $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$).
+From the definition of $g_{3}$, we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$.
 Hence $\ell=\ell_{1}+\ell_{2}$.
@@
 If $p\mid\operatorname{sf}\left(N_{\alpha}\right)$, then $p^{\ell_{1}+1}\parallel N_{\alpha}$.
-We must have $\ell_{2}=0$, since otherwise $\ell_{1}+1<\ell_{1}$ and we get a contradictions to $p^{\ell+1}|N_{\alpha}$.
+We must have $\ell_{2}=0$, since otherwise $\ell_{1}+1<\ell_{1}$ and we get a contradiction to $p^{\ell+1}|N_{\alpha}$.
 In this case, $p\nmid\left(x_{k}^{2}/g_{1}^{2}\right)$.
@@
 We have $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}\right)\right)\geq\ell+2$.
-Also $g_{2}=1$.
+Also $v_{2}\left(g_{2}\right)=1$.
 Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}/g_{3}\right)\geq\ell$, recalling that $g_{3}=1,2$ or $4$.
-(2.2) Suppose that $2|\left(N_{\alpha}\right)$.
-Again, we break this case into subcases.
-(2.2.1) If $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then we saw above that $v_{2}\left(N_{\alpha}\right)\geq\ell+1$.
+(2.2) Suppose that $2|\operatorname{sf}\left(N_{\alpha}\right)$.
+Here $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(\gcd\left(4x_{k}^{2},4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+1$.
+We break this case into subcases.
+(2.2.1) If $v_{2}\left(a^{2}\right)=v_{2}\left(db^{4}\right)=\ell$, then $v_{2}\left(N_{\alpha}\right)\geq\ell+1$.
 So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4\gcd\left(x_{k}^{2},N_{\alpha}/2\right)\right)\geq\ell+2$.
-Here $g_{2}\geq 1$, so as in case (2.1), we have $v_{2}\left(g^{2}\right)\geq\ell$, as required.
+Here $v_{2}\left(g_{2}\right)\geq 1$, so as in case (2.1), we have $v_{2}\left(g^{2}\right)\geq\ell$, as required.
 (2.2.2) If $v_{2}\left(a^{2}\right)\neq v_{2}\left(db^{4}\right)$, then $v_{2}\left(N_{\alpha}\right)=\ell$.
-So $v_{2}\left(g_{1}^{2}\right)=\ell+1$.
-Since $2|\operatorname{sf}\left(N_{\alpha}\right)$, it follows that $\ell$ must be odd.
+So $v_{2}\left(g_{1}^{2}\right)=v_{2}\left(4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)=\ell+1$, since $2|\operatorname{sf}\left(N_{\alpha}\right)$.
+From $v_{2}\left(g_{1}^{2}\right)=\ell+1$, it also follows that $\ell$ must be odd.
 From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$ and $v_{2}\left(N_{\alpha}\right)=\ell$, we have either $v_{2}\left(x_{k}^{2}\right)=\ell$ and $v_{2}\left(dy_{k}^{2}\right)\geq\ell+1$; or else $v_{2}\left(x_{k}^{2}\right)\geq\ell+1$ and $v_{2}\left(dy_{k}^{2}\right)=\ell$.
@@
 In the latter case, we must have $v_{2}(d)=\ell$ and $2\nmid y_{k}$.
-Also, $v_{2}\left(u_{1}/g_{1}\right)=(1/2)v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)\geq(1/2)\left(\ell+3-(\ell+1)\right)=1$, so $v_{2}\left(g_{2}\right)=1$.
-Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}g_{2}/g_{3}\right)=v_{2}\left(g_{1}^{2}g_{2}/4\right)=\ell$.
-∎
+Also,
+$v_{2}\left(u_{1}/g_{1}\right)=(1/2)v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)\geq(1/2)\left(\ell+3-(\ell+1)\right)=1,$
+so $v_{2}\left(g_{2}\right)=1$. Hence $v_{2}\left(g^{2}\right)=v_{2}\left(g_{1}^{2}g_{2}/g_{3}\right)=v_{2}\left(g_{1}^{2}g_{2}/4\right)=\ell$. ∎
 Lemma 3.12.
@@
 From that lemma, we can write $g^{2}=2^{m}\gcd\left(a^{2},db^{4}\right)$, where $m\geq 0$.
-So, squaring both sides of the equality in (3.35), applying the definitions of $b^{\prime}$ and $d^{\prime}$ and simplifying, it is equivalent to
+From (2.1) with $n=4$, we find that $\mathcal{N}_{d^{\prime},4}^{2}=2^{\min\left(v_{2}(d^{\prime}),6\right)}$.
+So, squaring both sides of the equality in (3.36), it is equivalent to
+$2^{\min\left(6+m,v_{2}(d^{\prime})+m\right)}=2^{\min\left(6,v_{2}(4b^{\prime})\right)}.$
+From the definition of $d^{\prime}$ and Lemma 3.11, we have $d^{\prime}=4N_{\alpha}/g^{2}=4N_{\alpha}/\left(2^{m}\gcd\left(a^{2},db^{4}\right)\right)$ Applying this, along with the definition of $b^{\prime}$, it follows that the equality in (3.36) is equivalent to
 $\min\left(6+m,v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)\right)=\min\left(6,v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)\right).$
@@
 From Lemma 3.10, we have $\gcd\left(x_{k},y_{k}\right)=1,2$.
-We will show that $\gcd\left(x_{k},y_{k}\right)=2$ is not possible here.
-From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $\left(x_{k}/2^{\ell/2+1}\right)^{2}-\left(d/2^{\ell}\right)\left(y_{k}/2\right)^{2}N_{\alpha}/2^{\ell+2}$.
-Since $v_{2}\left(N_{\alpha}\right)\geq\ell+5$, it follows that $x_{k}/2^{\ell/2+1}$ and $y_{k}/2$ are both odd.
+We will show that $y_{k}$ is odd and hence $\gcd\left(x_{k},y_{k}\right)=2$ is not possible here.
+We assume that $y_{k}$ is even and derive a contradiction.
+From $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $\left(x_{k}/2^{\ell/2+1}\right)^{2}-\left(d/2^{\ell}\right)\left(y_{k}/2\right)^{2}=N_{\alpha}/2^{\ell+2}$.
+Since $v_{2}\left(4N_{\alpha}/\gcd\left(a^{2},db^{4}\right)\right)>6$, we have $v_{2}\left(N_{\alpha}\right)\geq\ell+5$.
+It follows that $x_{k}/2^{\ell/2+1}$ and $y_{k}/2$ are both odd.
 Arguing mod $8$, we find that $d\equiv 1\pmod{8}$ and that $t$ and $u$ must both be even.
@@
 Since $t$ and $u$ are both even, it follows that $v_{2}\left(4y_{k}\right)=v_{2}\left(2du^{2}\pm 4\right)=2$.
-If $a$ is even, then $b$ is odd. So if $u$ is also odd, then $4|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=1$. Hence $2\nmid y_{k}$.
-If $u$ is even, then $16|(2atu)$, while $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$. Hence $2\nmid y_{k}$.
-We saw above that $\gcd\left(a^{2}\right)=\gcd\left(db^{4}\right)$, so if $a$ is odd, then so is $b$.
-Hence $v_{2}\left(b^{2}\left(2du^{2}\pm 4\right)\right)=2$, since $u$ is even.
-Here $8|(2atu)$, so $v_{2}\left(4y_{k}\right)=2$.
-In none of these cases did we find that $2|y_{k}$. So
-$v_{2}\left(g_{1}^{2}\right)=\min\left(v_{2}\left(4x_{k}^{2},4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4x_{k}^{2}\right)=\ell+2.$
+This contradicts the assumption that $y_{k}$ is even.
+From $y_{k}$ is odd and $\left(x_{k}/2^{\ell/2}\right)^{2}-\left(d/2^{\ell}\right)y_{k}^{2}=N_{\alpha}/2^{\ell}$, we have
+$v_{2}\left(g_{1}^{2}\right)=\min\left(v_{2}\left(4x_{k}^{2}\right),v_{2}\left(4N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)\right)=v_{2}\left(4x_{k}^{2}\right)=\ell+2.$
 This also implies that $v_{2}\left(4x_{k}^{2}/g_{1}^{2}\right)=0$, so $v_{2}\left(g_{2}\right)=0$.
@@
 Lemma 3.13.
-Suppose that $b=1$, $k\neq 0$, that $x_{k}$ and $y_{k}$ are both integers and that $g$ and $\mathcal{N}_{d^{\prime},4}$ are as above.
+Suppose that $b=1$, $x_{k}$ and $y_{k}$ are both integers with $y_{k}>1$ and that $g$ and $\mathcal{N}_{d^{\prime},4}$ are as above.
 If $d\geq 105$, then
@@
 (3.38) $\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>217.$
+Remark.
+Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
 Proof.
 We first obtain an analytic lower bound for $d$ such that (3.37) and (3.38) hold for larger $d$.
-We have $|g|\mathcal{N}_{d^{\prime},4}\geq 2^{1+\min\left(v_{2}(N_{\alpha})/2,2\right)}\geq 2$ from (3.35). Since $b=1$, from Lemma 3.5(c), we find that $y_{k}>\left|N_{\alpha}\right|/4$ holds, so
+We have $|g|\mathcal{N}_{d^{\prime},4}\geq 2^{1+\min\left(v_{2}(N_{\alpha})/2,2\right)}\geq 2$ from (3.36). Since $b=1$, from Lemma 3.5(c), we find that $y_{k}\geq\left|N_{\alpha}\right|/4$ holds, so
 $\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}>0.091\sqrt{d}.$
@@
 Similarly,
-$\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>\frac{21.12\sqrt{d}\,\left(\left|N_{\alpha}\right|/4\right)}{8\sqrt{\left|N_{\alpha}\right|}}\geq 0.66\sqrt{d},$
-where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.35) to establish the first inequality. For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
+$\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}\geq\frac{21.12\sqrt{d}\,\left(\left|N_{\alpha}\right|/4\right)}{8\sqrt{\left|N_{\alpha}\right|}}\geq 0.66\sqrt{d},$
+where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.36) and Lemma 3.5(c) with $u\geq 1$ to establish the first inequality.
+For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
 We complete the proof computationally, checking all the remaining pairs, $(a,d)$. Since $d$ is bounded and $N_{\alpha}<0$ (so $a^{2}<d$), there are only finitely many such pairs.
@@
 No counterexamples to (3.37) and (3.38) were found.
-The script took 112 seconds to run on a laptop with an Intel i7-9750H processor and 16gb of RAM.
+A PARI/GP program took 112 seconds to run on a Windows 10 laptop with an Intel i7-9750H processor and 16GB of RAM.
 The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, $k=-1$, $x_{k}=-59$ and $y_{k}=5$.
 The smallest value of the left-hand side of (3.38) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, $k=-1$, $x_{k}=-130$ and $y_{k}=11$.
-Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$. ∎
 4. Proposition 4.1 and its proof
@@
 $y_{k}>\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right).$
-Conjecture 1.2 would follow for all sequences with $b=1$ and $-N_{\alpha}$ a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$.
+Conjecture 1.2 would immediately follow when $-N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$.
 This is due to Lemma 3.5.
@@
 We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$.
+So $y_{\ell}>y_{k}\geq 4\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right\}$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
 We shall initially assume that
 (4.1) $d\geq 105.$
-This arises from Lemma 3.12 and will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof, in Subsection 4.6.
+This is the condition in Lemma 3.13, which will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof in Subsection 4.6.
 As in Lemma 3.9, we put $\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
@@
 From (3.31) in the proof of Lemma 3.9(b), we have
-$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.06.$
-Thus we can apply Lemma 3.6(a) with $c_{1}=0.06$ to find that
-(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
+$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
+Thus we can apply Lemma 3.7(a) with $c_{1}=0.127$ to find that
+(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
 The equality on the left-hand side is from the equalities in (3.30).
@@
 Substituting these quantities along with $\mathcal{D}_{4}=e^{1.68}$ from Lemma 2.3(a) into the definition of $E$ in (2.13) and applying (4.3), we have
-(4.5) $E$ $=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}>\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{4e^{1.68}\left|N_{\alpha}\right|}$
+(4.5) $E$ $=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{4e^{1.68}\left|N_{\alpha}\right|}$
 $>\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|x_{k}\right|+\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{10.74\left|N_{\alpha}\right|}>\frac{|g|\mathcal{N}_{d^{\prime},4}\left(1+\sqrt{0.9375}\right)\sqrt{d}\,y_{k}}{10.74\left|N_{\alpha}\right|}$
 $>\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}.$
-From Lemma 3.5(c), $|g|\mathcal{N}_{d^{\prime},4}\geq 2$ and (4.1), we have $E>1$, as required for its use with Lemma 2.1.
-Similarly, using (4.3) and (3.37) in Lemma 3.12 we have
+From (3.38) in Lemma 3.13, we have $E>1.13>1$, as required for its use with Lemma 2.1.
+Similarly, using (4.3) and (3.38) in Lemma 3.13, we have
 (4.6) $Q>\frac{2e^{1.68}\left(1+\sqrt{0.9375}\right)\sqrt{d}y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>217,$
@@
 Recall from (2.12) that we take $k_{0}=0.89$.
-Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.8(a), we can take
-(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.46\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
+Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
+(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
 Also from Lemma 3.9(a), we have $|\varphi|<0.6$, so the condition $|\omega-1|<1$ in Lemma 2.2 is satisfied.
@@
 We start by determining an upper bound for $y_{\ell}$ for all $r_{0}\geq 1$ when $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$, since we will also need such a result in Subsection 4.4.
-From equation (4.2), along with Lemma 2.1(b) and (3.25) with $b=1$, we have
-(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.99(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
+From the equality in (4.2), along with Lemma 2.1(b) and (4.9), we have
+(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
 Applying (2.12) and (4.7) to (4.10), we obtain
-$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{3.99(1-c)}{0.89\left(21.47\sqrt{d}\,y_{k}/\left(|g|\mathcal{N}_{d^{\prime},4}\right)\right)^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
+$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{3.959(1-c)}{0.89\left(21.47\sqrt{d}\,y_{k}/\left(|g|\mathcal{N}_{d^{\prime},4}\right)\right)^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
 After taking the fourth power of both sides and rearranging, we find that
-(4.11) $\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}>y_{\ell}^{3}.$
+(4.11) $\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}>y_{\ell}^{3}.$
 Specialising to the case when $r_{0}=1$ and using $|g|\mathcal{N}_{d^{\prime},4}\geq 2$ from (3.36), we have
-(4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
+(4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
 We will now combine (4.12) with the gap principle in Lemma 3.8(a). to show that this case cannot occur.
-Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.7(a), we find that
-$57.6^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
+Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
+$57.32^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
 which we can rewrite as
@@
 We will see in Subsection 4.4 that $c=0.75$ is a good choice.
-So $y_{k}<\left|N_{\alpha}\right|$/4, which contradicts Lemma 3.5(c).
+So $y_{k}<\left|N_{\alpha}\right|/4$, which contradicts parts (b) and (c) of Lemma 3.5.
 Hence this case cannot hold.
@@
 $\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega^{1/4}-p_{r_{0}}\right|$
-$=\left|\frac{N_{d,4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d,4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}\right|$
-$=\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
-(4.13) $\geq\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
+$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
+$=$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
+(4.13) $\geq$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
 From Lemma 2.2(d), we have
@@
 Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}(\omega)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
-$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.291}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
+$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
 Applying (4.2) with this inequality, we obtain
-$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{1.16}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1},$
+$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{1.154}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1},$
 so
-(4.15) $1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>y_{\ell}.$
+(4.15) $1.734r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>y_{\ell}.$
 We now specialise to the case of $r_{0}=1$ and apply the gap principle in Lemma 3.8.
 Our gap principle in Lemma 3.8(a) with $b=1$, along with (4.15), implies that
-$6.92\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.6\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.6\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
-since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible.
+$6.96\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.32\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.32\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
+since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
 4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
@@
 Applying (4.5) and (4.8), and then (4.3), to this inequality, we obtain
-$\left(y_{k}y_{\ell}\right)^{1/4}$ $>\frac{0.995c\left|x_{k}\right|}{0.46\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}$
-$>\frac{2.1c\sqrt{d}\,y_{k}}{\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}.$
+$\left(y_{k}y_{\ell}\right)^{1/4}$ $>\frac{0.995c\left|x_{k}\right|}{0.458\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}$
+$>\frac{2.103c\sqrt{d}\,y_{k}}{\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}.$
 Taking the fourth power of both sides and rearranging, we find that this inequality implies
-(4.17) $y_{\ell}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}.$
+(4.17) $y_{\ell}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}.$
 With this lower bound for $y_{\ell}$, we now focus for the rest of this subsection on when $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$.
@@
 (4.18) $\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}$
-$>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{12}c^{12}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{12r_{0}}d^{6r_{0}}y_{k}^{12r_{0}-3}.$
+$>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{12}c^{12}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{12r_{0}}d^{6r_{0}}y_{k}^{12r_{0}-3}.$
 Using elementary calculus, $c^{12}(1-c)^{4}$ is monotonically increasing for $0<c\leq 0.75$.
@@
 Applying this to (4.18) and simplifying, we have
-(4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
+(4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
 Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
-$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
+$\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
 and then
-$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
+$\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
 Using the equality in (3.36) of Lemma 3.12, we can compute that $\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}/d^{4}\leq 2^{20}/17^{4}$ (the max value occurs when $d=17a^{2}$ so that $\left|N_{\alpha}\right|=16a^{2}$) to obtain
@@
 We now combine our upper bound for $y_{\ell}$ in (4.15) with our lower bound for $y_{\ell}$ in (4.17). Thus
-$1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}$
+$1.734r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}$
 and so
-$1.73r_{0}^{1/2}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}^{2}f_{\ell}^{2}}}\right)^{4}c^{4}\left(\frac{0.1832^{4}|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{4\left|N_{\alpha}\right|^{3}}\right)^{r_{0}}d^{r_{0}}y_{k}^{2r_{0}-2}.$
-We can show that $0.1832^{4r_{0}}/r_{0}^{1/2}>0.175^{4r_{0}}$, with the minimum being attained at $r_{0}=3$. Applying this, along with $c=0.75$, yields
-$1>\frac{0.739d}{f_{k}^{2}f_{\ell}^{2}\left|N_{\alpha}\right|}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
+$1.734r_{0}^{1/2}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832^{4}|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{4\left|N_{\alpha}\right|^{3}}\right)^{r_{0}}d^{r_{0}}y_{k}^{2r_{0}-2}.$
+We can show that $0.1832^{4r_{0}}/r_{0}^{1/2}>0.175^{4r_{0}}$, with the minimum being attained at $r_{0}=3$.
+Applying this, along with $c=0.75$ and collecting the terms taken to the power $r_{0}-1$, yields
+$1>\frac{0.7405d}{f_{k}^{2}f_{\ell}^{2}\left|N_{\alpha}\right|}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
 Combined with $d>N_{\alpha}$, this implies that
-(4.22) $1>\frac{0.739}{f_{k}^{2}f_{\ell}^{2}}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
+(4.22) $1>\frac{0.7405}{f_{k}^{2}f_{\ell}^{2}}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
 We now proceed similarly to the way we did in Subsection 4.4.
 Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b).
-Also, as $r_{0}-1\geq 1$, it follows that $0.739\cdot 0.0002344^{r_{0}-1}>\left(0.0001732\right)^{r_{0}-1}$, the above inequality implies
-$1>\left(\frac{0.0001732|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}>\left(y_{k}^{2}\frac{d\left(|g|\mathcal{N}_{d,4}\right)^{4}}{5774\left|N_{\alpha}\right|^{3}}\right)^{r_{0}-1}.$
+Also, as $r_{0}-1\geq 1$, it follows that $0.7405\cdot 0.0002344^{r_{0}-1}>0.0001735^{r_{0}-1}$.
+So the above inequality implies
+$1>\left(\frac{0.0001735|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}>\left(y_{k}^{2}\frac{d\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{5764\left|N_{\alpha}\right|^{3}}\right)^{r_{0}-1}.$
 But if
-(4.23) $y_{k}\geq\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d,4}\right)^{2}},$
+(4.23) $y_{k}\geq\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}},$
 then the right-hand side is greater than $1$, so this is not possible.
@@
 By Theorem 1.1 of [1], our theorem holds for $N_{\alpha}=-1$, so we may also assume that $N_{\alpha}\leq-4$.
-There are 74 such pairs $(a,d)$.
+There are 59 such pairs $(a,d)$.
 For each such pair, we solved $x^{2}-dy^{4}=N_{\alpha}$ using Magma (version V2.28-2) [4] and its IntegralQuarticPoints() function.
-For the 74 equations, this calculation took 17.66 seconds using MAGMA’s online calculator.
+For the 59 equations, this calculation took 17.66 seconds using MAGMA’s online calculator.
 21 of the equations had at least two solutions in positive integers.
@@
 None of the equations had more solutions.
-For $x^{2}-17y^{4}=-16$, note that $\left(103+5^{2}\sqrt{17}\right)/\left(1+\sqrt{17}\right)=\left(161+39\sqrt{17}\right)/8$, so the solutions $(103,5)$ and $(1,1)$ arise from different sequences.
+For $x^{2}-17y^{4}=-16$, since $\left(103+5^{2}\sqrt{17}\right)/\left(1+\sqrt{17}\right)=\left(161+39\sqrt{17}\right)/8$, the solutions $(103,5)$ and $(1,1)$ arise from different sequences.
 Hence Proposition 4.1 holds for $(a,d)=(1,17)$.
@@
 If $N_{\alpha}$ is even, then $|g|\mathcal{N}_{d^{\prime},4}\geq 4$ by Lemma 3.12.
-So the right-hand side of the inequality in Proposition 4.1 is at most $4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
-But for $u\geq 5$, we have $y_{k}>25\left|N_{\alpha}\right|/4$, by Lemma 3.5(c).
+So the right-hand side of the inequality in Proposition 4.1 is $\max\left(1,4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$.
+But for $u\geq 5$ and $k\neq 0$, we have $y_{k}\geq 25\left|N_{\alpha}\right|/4$ by Lemma 3.5(c).
 So Theorem 1.4 holds for $u\geq 5$ and we need only consider $1\leq u\leq 4$.
-Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}>81\left|N_{\alpha}\right|/4>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$ and we need only consider $1\leq u\leq 8$.
+Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}\geq 81\left|N_{\alpha}\right|/4>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$ for $k\neq 0$ and we need only consider $1\leq u\leq 8$.
+Next we treat the case when $K<-1$, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
+Lemma 5.1.
+If $K<-1$, then there are at most two distinct integer squares among the $y_{k}$’s.
+Proof.
+From Lemma 3.6, we know that $K<-1$ can only happen in the following two cases:
+(i) $a\geq 1$, $d=a^{2}+4$, $t=a$ and $u=1$, where $\alpha=2\varepsilon$ and $N_{\alpha}=-4$,
+(ii) $a\geq 1$, $d=a^{2}+1$, $t=2a$ and $u=2$, where $\alpha=\varepsilon$ and $N_{\alpha}=-1$.
+In case (i), we have $x_{k}^{2}-dy_{k}^{2}=-4$.
+If $d=a^{2}+4$ is even, then it is divisible by $4$ and $x_{k}$ is also even.
+So the equation becomes $\left(x_{k}/2\right)^{2}-(d/4)y_{k}^{2}=-1$.
+From Theorem 1.1 of [1], there are at most two distinct squares among the $y_{k}$’s here, so the lemma holds in this case.
+If $d$ is odd, the conditions in Theorem 1 of [9] hold with $A=d=a^{2}+4$, $B=1$ both odd and the minimal solution of equation (1) there in odd positive integers being $(1,a)$.
+Hence there are at most two distinct squares among the $y_{k}$’s in this case too.
+In case (ii), we can also apply Theorem 1.1 of [1]. ∎
+For the remainder of this section, we may assume that $K=-1$.
 Lemma 5.2.
@@
 Proof.
-Suppose that $y_{k}$ is a square for some $|k|\geq 2$.
-From Lemma 3.5(c), $\left|N_{\alpha}\right|=d-a^{2}<d$ and that by Theorem 1.1 of [1] we may assume $\left|N_{\alpha}\right|\geq 4$, we have $y_{k}>\left|N_{\alpha}\right|^{2}u^{4}/8$.
-So, for $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 8\left|N_{\alpha}\right|>8\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
-However, we saw above that for $N_{\alpha}$ even, if $y_{k}>4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$.
+From Theorem 1.1 of [1], we cannot have three distinct squares among the $y_{k}$’s for $N_{\alpha}=-1$, so we may assume $\left|N_{\alpha}\right|\geq 4$ for $N_{\alpha}$ even and $\left|N_{\alpha}\right|\geq 9$ for $N_{\alpha}$ odd.
+We will suppose that $y_{k}$ is a square with $|k|>1$, but show that this is not possible.
+From Lemma 3.5(c) and $\left|N_{\alpha}\right|=d-a^{2}<d$, we obtain $y_{k}\geq\left|N_{\alpha}\right|^{2}u^{4}/10$.
+So, for $u\geq 2$ and $\left|N_{\alpha}\right|\geq 4$, we have $y_{k}\geq(8/5)\left|N_{\alpha}\right|^{2}\geq(32/5)\left|N_{\alpha}\right|>(32/5)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
+However, we saw at the start of this section that for $N_{\alpha}$ even, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$.
 So we need only consider $u=1$ when $N_{\alpha}$ is even.
-Similarly, for $N_{\alpha}$ odd, if $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 18\left|N_{\alpha}\right|>18\left|N_{\alpha}\right|^{2}>90\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
-But, we saw above that for $N_{\alpha}$ even, if $y_{k}>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$.
+Similarly, if $-N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
+But, we saw at the start of this section that for $N_{\alpha}$ odd, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$.
 So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
-For $u=2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}$ for $|k|\geq 2$.
-So if $\left|N_{\alpha}\right|\geq 11$ is odd, then we have $y_{k}\geq 22\left|N_{\alpha}\right|>20\left|N_{\alpha}\right|$.
+For $u=2$, we have $y_{k}>(8/5)\left|N_{\alpha}\right|^{2}$ for $|k|\geq 2$.
+So if $\left|N_{\alpha}\right|\geq 25$ is odd, then we have $y_{k}\geq(200/5)\left|N_{\alpha}\right|>40\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
 So, for odd $N_{\alpha}$, we need only consider $N_{\alpha}=-9$.
-In this case, $N(\varepsilon)=\left(t^{2}-4d\right)/4=\pm 1$, so $t^{2}-4a^{2}=-4N_{\alpha}\pm 4=32,40$.
-This means that $\left(t,a,N_{\alpha}\right)=(\pm 6,1,-9)$.
-Here we have $y_{\pm 2}\geq 493>20\left|N_{\alpha}\right|$.
+In this case, $N_{\varepsilon}=\left(t^{2}-4d\right)/4=\pm 1$, so $t^{2}-4a^{2}=-4N_{\alpha}\pm 4=32,40$.
+This means that $\left(t,a,d\right)=(6,1,10)$ (recall from Subsection 1.2 that we only consider positive values of $t$ and $u$).
+Here we have $y_{\pm 2}\geq y_{-2}=493>19.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
 For $u=1$, we proceed similarly.
-Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/8$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $N_{\alpha}$ is even and $9\leq\left|N_{\alpha}\right|\leq 121$ when $-N_{\alpha}$ is an odd square.
+Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $-N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $-N_{\alpha}$ is an odd square.
 For each value of $N_{\alpha}$, the norm of $\varepsilon$ leads us to an equation of the form $t^{2}-a^{2}=-N_{\alpha}\pm 4$.
-We solve each of these equations over the integers using PARI/GP and calculate $y_{k}$ and $y_{-k}$ for $k\geq 2$ until the lower bound from Proposition 4.1 is exceeded (we never had to go beyond $k=7$).
+We solve each of these equations over the integers using PARI/GP and calculate $y_{k}$ and $y_{-k}$ for $k\geq 2$ until the lower bound from Proposition 4.1 is exceeded (we never had to go beyond $|k|=5$).
 No squares were found.
@@
 Lemma 5.3.
-Suppose that $N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
+If both $y_{\pm 1}$ are squares, then $y_{k}$ is not a square for any $|k|>1$.
+Proof.
+From Lemmas 3.8(a) and  3.5(c), we have
+$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
+By Proposition 4.1 and Lemma 3.12, if the quantity on the right-hand side exceeds $\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$, then the lemma follows.
+Since $d>\left|N_{\alpha}\right|$, this holds if $d^{2}u^{6}>19\cdot 64/57.32$, which holds unless $d^{2}<21.3$ and $u=1$.
+This leaves $d=2$ or $3$, but for neither of these does $u=1$ yield a unit.
+∎
+Lemma 5.4.
+Suppose that $-N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
 (a) If $t^{2}-du^{2}=4$, then $t\equiv 2\pmod{4}$ and $u\equiv 0\pmod{4}$.
@@
 If $a$ is even, then $d\equiv 1\pmod{8}$ and $u\equiv 8\pmod{16}$, which is not possible. ∎
-We now prove a lemma concerning $N_{\alpha}$ odd.
 Lemma 5.5.
-If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $\gcd\left(a^{2},d\right)=1$ and $u=2$, $t^{2}-du^{2}=-4$.
+If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
 Proof.
-(1) We start by considering $t^{2}-du^{2}=4$.
+We proceed in steps.
+(1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
 From Lemma 5.4(a), if $t^{2}-du^{2}=4$ and $-N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
@@
 So $t\equiv 0\pmod{3}$.
-Since $d\equiv 5\pmod{9}$), we have $y_{\pm 1}=1+32d\pm 4at\equiv 2\pm 4at\pmod{3}$.
+Since $d\equiv 5\pmod{9}$, we have $y_{\pm 1}=1+32d\pm 4at\equiv 2\pm 4at\pmod{3}$.
 Since $3|t$, it follows that $y_{\pm 1}\equiv 2\pmod{3}$, which can never be square.
@@
 In fact, using this same code, we could have eliminated case (1-i) above too.)
-Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 36$.
+Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 36$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>16\left|N_{\alpha}\right|$ here.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 16\left|N_{\alpha}\right|$ here.
 So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
@@
 (2-i) From Lemma 5.4(b), we have $u\equiv 2\pmod{4}$. For $u=6$, $t^{2}-du^{2}=-4$ is not possible modulo $9$. So we must have $u=2$ if $t^{2}-du^{2}=-4$.
-(2-ii) Finally, we show that $\gcd\left(a^{2},d\right)=1$ in this case.
+(2-ii) We show that $\gcd\left(a^{2},d\right)=1$ when $u=2$.
 We suppose otherwise.
@@
 Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$.
-Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 100$.
+Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 100$.
 Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|$ here.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$ here.
 So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$.
@@
 Lemma 5.6.
-If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
+If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.
 Proof.
 As in the proof of Lemma 5.5, we proceed in steps.
-(1) We start by considering $t^{2}-du^{2}=4$.
-(1-i) If $u=4$ and $t^{2}-du^{2}=4$, then we can argue as for $u=8$ in the proof of Lemma 5.3 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$.
-Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 144$.
+(1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
+(1-i) We show that $u=4$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
+Here we can argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$.
+Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 144$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq(76/144)\left|N_{\alpha}\right|$.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>4\left|N_{\alpha}\right|$ here.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 4\left|N_{\alpha}\right|$ here.
 So we can exclude $u=4$ and $t^{2}-du^{2}=4$ from consideration.
-(1-ii) Suppose that $u=3$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$.
+(1-ii) We show that $u=3$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 We will again argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $32$ to show that $64|\gcd\left(a^{2},d\right)$.
-Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 256$.
+Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 256$.
 Hence the theorem holds if $y_{\pm 1}>(76/256)\left|N_{\alpha}\right|$, by Proposition 4.1.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>(9/4)\left|N_{\alpha}\right|$ here.
-(1-iii) Suppose that $u=2$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq(9/4)\left|N_{\alpha}\right|$ here.
+(1-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 Using the argument for $u=8$ in the proof of Lemma 5.5 and the Maple program mentioned there modulo $16$, we obtain $16|\gcd\left(a^{2},d\right)$.
 Doing the same modulo $9$, we obtain $9|\gcd\left(a^{2},d\right)$.
-Combining this with $16|\gcd\left(a^{2},d\right)$ and applying Lemma 3.11, we have $|g|\mathcal{N}_{d,4}\geq 24$.
+Combining this with $16|\gcd\left(a^{2},d\right)$ and applying Lemma 3.12, we have $|g|\mathcal{N}_{d^{\prime},4}\geq 24$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/24^{2}\right)\left|N_{\alpha}\right|$.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|u^{2}/4=\left|N_{\alpha}\right|$ here.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|u^{2}/4=\left|N_{\alpha}\right|$ here.
 So we can exclude $u=2$ and $t^{2}-du^{2}=4$ from consideration.
-(1-iv) Suppose that $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$.
+(1-iv) We show that $u=1$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 Here too we use the argument and Maple program from the proof of Lemma 5.5, first modulo $9$ and then modulo $16$.
 From the latter, we obtain $64|\gcd\left(a^{2},d\right)$.
-So by Lemma 3.11, $|g|\mathcal{N}_{d,4}\geq 48$.
+So by Lemma 3.12, $|g|\mathcal{N}_{d^{\prime},4}\geq 48$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/48^{2}\right)\left|N_{\alpha}\right|$.
-But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|/4$ here.
+But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|/4$ here.
 So we can exclude $u=1$ and $t^{2}-du^{2}=4$ from consideration.
-(2) We now suppose that $t^{2}-du^{2}=-4$.
+(2) We now consider $t^{2}-du^{2}=-4$.
 Independent of the parity of $N_{\alpha}$, for $u=3$ and $4$, $t^{2}-du^{2}=-4$ is not possible modulo $9$ and $16$, respectively. So we can only have $u=1,2$ with $t^{2}-du^{2}=-4$
-(2-i) Consider $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$.
+(2-i) We show that if $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
 We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $-N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
-If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d,4}\geq 16$.
+If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d^{\prime},4}\geq 16$.
 So $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}\leq\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
@@
 Hence $b_{1}\equiv 2\pmod{4}$ and so $-N_{\alpha}\equiv 4\pmod{16}$.
-Now we show that $\gcd\left(a^{2},d\right)=1$ or $4$.
+(2-ii) We show that if $u=1$, $-N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
 We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $-N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 00 in the second case.
 We consider first $\gcd\left(a^{2},d\right)>4$ odd.
-So by Lemma 3.11 and the argument in the previous paragraph, $|g|\mathcal{N}_{d,4}=4\sqrt{\gcd\left(a^{2},d\right)}$.
+So by Lemma 3.12 and the argument in the previous paragraph, $|g|\mathcal{N}_{d^{\prime},4}=4\sqrt{\gcd\left(a^{2},d\right)}$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$.
@@
 Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$).
-So by Lemma 3.11 and the argument above, $|g|\mathcal{N}_{d,4}=2\sqrt{\gcd\left(a^{2},d\right)}$.
+So by Lemma 3.12 and the argument above, $|g|\mathcal{N}_{d^{\prime},4}=2\sqrt{\gcd\left(a^{2},d\right)}$.
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
@@
 If $\gcd\left(a^{2},d\right)\geq 100$, then $\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}\leq 0.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/4$, the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
-(2-ii) Lastly, we consider $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$.
+(2-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
 Arguing modulo $64$, we have $d\equiv a^{2}\equiv 1\pmod{16}$ and $8|t$.
@@
 ∎
-Theorem 1.4 now follows from the bounds on $u$ from Proposition 4.1, along with Lemmas 5.1, lem:nrmA-odd and lem:nrmA-even.
+Theorem 1.4(c) now follows from the bounds on $u$ that we obtained from Proposition 4.1 at the start of Section 5, along with Lemmas 5.1, 5.2, 5.5 and 5.6.
+Furthermore, from Lemma 5.3, if $y_{-1}$ and $y_{1}$ are both squares, then there are no further squares.
+So we may assume that precisely one of $y_{\pm 1}$ is a square.
+From Lemma 3.5(c), we obtain $y_{k}>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$ for $|k|\geq 2$, provided that $d>190$ when $u=1$ and that $d>11$ when $u=2$.
+This leaves $26$ values of $d$.
+For each of these, we compute $y_{-2}$ directly for each possibility of $N_{\alpha}$ and compare the value to the bound in Proposition 4.1.
+Where the bound was exceeded, Proposition 4.1 tells us that there are no further squares, completing the proof for them.
+There were only four cases where $y_{-2}$ did not exceed that bound: $(a,b,d,t,u)=(1,1,2,2,2)$, $(1,1,5,1,1)$, $(2,1,8,2,1)$, $(3,1,13,3,1)$.
+These were treated in Subsection 4.6.
 6. Examples
@@
 6.1. Examples for Conjecture 1.1
-For $d=2$ with $t=u=2$.
 $a$ $b$ indices, $k$ $\sqrt{y_{k}}$
-1 $3^{2}$ [0, -1, 3, -5] [3, 5, 31, 167]
-1019 $27^{2}$ [0, 1, -3, -7] [27, 65, 29, 983]
-167 $13^{2}$ [0, 1, -3, 4] [13,29,71, 407]
-157 $29^{2}$ [0, -1, 3, -4] [29, 47, 307, 649]
-1 $41^{2}$ [0, -1, -9, 11] [41, 71, 80753, 470861]
-1633 $65^{2}$ [0, -1, -4, 7] [65, 97, 1331, 24791]
-48479 $211^{2}$ [0, -3, 4, -7] [211, 1007, 6743, 34205]
-45649 $677^{2}$ [0, -1, -4, 4] [677, 1133, 15679, 16825]
-1940147 $1217^{2}$ [0, -3, 4, -11] [1217, 3289, 40573, 3794239]
-600589 $2213^{2}$ [0, -1, -4, 4] [2213, 3673, 50801, 55415]
-20509501 $8689^{2}$ [0, -1, 3, -4] [8689, 13619, 94393, 187603]
-255488029 $13457^{2}$ [0, -1, 3, -4] [13457, 5683, 189241, 15821]
-409660129 $17023^{2}$ [0, -1, -4, -8] [17023, 7073, 7949, 269495]
-3032771269 $46313^{2}$ [0, -1, -4, -8] [46313, 19213, 15269, 516625]
-We also found such examples with $(d,t,u,a,b)=(3,4,2,672,8281)$, $(6,10,4,78,49)$, $(6,10,4,34986,22201)$, $(6,10,4,3663828,5094049)$, $(30,22,4,826320,1234321)$ and $(37,12,2,138,25)$.
+1 $3$ [0, -1, 3, -5] [3, 5, 31, 167]
+1019 $27$ [0, 1, -3, -7] [27, 65, 29, 983]
+167 $13$ [0, 1, -3, 4] [13,29,71, 407]
+157 $29$ [0, -1, 3, -4] [29, 47, 307, 649]
+1 $41$ [0, -1, -9, 11] [41, 71, 80753, 470861]
+1633 $65$ [0, -1, -4, 7] [65, 97, 1331, 24791]
+48479 $211$ [0, -3, 4, -7] [211, 1007, 6743, 34205]
+45649 $677$ [0, -1, -4, 4] [677, 1133, 15679, 16825]
+1940147 $1217$ [0, -3, 4, -11] [1217, 3289, 40573, 3794239]
+600589 $2213$ [0, -1, -4, 4] [2213, 3673, 50801, 55415]
+20509501 $8689$ [0, -1, 3, -4] [8689, 13619, 94393, 187603]
+255488029 $13457$ [0, -1, 3, -4] [13457, 5683, 189241, 15821]
+409660129 $17023$ [0, -1, -4, -8] [17023, 7073, 7949, 269495]
+3032771269 $46313$ [0, -1, -4, -8] [46313, 19213, 15269, 516625]
+Table 1. Examples for $d=2$ with $t=u=2$
+In addition to the examples for $d=2$ in Table 1, we also found examples with four squares for $(d,t,u,a,b)=(3,4,2,672,91)$, $(6,10,4,78,7)$, $(6,10,4,34986,149)$,
+$(6,10,4,3663828,2257)$, $(30,22,4,826320,1111)$ and $(37,12,2,138,5)$.
 At least for $d=2$ and $d=6$, it appears there may be infinitely many such examples.
@@
 The examples in Subsection 6.1 also apply in this more general case to show there can be at least four distinct squares.
+Here are two examples showing that there are at least three distinct squares when $\left|N_{\alpha}\right|$ is a square:
+$(d,t,u,a,b)=(5,1,1,43,3)$: $N_{\alpha}=38^{2}$, $y_{-3}=5^{2}$, $y_{0}=3^{2}$, $y_{11}=53^{2}$;
+$(d,t,u,a,b)=(10,6,2,1,1)$: $N_{\alpha}=-3^{2}$, $y_{0}=1^{2}$, $y_{1}=2^{2}$, $y_{2}=5^{2}$.
+Here are two examples showing that the same is true when $\left|N_{\alpha}\right|$ is a prime power:
+$(d,t,u,a,b)=(5,1,1,153,4)$: $N_{\alpha}=22129$, $y_{-3}=11^{2}$, $y_{0}=4^{2}$, $y_{30}=8862^{2}$;
+$(d,t,u,a,b)=(51,100,14,2,1)$: $N_{\alpha}=-47$, $y_{-1}=6^{2}$, $y_{0}=1^{2}$, $y_{1}=8^{2}$.
+Here is an example showing that if $\left|N_{\alpha}\right|$ is not a prime power, but $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$ is a prime, then there can be distinct four squares (unlike in Conjecture 1.1):
+$(d,t,u,a,b)=(5,1,1,7,1)$: $N_{\alpha}=2^{2}\cdot 11$, $y_{-9}=9^{2}$, $y_{0}=1$, $y_{1}=2^{2}$, $y_{3}=3^{2}$.
+And finally, an example showing that if $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$ is twice an odd prime, then there can be distinct four squares (again unlike in Conjecture 1.1):
+$(d,t,u,a,b)=(6,10,4,2,3)$: $N_{\alpha}=-2\cdot 241$, $y_{-3}=63^{2}$, $y_{0}=3^{2}$, $y_{1}=7^{2}$, $y_{3}=69^{2}$.
+6.4. Examples for Theorem 1.4(a)
+From (3.14) with $b=u=1$, we find that $(2t+a)^{2}-8y_{1}=a^{2}-8$.
+Since $-N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $22t+$ 2 is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$.
+The corresponding values of $\sqrt{y_{1}}$ are the elements of the sequence $\left(u_{n}\right)_{n\geq 0}$ with $u_{0}=5$, $u_{1}=169$ and $u_{n+1}=34u_{n}-u_{n-1}$.
+Then $y_{1}$ is approximately $t^{2}/2\approx\left|N_{\alpha}\right|/2\approx\left|N_{\alpha}\right|^{3/2}/\left(2\sqrt{d}\right)$.
+In fact, such $y_{1}$’s are the ones that are squares satisfying the conditions in Theorem 1.4(a) with the smallest ratio compared to $\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, the quantity in Proposition 4.1.
 6.5. Examples for Lemma 3.5
@@
 There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $-N_{\alpha}$ is a square too.
+The sequences given in Lemma 3.6 show that for $k=-1$, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.

### v3

@@
 Observe that $x_{0}=a$, $y_{0}=b^{2}$,
-[…]
+(1.2) $y_{1}=\frac{b^{2}\left(t^{2}+du^{2}\right)+2atu}{4},\hskip 8.53581pty_{-1}=\frac{b^{2}\left(t^{2}+du^{2}\right)-2atu}{4}$
 and that both sequences satisfy the recurrence relation
@@
 Suppose that $b=1$ and $N_{\alpha}<0$. There are at most three distinct integer squares among the $y_{k}$’s.
-If $-N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
+If $N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
 In fact, a more general result than Conjecture 1.1 also appears to be true.
@@
 1.4. Results
-In this paper, we prove Conjecture 1.2 when $-N_{\alpha}$ is a square, except for a very limited set of sequences, where we prove a slightly weaker result.
+In this paper, we prove Conjecture 1.2 when $N_{\alpha}$ is a square, except for a very limited set of sequences, where we prove a slightly weaker result.
 Theorem 1.4.
-Let $b=1$, $a$ and $d$ be positive integers, where $d$ is not a square, $N_{\alpha}<0$ and $-N_{\alpha}$ is a square.
+Let $b=1$, $a$ and $d$ be positive integers, where $d$ is not a square, $N_{\alpha}<0$ and $N_{\alpha}$ is a square.
 (a) If $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$, $\gcd\left(a^{2},d\right)=1,4$ and one of $y_{\pm 1}$ is a perfect square, then there are at most three distinct squares among the $y_{k}$’s.
@@
 Suppose that there exist $k_{0},\ell_{0}>0$ and $E,Q>1$ such that for all non-negative integers $r$, there are algebraic integers $p_{r}$ and $q_{r}$ in $\mathbb{K}$ with $\left|q_{r}\right|<k_{0}Q^{r}$ and $\left|q_{r}\theta-p_{r}\right|\leq\ell_{0}E^{-r}$ satisfying $p_{r}q_{r+1}\neq p_{r+1}q_{r}$.
-For any algebraic integers $p$ and $q$ in $\mathbb{K}$, let $r_{0}$ be the smallest positive integer such that $\left(Q-1/E\right)\ell_{0}|q|/\left(Q-1\right)<cE^{r_{0}}$, where $10<c<$ 1.
+For any algebraic integers $p$ and $q$ in $\mathbb{K}$, let $r_{0}$ be the smallest positive integer such that $\left(Q-1/E\right)\ell_{0}|q|/\left(Q-1\right)<cE^{r_{0}}$, where $0<c<1$.
 (a) We have
@@
 Let $t^{\prime}$, $u_{1}$ and $u_{2}$ be rational integers with $t^{\prime}<0$  [^1] To avoid confusion with $t$ in the expression for $\varepsilon$, we use $t^{\prime}$ here to denote what was $t$ in our previous works like [16].
-Similarly, we will use $d^{\prime}$ where $d$ was used in previous works..
-We let $u=\left(u_{1}+u_{2}\sqrt{t^{\prime}}\right)/2$ be an algebraic integer in $\mathbb{K}=\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$ with $\sigma(u)=\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/2$ as its algebraic (and complex) conjugate.
-Put $\omega=u/\sigma(u)$ and write $\omega=e^{i\varphi}$, where $-\pi<\varphi\leq\pi$.
+Similarly, we will use $d^{\prime}$ where $d$ was used in previous works.
+such that $u=\left(u_{1}+u_{2}\sqrt{t^{\prime}}\right)/2$ be an algebraic integer in $\mathbb{K}=\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$ with $\sigma(u)=\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/2$ as its algebraic (and complex) conjugate.
+Put $\omega=u/\sigma(u)$ and write $\omega=e^{i\varphi}$, where $\pi<\varphi\leq\pi$.
 For any real number $\nu$, we shall put $\omega^{\nu}=e^{i\nu\varphi}$ – unless otherwise stated, we will use this convention throughout this paper.
@@
 For integers $m$ and $n$ with $0<m<n$, $(m,n)=1$ and $r$ a non-negative integer, put $\nu=m/n$ and
-$X_{m,n,r}(z)={}_{2}F_{1}(-r-\nu,-r,1-\nu,z),\quad Y_{m,n,r}=z^{r}X_{m,n,r}\left(z^{-1}\right)$
+$X_{m,n,r}(z)={}_{2}F_{1}(-r-\nu,-r,1-\nu,z),\quad Y_{m,n,r}(z)=z^{r}X_{m,n,r}\left(z^{-1}\right)$
 and
@@
 $g_{2}$ $=\gcd\left(u_{1}/g_{1},t^{\prime}\right),$
-(2.8) $g_{3}$ =$=\left\{\begin{array}[]{ll}1&\text{if $t^{\prime}\equiv 1\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 2&\text{if $t^{\prime}\equiv 3\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 4&\text{otherwise,}\end{array}\right.$
+(2.8) $g_{3}$ $=\left\{\begin{array}[]{ll}1&\text{if $t^{\prime}\equiv 1\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 2&\text{if $t^{\prime}\equiv 3\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 4&\text{otherwise,}\end{array}\right$.
 $g$ $=g_{1}\sqrt{g_{2}/g_{3}}.$
@@
 $=\frac{D_{4,r}}{N_{d^{\prime},4,r}}\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r}X_{1,4,r}\left(1-u_{2}\sqrt{t^{\prime}}\frac{-2}{u_{1}-u_{2}\sqrt{t^{\prime}}}\right).$
-Since $d^{\prime}=u_{2}^{2}t^{\prime}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$.
+Since $d^{\prime}=u_{2}^{2}t^{\prime}/g^{2}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$.
 The analogous expression for $q_{r}$ shows that it is also an algebraic integer.
@@
 and
-$\ell_{0}=\mathcal{C}_{4,2}|\varphi|=0.2|\varphi|.$
+(2.14) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|=0.2|\varphi|.$
 3. Lemmas about $\left(x_{k}\right)_{k=-\infty}^{\infty}$ and $\left(y_{k}\right)_{k=-\infty}^{\infty}$
@@
 Proposition 3.1.
-Let $a\neq 0$,  [^2] PV: why do I want $a\neq 0$?
-$a=0$ will give me the squares among the $x_{k}$’s from $\varepsilon^{2k}$.
-$b>0$ and $d$ be rational integers such that $d$ is not a square.
+Let $a\neq 0$, $b>0$ and $d$ be rational integers such that $d$ is not a square.
 Put $\alpha=a+b^{2}\sqrt{d}$ and denote $N_{\mathbb{Q}(\sqrt{d})/\mathbb{Q}}(\alpha)$ by $N_{\alpha}$.
@@
 where $f^{\prime}|\operatorname{sf}\left(N_{\alpha}\right)$ and $0<f^{\prime}<\max\left(2,\sqrt{\left|\operatorname{sf}\left(N_{\alpha}\right)\right|}\right)$.
-(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
-±$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
+(b) If $N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
+$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
 (c) If $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)=2^{\ell}p^{m}$ where $\ell,m\in\{0,1\}$ with $\ell+m\geq 1$ and $p$ is an odd prime, then we have $f\mid\left(4b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ when $N_{\alpha}\equiv 1\pmod{4}$ and $4|d$, and $f\mid\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right),N_{\varepsilon}\right)\right)\right)$ otherwise.
@@
 With $g_{1}^{2}=\gcd\left(4b^{2}r_{1}/\operatorname{sf}\left(r_{1}\right),r_{1}^{2}\right)$, we have
-(3.3) $v_{p}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2},&\text{if $k_{2}$ is even and $2k_{1}\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}\geq k_{2}$,}\\ 2k_{1}+k_{2}-1,&\text{if $k_{2}$ is odd and $2k_{1}-1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
+[…]
 From the definitions of $r_{1}$ and $k_{2}$, it follows that $v_{p}(au)\geq\min\left(k_{2},v_{p}\left(tb^{2}\pm 2by\right)\right)$.
@@
 The proof when $p=2$ is similar. But we will use the following information later in the proof of Proposition 3.1. We have
-(3.5) $v_{2}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2}+2,&\text{if$k_{2}$ is even and $2k_{1}+2\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}+2\geq k_{2}$,}\\ 2k_{1}+k_{2}+1,&\text{if $k_{2}$ is odd and $2k_{1}+1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
+[…]
 As above, we have $v_{2}(au)\geq\min\left(k_{2},v_{2}\left(tb^{2}\pm 2by\right)\right)$.
 Applying this and the right-hand expression in (3.4) for $4y^{2}$, we obtain
-$2\displaystyle 2v_{2}(y)+$ 2 $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
+$2v_{2}(y)+2$ $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
 $\geq\min\left(v_{2}\left(2t^{2}b^{2}\pm 4bty\right),k_{2}+v_{p}(t)+1,v_{2}\left(2b^{2}\right)\right)$
@@
 As in the statement of Lemma 3.3, we let $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$ be the two possibilities.
+We let $g_{1}^{\prime}$ and $g_{1}^{\prime\prime}$ be the associated values of $g_{1}$.
 We will use Lemma 3.3(b).
 We divide the prime divisors of $\operatorname{sf}\left(N_{\alpha}\right)$ into three disjoint sets:
-$I^{\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
-$I^{\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
-$I^{\prime\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
+$I^{\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
+$I^{\prime\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
+$I^{\prime\prime\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
 With $d^{\prime}=\prod_{p\in I^{\prime}}p$, $d^{\prime\prime}=\prod_{p\in I^{\prime\prime}}p$ and $d^{\prime\prime\prime}=\prod_{p\in I^{\prime\prime\prime}}p$, we have $d^{\prime}d^{\prime\prime}d^{\prime\prime\prime}=\operatorname{sf}\left(N_{\alpha}\right)$.
@@
 Here we will replace $r$ and $s$.
-We can write $(r+si)/(1+i)=(r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
-$-\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
+We can write (r+si)/(1+i)=(r+s)/2+(sr)i/2(r+si)/(1+i)=(r+s)/2+(s-r)i/2. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
+$\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
 Since $4b^{2}r/g_{1}^{2}$ is even, we know that $2b^{2}r/g_{1}^{2}\in\mathbb{Z}$.
-Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$.
+Since $N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$.
 Hence $2b^{2}r/g_{1}^{2}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ and so we take $f=2b^{2}r/g_{1}^{2}$ in this case.
@@
 A consequence of these two inequalities is that $64|\left(u^{2}N_{\alpha}\right)$.
-From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., $atu$ and $b^{2}du^{2}$ have the same parity.
+From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., atuatu and $b^{2}du^{2}$ have the same parity.
 (b-iv-1) If $b^{2}du^{2}$ is odd, then $a$ and $t$ are also odd. Since $4N_{\varepsilon}=t^{2}-du^{2}$ is divisible by $4$ and $t$ and $u$ are both odd, it must be the case that $d\equiv 1\pmod{4}$.
 If $d\equiv 1\pmod{8}$, then $4N_{\varepsilon}=t^{2}-du^{2}\equiv 0\pmod{8}$ and so $N_{\varepsilon}$ is even.
-Furthermore, since $abd$ is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
+Furthermore, since abdabd is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
 If $d\equiv 5\pmod{8}$, then $N_{\alpha}=a^{2}-db^{4}\equiv 4\pmod{8}$. Since $u$ is odd, we get a contradiction with $64|\left(u^{2}N_{\alpha}\right)$.
@@
 Lemma 3.5.
-[…]
+Let the $y_{k}$’s be defined by (1.1) with the notation and assumptions there. Suppose that $N_{\alpha}<0$. Let $K$ be the largest negative integer such that $y_{K}>b^{2}$.
 (a) Put $\overline{\alpha}=a-b^{2}\sqrt{d}$. We have
@@
 (3.13) $y_{k}\geq\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{\max(0,K-k)}&\text{for $k<0$.}\end{array}\right$.
-In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $52du^{2}/$ 5 by $85du^{2}/$ 8 and if $N_{\varepsilon}=1$, then we can replace $52du^{2}/$ 5 by $du^{2}$.
+In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $2du^{2}/5$ by $5du^{2}/8$ and if $N_{\varepsilon}=1$, then we can replace $2du^{2}/5$ by $du^{2}$.
 Remark.
@@
 (3.14) $4b^{2}y_{1}=b^{4}\left(t^{2}+du^{2}\right)+2ab^{2}tu=\left(b^{2}t+au\right)^{2}-N_{\alpha}u^{2}.$
-From the second equality in (3.14), equation (3.13) follows for $k=1$.
-Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence.
+From the second equality in (3.14) and $N_{\alpha}<0$, equation (3.13) follows for $k=1$.
+Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u/2$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence.
 Hence, for $k<0$,
@@
 Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain equation (3.13) for $k\leq K$ in the same way. ∎
-We will also need to know when $K$ in Lemma 3.5 is not equal to $-1$ for $b=1$.
+We will also need to know when $K$ in Lemma 3.5 is not equal to 1-1 for $b=1$.
 Lemma 3.6.
-[…]
+Let the $y_{k}$’s be defined by (1.1) with the notation and assumptions there. Suppose that $b=1$ and $N_{\alpha}<0$.
 We have $y_{-1}>1$, except for
@@
 (ii) $a\geq 1$, $d=a^{2}+1$, $t=2a$ and $u=2$, where $\alpha=\varepsilon$ and $N_{\alpha}=-1$.
-In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, $K=-2$.
+In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, K=2K=-2.
 Proof.
@@
 $a^{2}-t^{2}=-8$ can only occur for $a=1$ and $t=3$ (since we are assuming that $a$ and $t$ are positive integers).
-But in this case, $b^{2}t-au=2$, so $84y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=$ 8.
+But in this case, $b^{2}t-au=2$, so $4y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=8$.
 Otherwise, we have $t=a$. So $d=a^{2}+4$, $\alpha=2\varepsilon$ and $N_{\varepsilon}=-1$. Here $b^{2}t-au=0$, so $y_{-1}=1$.
 In case (2), we have $d=\left(t^{2}\pm 4\right)/4$ and $a^{2}-d=-1$.
-We proceed similarly, obtaining $84a^{2}-t^{2}=-$ 8 (which is never possible) when $N_{\varepsilon}=1$ and $04a^{2}-t^{2}=$ 0 when $N_{\varepsilon}=-1$.
+We proceed similarly, obtaining $4a^{2}-t^{2}=-8$ (which is never possible) when $N_{\varepsilon}=1$ and $4a^{2}-t^{2}=0$ when $N_{\varepsilon}=-1$.
 Here we have $t=2a$ and $d=a^{2}+1$, so $\alpha=\varepsilon$.
@@
 Lemma 3.7.
-Let $\omega=e^{i\theta}$ with $-\pi<\theta\leq\pi$.
+Let $\omega=e^{i\theta}$ with $\pi<\theta\leq\pi$.
 (a) Put $\omega^{1/4}=e^{i\theta/4}$. If
@@
 $\prod_{k=1}^{3}\left|e^{2\pi ik/4-i\theta/4}z-1\right|=\left|e^{3i\varphi}+e^{2i\varphi}+e^{i\varphi}+1\right|,$
-for $-\pi<\varphi\leq\pi$ such that $e^{i\varphi}=e^{-i\theta/4}z$. Squaring this quantity and simplifying, we obtain
+for $\pi<\varphi\leq\pi$ such that $e^{i\varphi}=e^{-i\theta/4}z$. Squaring this quantity and simplifying, we obtain
 (3.15) $8\cos^{2}(\varphi)\left(\cos(\varphi)+1\right).$
@@
 $=\frac{2-2\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
-Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$.
+Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $16\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=0$ and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$.
 Part (b) now follows and follows with inequalities too, since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$.
@@
 Lemma 3.8.
-[…]
-(a) Suppose that $-N_{\alpha}$ is a square.
-If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right\}$, then
+Let the $y_{k}$’s be defined as in (1.1) with $N_{\alpha}<0$.
+(a) Suppose that $N_{\alpha}$ is a square.
+If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right)$, then
 $y_{j}>57.32\left(\frac{d}{b^{2}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
-(b) Suppose $-N_{\alpha}$ is not a square.
+(b) Suppose $N_{\alpha}$ is not a square.
 If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $k_{1},k_{2},k_{3}\neq 0$ and $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
@@
 where $f_{i}$ and $f_{j}$ are the values of $f$ in Proposition 3.1 associated with $y_{i}$ and $y_{j}$, respectively.
-If $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right\}$, then we can replace $15.36$ with $182$.
+If $y_{j}>y_{i}\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right)$, then we can replace $15.36$ with $182$.
 Proof.
@@
 Since we have assumed that $i\neq 0$ and $j\neq 0$, we can apply Proposition 3.1 to show that there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
-±$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
+$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
 (3.16) $f_{j}\sqrt{y_{j}}$ $=b\left(r_{j}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{j}^{2}\right),$
 where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$.
-Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon}=\pm 1$.
-For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always $+$. That is,
-(3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
-$f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
+Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon^{j}}=\pm 1$.
+For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always ++. That is,
+(3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon^{i}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
+$f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
 as the argument for the other cases is exactly the same. Subtracting the complex conjugate of one of these equations from the equation itself, we obtain
 (3.18) $\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
-$=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
+$=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
 and the analogous equation with the index $j$ replaced by $i$ also holds.
 Putting $\omega=\left(a-\sqrt{N_{\alpha}}\right)/\left(a+\sqrt{N_{\alpha}}\right)$, by (3.18) and the relationship for $f_{j}\sqrt{y_{j}}$ in Proposition 3.1(a), we have
-(3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ =|±$=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
+(3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ $=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
 $=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}\leq\frac{1}{2},$
@@
 From (3.17), it follows that
-$\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{j}=\pm\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{i}.$
-Comparing the imaginary parts of both sides of this equation, we find that $y_{i}=y_{j}$, but this contradicts our assumption that $y_{j}>y_{i}$. Hence the left-hand side of (3.22) cannot be 00.
+$\left(x_{i}+N_{\varepsilon^{i}}\sqrt{N_{\alpha}}\right)y_{j}=\pm\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)y_{i}.$
+Comparing the imaginary parts of both sides of this equation, we find that $y_{i}=y_{j}$, but this contradicts our assumption that $y_{j}>y_{i}$. Hence the left-hand side of (3.22) cannot be 0.
 Let $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$.
@@
 At this point, our proofs of the two parts of the lemma separate.
-(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
+(a) Suppose that $N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
 $\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
 using (3.25) above.
-Furthermore, from Proposition 3.1(b), we know that $f_{i},f_{j}|b^{2}$, since $-N_{\alpha}$ is a square. So
+Furthermore, from Proposition 3.1(b), we know that $f_{i},f_{j}|b^{2}$, since $N_{\alpha}$ is a square. So
 $\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\geq\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}.$
@@
 (b) We again use (3.23) and (3.24).
-Notice that $x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}$ on the left-hand side of (3.24) for $\zeta_{4}^{(j)}=\pm\sqrt{-1}\zeta_{4}^{(i)}$ can be as small as $1/\left(2\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}y\right)$ when $-N_{\alpha}$ is not a square, so the proof of part (a) breaks down here.
-Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other.
+Notice that $x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}$ on the left-hand side of (3.24) for $\zeta_{4}^{(j)}=\pm\sqrt{-1}\zeta_{4}^{(i)}$ can be as small as $1/\left(2\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}y\right)$ when $N_{\alpha}$ is not a square, so the proof of part (a) breaks down here.
+Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is $\pm 1$ times the other.
 Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $y_{i}<y_{j}$.
@@
 Lemma 3.9.
-[…]
+Let the $y_{k}$’s be defined as in (1.1) with the notation and assumptions there. Suppose that $N_{\alpha}<0$.
 (a) Let $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ be a square and put
-$\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)=e^{i\varphi}$
-with $-\pi<\varphi\leq\pi$. Then
-$\left|\varphi\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
-(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega$ be as in part (a) and put
+$\omega_{k}=\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)=e^{i\varphi_{k}}$
+with $\pi<\varphi_{k}\leq\pi$. Then
+$\left|\varphi_{k}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}<0.6.$
+(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $k,\ell\neq 0$, $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega_{k}$ be as in part (a) and put
 $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$
-with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1.
+with $_{k}, rr_{\ell}$, $s_{k}$ and $s_{\ell}$ as in Proposition 3.1.
 Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.8 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$.
 Then
-$\min_{0\leq j\leq 3}\left|\omega^{1/4}-\zeta_{4}^{j}\frac{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
-occurs for either $j=0$ or $j=2$.
+$\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-\zeta_{4}^{j}\frac{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
+occurs for either $j=0$ or $j=2$, where $\zeta_{4}$ is a primitive $4$-th root of unity.
 Proof.
-(a) We can write $\omega=\left(\left|x_{k}\right|\pm\sqrt{N_{\alpha}}\right)^{2}/\left(x_{k}^{2}+\left|N_{\alpha}\right|\right)$, so with $\omega=e^{i\varphi}$, we have $\left|\tan\left(\varphi\right)\right|=\left|2x_{k}\sqrt{\left|N_{\alpha}\right|}/\left(x_{k}^{2}+N_{\alpha}\right)\right|$.
+(a) We can write $\omega_{k}=\left(\left|x_{k}\right|\pm\sqrt{N_{\alpha}}\right)^{2}/\left(x_{k}^{2}+\left|N_{\alpha}\right|\right)$, so with $\omega_{k}=e^{i\varphi_{k}}$, we have $\left|\tan\left(\varphi_{k}\right)\right|=\left|2x_{k}\sqrt{\left|N_{\alpha}\right|}/\left(x_{k}^{2}+N_{\alpha}\right)\right|$.
 From (1.1) and $N_{\alpha}>-b^{4}d$, we have
@@
 since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ and $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}<0$.
-From $|\varphi|\leq|\tan(\varphi)|$, we have
-$|\varphi|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
+From $\left|\varphi_{k}\right|\leq\left|\tan\left(\varphi_{k}\right)\right|$, we have
+$\left|\varphi_{k}\right|\leq\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{x_{k}^{2}+N_{\alpha}}\right|<\left|\frac{2\sqrt{\left|N_{\alpha}\right|}x_{k}}{0.875x_{k}^{2}}\right|<\frac{2.29\sqrt{\left|N_{\alpha}\right|}}{\left|x_{k}\right|}.$
 Since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have $dy_{k}^{2}\geq 16\left|N_{\alpha}\right|$ and so
 $x_{k}^{2}=dy_{k}^{2}+N_{\alpha}\geq 15\left|N_{\alpha}\right|.$
-Combining this with the inequality above it yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
+Combining this with the inequality above it yields $\left|\varphi_{k}\right|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
 (b) Recall that from (3.18) in the proof of Lemma 3.8, we can write
 $\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
-=±$=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
+$=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
 for $i=k,\ell$. There are other cases according to the signs in the result in Proposition 3.1, but the argument for them is identical to the argument that follows below.
-Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
-$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
-$-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+Using this relationship for $i=\ell$ and using our expressions in Proposition 3.1 for $x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}$ and $f_{k}\sqrt{y_{k}}$, we have
+$\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
+$-\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
 $=$ $\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}}{f_{k}^{2}}\left[\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right$.
@@
 Applying equation (3.18) with $j=\ell$ to the last expression and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
-(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
-$\left.-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
+(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right$.
+$\left.-\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
 $=$ $\frac{2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}}\sqrt{\left|N_{\alpha}\right|}.$
 Let $\zeta_{4}$ be the $4$-th root of unity satisfying
-$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
-From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (1.1) (which implies that $\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
-(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
-$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
+$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
+From (3.29), our expression in the statement of this lemma for $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$, the expressions for $y_{k}$ and $y_{\ell}$ in Proposition 3.1, and (1.1) (which implies that $\left|x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right|^{2}=x_{k}^{2}-N_{\alpha}=dy_{k}^{2}$), we have
+(3.30) $\left|\omega_{k}-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
+$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
 $=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}\leq\frac{1}{2}.$
@@
 By Lemma 3.7(b) with $c_{0}=1/2+0.0001$,
-(3.31) $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
+(3.31) $\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.1263.$
 From (3.22), along with $y_{k},y_{\ell}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have
@@
 $<0.5051b^{2}\frac{\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}}\left(\frac{1}{y_{k}}+\frac{1}{y_{\ell}}\right)<0.253.$
-From part (a), we have $|\varphi|<0.6$, so $\left|\omega^{1/4}-1\right|<0.15$ and
-$\left|\omega^{1/4}-\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\leq\left|\omega^{1/4}-1\right|+\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|<0.403.$
+From part (a), we have $\left|\varphi_{k}\right|<0.6$, so $\left|\omega_{k}^{1/4}-1\right|<0.15$ and
+$\left|\omega_{k}^{1/4}-\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|\leq\left|\omega_{k}^{1/4}-1\right|+\left|\frac{\zeta_{4}^{(k)}}{\zeta_{4}^{(\ell)}}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-1\right|<0.403.$
 Recalling (3.31), it follows that
-$\left|\omega^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
-$=$ $\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
-$\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
+$\left|\omega_{k}^{1/4}-\zeta_{4}^{\prime}\frac{\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}{\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)}\right|$
+$=$ $\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}+\left(\zeta_{4}-\zeta_{4}^{\prime}\right)\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
+$\geq\left|\zeta_{4}-\zeta_{4}^{\prime}\right|-\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\sqrt{2}-0.127,$
 for any $4$-th root of unity, $\zeta_{4}^{\prime}$, with $\zeta_{4}^{\prime}\neq\zeta_{4}$.
@@
 Let $\left(x_{k}\right)_{k=0}^{\infty}$ and $\left(y_{k}\right)_{k=0}^{\infty}$ be sequences defined by $x_{k}+y_{k}\sqrt{d}=\left(a+b\sqrt{d}\right)\varepsilon^{k}$, where $a,b,d$ are positive integers, $d$ is not a square and $\varepsilon=\left(t+u\sqrt{d}\right)/2\neq\pm 1$, with $t$ and $u$ integers, is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$.
-If $x_{k}$ and $y_{k}$ are both integers, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
+If $x_{k}$ and $y_{k}$ are both integers and $\gcd(a,b)$ is odd, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1$ or $2$.
+The additional hypothesis that $\gcd(a,b)$ is odd was omitted from the published version of this lemma. If $\gcd(a,b)$ is even, then $\gcd\left(x_{k},y_{k}\right)/\gcd(a,b)=1/2$ also occurs.
 Proof.
@@
 Otherwise, consider below the sequences $\left(x_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ and $\left(y_{k}/\gcd(a,b)\right)_{k=0}^{k=\infty}$ instead of $\left(x_{k}\right)_{k=0}^{k=\infty}$ and $\left(y_{k}\right)_{k=0}^{k=\infty}$.
+Since $x_{0}=a$, $x_{1}=\left(at+bdu\right)/2$, $x_{-1}=\pm\left(at-dbu\right)/2$, $y_{0}=b$, $y_{1}=\left(au+bt\right)/2$, $y_{-1}=\pm\left(au-bt\right)/2$, the recurrence relation satisfied by both sequences has integer coefficients and $\gcd(a,b)$ is odd, we see that $x_{k}/\gcd(a,b)$ and $y_{k}/\gcd(a,b)$ are integers if and only if $x_{k}$ and $y_{k}$ are integers.
 We start with some relationships that we will require for the proof itself.
@@
 Hence
-$\gcd\left(x_{k},y_{k}\right)\text{ divides }\left\{\begin{array}[]{ll}\gcd\left(x_{0},y_{0}\right)=1&\text{if $k$ is even},\\ \gcd\left(x_{1},y_{1}\right)|\left(2\gcd\left(x_{0},y_{0}\right)\right)=2&\text{if $k$ is odd}.\end{array}\right.$
+[…]
 (iv) We cannot have $t$ odd and $u$ even because $t^{2}-du^{2}\equiv 0\pmod{4}$. ∎
@@
 (1) Suppose that $p>2$ is prime.
-Put $v_{p}\left(g^{2}\right)=\ell\geq 0$.
+Put $v_{p}\left(g^{2}\right)=v_{p}\left(g_{1}^{2}g_{2}/g_{3}\right)=\ell$.
+Since $g_{3}$ is a power of $2$, it follows that $v_{p}\left(g_{3}\right)=0$ and hence $\ell\geq 0$.
 We will prove that $v_{p}\left(\gcd\left(a^{2},d\right)\right)=\ell$.
@@
 Note that $\ell_{2}=0$ or $1$, since $g_{2}|\operatorname{sf}\left(N_{\alpha}\right)$ (from the definition of $g_{2}$, we have $g_{2}|t^{\prime}$ and here $t^{\prime}=\operatorname{sf}\left(N_{\alpha}\right)$).
-From the definition of $g_{3}$, we have $g_{3}|4$, so $v_{p}\left(g_{3}\right)=0$ for odd primes, $p$.
-Hence $\ell=\ell_{1}+\ell_{2}$.
+Since $v_{p}\left(g_{3}\right)=0$, we have $\ell=\ell_{1}+\ell_{2}$.
 Since $g_{1}|u_{2}$, by the definition of $g_{1}$, and from the expression for $u_{2}$ here, we have $p^{\ell_{1}}|\left(N_{\alpha}/\operatorname{sf}\left(N_{\alpha}\right)\right)$.
@@
 So $p^{\ell}|N_{\alpha}$.
-In the same way, we find that $p^{\ell_{1}+2\ell_{2}}|x_{k}^{2}$.
-By Lemma 3.10, since $p$ is odd, $p\nmid\gcd\left(x_{k},y_{k}\right)$.
+In the same way, we find that $p^{\ell_{1}+2\ell_{2}}|x_{k}^{2}$. So $p^{\ell}|x_{k}^{2}$ too.
+By Lemma 3.10 and $\gcd(a,b)=1$, since $p$ is odd, $p\nmid\gcd\left(x_{k},y_{k}\right)$.
 Hence, from $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}$, we have $p^{\ell}|d$.
@@
 Lemma 3.12.
-[…]
-[…]
+Suppose that $a$, $b$, $d$, $g$, $t^{\prime}$, $u_{1}$, $u_{2}$, $x_{k}$ and $y_{k}$ are as in Lemma 3.11, $d^{\prime}=u_{2}^{2}t^{\prime}/g^{2}$, as defined in (2.10), $\mathcal{N}_{d^{\prime},4}$ is as defined in (2.1) and put $b^{\prime}=N_{\alpha}/\gcd\left(a^{2},db^{4}\right)$.
+Then
+(3.36) $|g|\mathcal{N}_{d^{\prime},4}=\sqrt{\gcd\left(a^{2},db^{4}\right)}\,2^{1+\min\left(2,v_{2}(b^{\prime})/2\right)}\geq 2^{1+\min\left(2,v_{2}\left(N_{\alpha}\right)/2\right)}.$
 Proof.
@@
 Remark.
-Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
+Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, k=1k=-1, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$.
 Proof.
@@
 $\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}>0.091\sqrt{d}.$
-For $d\geq 155$, we find that $1.130.091\sqrt{d}>$ 1.13 holds.
+For $d\geq 155$, we find that $0.091\sqrt{d}>1.13$ holds.
 Similarly,
@@
 where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.36) and Lemma 3.5(c) with $u\geq 1$ to establish the first inequality.
-For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
+For $d\geq 109,000$, we find that $0.66\sqrt{d}>217$ holds.
 We complete the proof computationally, checking all the remaining pairs, $(a,d)$. Since $d$ is bounded and $N_{\alpha}<0$ (so $a^{2}<d$), there are only finitely many such pairs.
-For each pair, we check (3.37) and (3.38) by computing their left-hand sides for all $k\neq 0$ such that $20.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<$ 2 and $30021.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<$ 300.
+For each pair, we check (3.37) and (3.38) by computing their left-hand sides for all $k\neq 0$ such that $0.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<2$ and $21.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<300$.
 No counterexamples to (3.37) and (3.38) were found.
 A PARI/GP program took 112 seconds to run on a Windows 10 laptop with an Intel i7-9750H processor and 16 GB of RAM.
-The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, $k=-1$, $x_{k}=-59$ and $y_{k}=5$.
-The smallest value of the left-hand side of (3.38) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, $k=-1$, $x_{k}=-130$ and $y_{k}=11$.
+The smallest value of the left-hand side of (3.37) found for $d\geq 105$ was $1.139\ldots$, for $a=11$, $d=140$, $N_{\alpha}=-19$, k=1k=-1, $x_{k}=-59$ and $y_{k}=5$.
+The smallest value of the left-hand side of (3.38) found for $d\geq 105$ was $217.3\ldots$, for $a=10$, $d=140$, $N_{\alpha}=-40$, k=1k=-1, $x_{k}=-130$ and $y_{k}=11$.
 4. Proposition 4.1 and its proof
@@
 Let $\left(y_{k}\right)_{k=-\infty}^{\infty}$ be defined by (1.1).
-If $b=1$ and $-N_{\alpha}$ is a square, then there is at most one integer square among all distinct elements of $\left(y_{k}\right)_{k=-\infty}^{\infty}$ which satisfies
+If $b=1$ and $N_{\alpha}$ is a square, then there is at most one integer square among all distinct elements of $\left(y_{k}\right)_{k=-\infty}^{\infty}$ which satisfies
 $y_{k}>\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right).$
-Conjecture 1.2 would immediately follow when $-N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$.
+Conjecture 1.2 would immediately follow when $N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$.
 This is due to Lemma 3.5.
@@
 We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$.
-So $y_{\ell}>y_{k}\geq 4\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right\}$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
+So $y_{\ell}>y_{k}\geq 4\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right)$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
 We shall initially assume that
@@
 This is the condition in Lemma 3.13, which will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof in Subsection 4.6.
-As in Lemma 3.9, we put $\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
-$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|,$
-where $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k}$, $r_{\ell}$, $s_{k}$ and $s_{\ell}$ as in Proposition 3.1, which are associated with $\left(x_{k},y_{k}\right)$ and $\left(x_{\ell},y_{\ell}\right)$, respectively.
+As in Lemma 3.9, we put $\omega_{k}=\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
+$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|,$
+where $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $\left(r_{k},s_{k}\right)$ and $\left(r_{\ell},s_{\ell}\right)$ as in Proposition 3.1, which are associated with $\left(x_{k},y_{k}\right)$ and $\left(x_{\ell},y_{\ell}\right)$, respectively.
 From (3.31) in the proof of Lemma 3.9(b), we have
-$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
+$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
 Thus we can apply Lemma 3.7(a) with $c_{1}=0.127$ to find that
-(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
+(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega_{k}-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
 The equality on the left-hand side is from the equalities in (3.30).
-By our choice of $k$ and $\ell$, and by Lemma 3.9(b) when $-N_{\alpha}$ is not a square, $\zeta_{4}=\pm 1\in\mathbb{Q}\left(\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$.
+By our choice of $k$ and $\ell$, and by Lemma 3.9(b) when $N_{\alpha}$ is not a square, $\zeta_{4}=\pm 1\in\mathbb{Q}\left(\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$.
 This is important for us here as $\zeta_{4}\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)/\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ must be in an imaginary quadratic field in order to apply Lemma 2.1 to obtain a lower bound for the rightmost quantity in (4.2).
@@
 Recall from (2.12) that we take $k_{0}=0.89$.
-Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
-(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
-Also from Lemma 3.9(a), we have $|\varphi|<0.6$, so the condition $|\omega-1|<1$ in Lemma 2.2 is satisfied.
+Writing $\omega_{k}=e^{i\varphi_{k}}$, with $\pi<\varphi_{k}\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
+(4.8) $\ell_{0}=\mathcal{C}_{4,2}\left|\varphi_{k}\right|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
+Also from Lemma 3.9(a), we have $\left|\varphi_{k}\right|<0.6$, so the condition $\left|\omega_{k}-1\right|<1$ in Lemma 2.2 is satisfied.
 Let $q=x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ and $p=x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$.
@@
 (4.9) $|q|=\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}.$
-We are now ready to deduce the required contradiction from the assumption that there are two sufficiently large squares when $-N_{\alpha}$ is a square. We will break the proof into five parts.
+We are now ready to deduce the required contradiction from the assumption that there are two sufficiently large squares when $N_{\alpha}$ is a square. We will break the proof into five parts.
 With $r_{0}$ as in Lemma 2.1, we separate the case of $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$, from the case of $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$.
@@
 We start by determining an upper bound for $y_{\ell}$ for all $r_{0}\geq 1$ when $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$, since we will also need such a result in Subsection 4.4.
-From the equality in (4.2), along with Lemma 2.1(b) and (4.9), we have
-(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
+From (4.2), along with Lemma 2.1(b) and (4.9), we have
+(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
 Applying (2.12) and (4.7) to (4.10), we obtain
@@
 (4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
-We will now combine (4.12) with the gap principle in Lemma 3.8(a). to show that this case cannot occur.
-Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
+We will now combine (4.12) with the gap principle in Lemma 3.8(a) to show that this case cannot occur.
+Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
 $57.32^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
@@
 The last inequality holds because $\left|N_{\alpha}\right|=d-a^{2}<d$.
-By (4.1), we have $0.250.24/\left((1-c)d^{1/2}\right)<$ 0.25 if $c<0.9$.
+By (4.1), we have $0.24/\left((1-c)d^{1/2}\right)<0.25$ if $c<0.9$.
 We will see in Subsection 4.4 that $c=0.75$ is a good choice.
@@
 From the definitions of $p_{r_{0}}$, $q_{r_{0}}$ and $R_{r_{0}}$ in (2.9), along with parts (a) and (e) of Lemma 2.2, we have
-$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega^{1/4}-p_{r_{0}}\right|$
-$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
-$=$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
-(4.13) $\geq$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
+$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{p}{q}\right|=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega_{k}^{1/4}-p_{r_{0}}\right|$
+$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega_{k})\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega_{k})\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
+$=$ $\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega_{k})}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega_{k}\right)\right|$
+(4.13) $\geq$ $\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega_{k})}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
 From Lemma 2.2(d), we have
-(4.14) $\left|Y_{1,4,r_{0}}(\omega)\right|<1.072\frac{r_{0}!\Gamma(3/4)}{\Gamma(r_{0}+3/4)}\left|1+\sqrt{\omega}\right|^{2r_{0}}.$
-We can write $\omega-1$ as $\left(2N_{\alpha}+2x_{k}N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}^{2}-N_{\alpha}\right)$, so
-$|\omega-1|=\sqrt{\frac{4\left|N_{\alpha}\right|}{x_{k}^{2}-N_{\alpha}}}=2\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
-Since $\left|\sqrt{\omega}+1\right|<2$, it follows that
-$\left|\sqrt{\omega}-1\right|=\frac{\left|\omega-1\right|}{\left|\sqrt{\omega}+1\right|}>\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
+(4.14) $\left|Y_{1,4,r_{0}}\left(\omega_{k}\right)\right|<1.072\frac{r_{0}!\Gamma(3/4)}{\Gamma(r_{0}+3/4)}\left|1+\sqrt{\omega_{k}}\right|^{2r_{0}}.$
+We can write $\omega_{k}-1$ as $\left(2N_{\alpha}+2x_{k}N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}^{2}-N_{\alpha}\right)$, so
+$\left|\omega_{k}-1\right|=\sqrt{\frac{4\left|N_{\alpha}\right|}{x_{k}^{2}-N_{\alpha}}}=2\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
+Since $\left|\sqrt{\omega_{k}}+1\right|<2$, it follows that
+$\left|\sqrt{\omega_{k}}-1\right|=\frac{\left|\omega_{k}-1\right|}{\left|\sqrt{\omega_{k}}+1\right|}>\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
 From these two inequalities, we also obtain
-$\left|\frac{(\omega-1)^{2r_{0}+1}}{\left(1+\sqrt{\omega}\right)^{2r_{0}}}\right|=\left|(\omega-1)\left(\sqrt{\omega}-1\right)^{2r_{0}}\right|>2\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
-Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}(\omega)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
-$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
+$\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{\left(1+\sqrt{\omega_{k}}\right)^{2r_{0}}}\right|=\left|\left(\omega_{k}-1\right)\left(\sqrt{\omega_{k}}-1\right)^{2r_{0}}\right|>2\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
+Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}\left(\omega_{k}\right)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
+$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
 Applying (4.2) with this inequality, we obtain
@@
 $6.96\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.32\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.32\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
-since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
-4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
+since $N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
+$4.$ $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
 Here we establish a stronger gap principle for $y_{k}$ and $y_{\ell}$ than the one in Lemma 3.8. We then use this with the upper bound for $y_{\ell}$ in (4.11) to obtain a contradiction.
@@
 (4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
-Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
+Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
 $\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
@@
 We now proceed similarly to the way we did in Subsection 4.4.
-Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b).
+Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b).
 Also, as $r_{0}-1\geq 1$, it follows that $0.7405\cdot 0.0002344^{r_{0}-1}>0.0001735^{r_{0}-1}$.
@@
 To complete the proof of Proposition 4.1, we need to remove the assumption on $d$ in (4.1).
-We check directly all pairs $(a,d)$ of positive integers with $2\leq d\leq 104$ not a square and $-N_{\alpha}=d-a^{2}$ a square.
+We check directly all pairs $(a,d)$ of positive integers with $2\leq d\leq 104$ not a square and $N_{\alpha}=d-a^{2}$ a square.
 Note that there are only finitely many such pairs, since $N_{\alpha}=a^{2}-d<0$.
@@
 Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}\geq 81\left|N_{\alpha}\right|/4>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$ for $k\neq 0$ and we need only consider $1\leq u\leq 8$.
-Next we treat the case when $K<-1$, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
+Next we treat the case when K<1K<-1, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
 Lemma 5.1.
-If $K<-1$, then there are at most two distinct integer squares among the $y_{k}$’s.
+If K<1K<-1, then there are at most two distinct integer squares among the $y_{k}$’s.
 Proof.
-From Lemma 3.6, we know that $K<-1$ can only happen in the following two cases:
+From Lemma 3.6, we know that K<1K<-1 can only happen in the following two cases:
 (i) $a\geq 1$, $d=a^{2}+4$, $t=a$ and $u=1$, where $\alpha=2\varepsilon$ and $N_{\alpha}=-4$,
@@
 In case (ii), we can also apply Theorem 1.1 of [1]. ∎
-For the remainder of this section, we may assume that $K=-1$.
+For the remainder of this section, we may assume that K=1K=-1.
 Lemma 5.2.
@@
 So we need only consider $u=1$ when $N_{\alpha}$ is even.
-Similarly, if $-N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
+Similarly, if $N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
 But, we saw at the start of this section that for $N_{\alpha}$ odd, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$.
@@
 For $u=1$, we proceed similarly.
-Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $-N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $-N_{\alpha}$ is an odd square.
+Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $N_{\alpha}$ is an odd square.
 For each value of $N_{\alpha}$, the norm of $\varepsilon$ leads us to an equation of the form $t^{2}-a^{2}=-N_{\alpha}\pm 4$.
@@
 From Lemmas 3.8(a) and  3.5(c), we have
-$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
+$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}^{3}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
 By Proposition 4.1 and Lemma 3.12, if the quantity on the right-hand side exceeds $\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$, then the lemma follows.
@@
 Lemma 5.4.
-Suppose that $-N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
+Suppose that $N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
 (a) If $t^{2}-du^{2}=4$, then $t\equiv 2\pmod{4}$ and $u\equiv 0\pmod{4}$.
@@
 If $4\mid t$, then $du^{2}\equiv 12\pmod{16}$.
-Since $-N_{\alpha}$ is an odd square, we have $-N_{\alpha}\equiv 1,9\pmod{16}$.
+Since $N_{\alpha}$ is an odd square, we have $N_{\alpha}\equiv 1,9\pmod{16}$.
 If $a$ is odd, then $a^{2}\equiv 1\pmod{8}$.
@@
 If $a$ is even, then $a^{2}\equiv 0,4\pmod{16}$.
-Since $-N_{\alpha}=d-a^{2}\equiv 1,9\pmod{16}$, it follows that $d\equiv 1\pmod{8}$.
+Since $N_{\alpha}=d-a^{2}\equiv 1,9\pmod{16}$, it follows that $d\equiv 1\pmod{8}$.
 So $u^{2}\equiv 12\pmod{16}$, which is also impossible.
@@
 From $t\equiv 2\pmod{4}$ and $t^{2}-du^{2}=4$, we have $du^{2}\equiv 0\pmod{16}$.
-Since $N_{\alpha}$ is odd, we have $-N_{\alpha}\equiv 1,9\pmod{16}$. Also $a^{2}\equiv 0,1,4,9\pmod{16}$. Hence $d\equiv 1,2,5,9,10,13\pmod{16}$ and so $8|u^{2}$. Thus $4|u$.
+Since $N_{\alpha}$ is odd, we have $N_{\alpha}\equiv 1,9\pmod{16}$. Also $a^{2}\equiv 0,1,4,9\pmod{16}$. Hence $d\equiv 1,2,5,9,10,13\pmod{16}$ and so $8|u^{2}$. Thus $4|u$.
 (b) We proceed similarly.
@@
 Lemma 5.5.
-If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
+If $N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
 Proof.
@@
 (1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
-From Lemma 5.4(a), if $t^{2}-du^{2}=4$ and $-N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
+From Lemma 5.4(a), if $t^{2}-du^{2}=4$ and $N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
 (1-i) First, we eliminate $u=4$.
@@
 Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$.
-From this, $-N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
+From this, $N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
 We will show that $d\equiv 5\pmod{9}$ is not possible.
@@
 But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 16\left|N_{\alpha}\right|$ here.
-So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
+So the case where $u=8$, $N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
 Thus for $N_{\alpha}$ odd, we do not need to consider $t^{2}-du^{2}=4$.
@@
 Hence $p\equiv 1\pmod{4}$.
-Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$.
+Since $p^{2}|a^{2}$ and $N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$.
 Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 100$.
@@
 But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$ here.
-So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$.
+So the theorem holds if $u=2$, $N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$.
 ∎
@@
 Lemma 5.6.
-If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.
+If $N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.
 Proof.
@@
 (1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
-(1-i) We show that $u=4$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
+(1-i) We show that $u=4$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 Here we can argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$.
@@
 So we can exclude $u=4$ and $t^{2}-du^{2}=4$ from consideration.
-(1-ii) We show that $u=3$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
+(1-ii) We show that $u=3$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 We will again argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $32$ to show that $64|\gcd\left(a^{2},d\right)$.
@@
 But by Lemma 3.5(c), we know that $y_{\pm 1}\geq(9/4)\left|N_{\alpha}\right|$ here.
-(1-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
+(1-iii) We show that $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 Using the argument for $u=8$ in the proof of Lemma 5.5 and the Maple program mentioned there modulo $16$, we obtain $16|\gcd\left(a^{2},d\right)$.
@@
 So we can exclude $u=2$ and $t^{2}-du^{2}=4$ from consideration.
-(1-iv) We show that $u=1$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
+(1-iv) We show that $u=1$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
 Here too we use the argument and Maple program from the proof of Lemma 5.5, first modulo $9$ and then modulo $16$.
@@
 Independent of the parity of $N_{\alpha}$, for $u=3$ and $4$, $t^{2}-du^{2}=-4$ is not possible modulo $9$ and $16$, respectively. So we can only have $u=1,2$ with $t^{2}-du^{2}=-4$
-(2-i) We show that if $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
-We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $-N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
+(2-i) We show that if $u=1$, $N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
+We find that $y_{\pm 1}\equiv 0,1,4,9\pmod{16}$ is not possible if $N_{\alpha}=b_{1}^{2}$ where $b_{1}\equiv 4\pmod{8}$.
 If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d^{\prime},4}\geq 16$.
@@
 So we can exclude $8|b_{1}$, provided $d\geq 80$.
-For $d<80$ with $u=1$ and $-N_{a}$ an even square divisble by $64$, there is just one possibility: $d=68$, $t=8$, $u=1$, $a=2$, so $N_{\alpha}=-64$ and $y_{-1}=25$.
+For $d<80$ with $u=1$ and $N_{a}$ an even square divisble by $64$, there is just one possibility: $d=68$, $t=8$, $u=1$, $a=2$, so $N_{\alpha}=-64$ and $y_{-1}=25$.
 But here, $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}<\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}<19$, so this case is excluded too and hence we can exclude $8|b_{1}$ altogether.
-Hence $b_{1}\equiv 2\pmod{4}$ and so $-N_{\alpha}\equiv 4\pmod{16}$.
-(2-ii) We show that if $u=1$, $-N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
-We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $-N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 00 in the second case.
+Hence $b_{1}\equiv 2\pmod{4}$ and so $N_{\alpha}\equiv 4\pmod{16}$.
+(2-ii) We show that if $u=1$, $N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
+We have $v_{2}\left(\gcd\left(a^{2},d\right)\right)=0$ or $2$. Since $N_{\alpha}\equiv 4\pmod{16}$, we have $v_{2}\left(b^{\prime}\right)=2$ in the first case and 0 in the second case.
 We consider first $\gcd\left(a^{2},d\right)>4$ odd.
@@
 Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$.
-If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $40.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/$ 4, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
+If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $0.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/4$, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
 Since $d=t^{2}+4$, arguing modulo $3$, we see that $3\nmid d$ and hence $\gcd\left(a^{2},d\right)=9$ is not possible.
-Now consider $\gcd\left(a^{2},d\right)>4$ even.
-Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$).
-So by Lemma 3.12 and the argument above, $|g|\mathcal{N}_{d^{\prime},4}=2\sqrt{\gcd\left(a^{2},d\right)}$.
-Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
+[…]
 We saw above that $3\nmid d$, so we can ignore $\gcd\left(a^{2},d\right)=6^{2}$.
 If $\gcd\left(a^{2},d\right)\geq 100$, then $\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}\leq 0.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/4$, the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
-(2-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
+(2-iii) We show that $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
 Arguing modulo $64$, we have $d\equiv a^{2}\equiv 1\pmod{16}$ and $8|t$.
@@
 But $76\left|N_{\alpha}\right|/64=1.1875\left|N_{\alpha}\right|$.
-So we can exclude the case where $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$.
+So we can exclude the case where $u=2$, $N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$.
 ∎
@@
 $a$ $b$ indices, $k$ $\sqrt{y_{k}}$
-1 $3$ [0, -1, 3, -5] [3, 5, 31, 167]
+1 33 [0, -1, 3, -5] [3, 5, 31, 167]
 1019 $27$ [0, 1, -3, -7] [27, 65, 29, 983]
@@
 Let $n\geq 5$ be an odd integer.
-Put $a=\left(n^{2}-9\right)/4$, $d=\left(n^{4}-2n^{2}+17\right)/16$, so that $N_{\alpha}=4-n^{2}$.
+Put $a=\left(n^{2}-9\right)/4$ and $d=\left(n^{4}-2n^{2}+17\right)/16$, so that $N_{\alpha}=4-n^{2}$.
 With $t=\left(n^{2}-1\right)/2$ and $u=2$, $\varepsilon=\left(t+u\sqrt{d}\right)/2$ is a unit in $\mathcal{O}_{\mathbb{Q}\left(\sqrt{d}\right)}$.
@@
 From (3.14) with $b=u=1$, we find that $(2t+a)^{2}-8y_{1}=a^{2}-8$.
-Since $-N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $22t+$ 2 is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$.
+Since $N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $2t+2$ is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$.
 The corresponding values of $\sqrt{y_{1}}$ are the elements of the sequence $\left(u_{n}\right)_{n\geq 0}$ with $u_{0}=5$, $u_{1}=169$ and $u_{n+1}=34u_{n}-u_{n-1}$.
@@
 Then $(t,u)=(a+1,1)$, $y_{-1}=n^{2}$ and so $y_{-1}/\left|N_{\alpha}\right|\rightarrow 1/4$ from above as $n\rightarrow+\infty$.
-There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $-N_{\alpha}$ is a square too.
-The sequences given in Lemma 3.6 show that for $k=-1$, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.
+There exist families with the same limit for $y_{-1}/\left|N_{\alpha}\right|$ when $N_{\alpha}$ is a square too.
+The sequences given in Lemma 3.6 show that for k=1k=-1, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.